### v2

# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

Abstract
This paper presents a novel approach to four fundamental problems in the field of cosmological perturbation theory. Firstly, the issue of gauge dependence has been addressed by demonstrating the existence of unique and gauge-invariant quantities corresponding to the actual perturbations. Secondly, the formation of primordial structures after decoupling of matter and radiation is dependent on the existence of local fluid flows resulting from local pressure gradients. To take pressure gradients into account, it is necessary to consider both the energy density and the particle number density. Thirdly, the novel relativistic perturbation theory applies to an open, flat, and closed Friedmann-Lemaître-Robertson-Walker universe. The derivation of the novel perturbation theory definitively reveals the inherent limitations of Newtonian gravitation as a framework for investigating cosmological density perturbations. Finally, the application of the perturbation theory to a flat universe demonstrates that, prior to decoupling, perturbations in both Cold Dark Matter and ordinary matter are coupled to perturbations in radiation. Therefore, the universe’s earliest structures formed only after decoupling, at which point local nonadiabatic random pressure fluctuations became a significant factor. Negative nonadiabatic pressure fluctuations resulted in a brief, rapid growth of density fluctuations until the total pressure fluctuations became positive. In contrast, positive nonadiabatic pressure fluctuations led to the formation of voids. Perturbations with masses about $2.2\times 10^{4}\,\text{M}_{\odot}$ became nonlinear already $60$ million years after the Big Bang, and perturbations with masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$ became nonlinear within about $600$ million years.
Cosmology; Early Universe; Structure Formation; Perturbation Theory; Gauge Problem.
pacs:
98.80.-k; 98.80.Cq; 97.10.Bt; 04.25.Nx.
I Introduction
The linearized Einstein equations and conservation laws of Friedmann-Lemaître-Robertson-Walker (flrw) universes are crucial to cosmology. They can explain the formation of all kinds of structures in the expanding universe, including stars, galaxies, and microwave background fluctuations.
Lifshitz [1] and Lifshitz & Khalatnikov [2] were the first researchers to develop a cosmological perturbation theory to study the evolution of density perturbations in the universe. They encountered the problem that the solutions of the linearized Einstein equations and conservation laws can be modified by linear coordinate transformations, which are known as gauge transformations. Consequently, the solutions are contingent upon the selection of a coordinate system, rendering them gauge-dependent and thus devoid of physical significance. This is the notorious gauge problem of cosmology, which demands urgent resolution. A multitude of potential solutions to this problem have been proposed in the literature. Stewart [3] noted in his paper, […] perturbation theory within general relativity is generally regarded as a somewhat arcane subject […] . This observation is further substantiated by the lack of progress observed to date in this field. The purpose of this paper is to elucidate the hitherto puzzling aspects of cosmological perturbation theory.
Before delving into an examination of the approaches outlined in the extant literature, it is imperative to acknowledge the fundamental criteria that a comprehensive and precise cosmological perturbation theory must fulfill. Primarily, as Bardeen elucidated in his seminal paper [4], a quantity signifying the real physical energy density perturbation must be manifestly gauge-invariant, that is, it must be defined independently of any reference system. Secondly, in the nonrelativistic limit, the linearized energy density constraint equation, when considered together with the linearized momentum constraint equation, must yield the time-independent Poisson equation of Newtonian theory with the aforementioned gauge-invariant quantity as the source term. Thirdly, in the nonrelativistic limit, it is essential that the freedom to select a coordinate system is present. That is to say, relativistic gauge transformations must converge to Newtonian gauge transformations in the nonrelativistic limit. Finally, and equally important, the perturbation theory must encompass all previously established and well-known physical solutions for large-scale density perturbations.
The present paper proposes a novel cosmological perturbation theory that has been demonstrated to satisfy all four of the aforementioned criteria. A review of the existing literature reveals that there is no theory that meets all four of the above criteria.
Bardeen [4] was the first to demonstrate that the linearized Einstein equations and conservation laws can be recast into evolution equations that have solutions independent of the choice of a coordinate system by using gauge-invariant quantities. In his paper, Bardeen constructed two distinct gauge-invariant quantities for the perturbation to the energy density, designated as $\epsilon_{m}$ and $\epsilon_{g}$. In the limit of small scales, these quantities converge to the usual gauge-dependent quantity, denoted as $\varepsilon_{{\scriptscriptstyle(1)}}$. Bardeen presupposed that on small scales, Newtonian gravitation is a valid description of cosmological density perturbations, and therefore, the gauge-dependent quantity, $\varepsilon_{{\scriptscriptstyle(1)}}$, becomes equal to its Newtonian counterpart. It is, however, a well-known fact that in Newtonian gravitation the coordinates of time and space can be chosen at will. Correspondingly, this must also be true in the nonrelativistic limit of a cosmological perturbation theory. It can thus be concluded that $\varepsilon_{{\scriptscriptstyle(1)}}$, which depends on the choice of coordinates in general relativity, is also dependent on the choice of coordinates in Newtonian gravitation. Thus, quantities that lack physical significance in general relativity are similarly devoid of physical meaning in Newtonian gravitation. It follows that the small-scale limit is not equivalent to the nonrelativistic limit. In consequence, the quantities $\epsilon_{m}$ and $\epsilon_{g}$ do not correspond to the actual density perturbations.
Bardeen’s paper has inspired a number of works, including those in Refs. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These researchers proposed alternative perturbation theories using gauge-invariant quantities that differ from those used by Bardeen.
The authors of Refs. [5, 6, 7, 9, 11] define gauge-invariant density perturbations by gradients of the energy density. For a pressureless fluid, their equations are identical to the standard perturbation equation derived from Newton’s theory of gravitation. This equation yields the well-known solutions of the standard equation (150) with $w=0$ and $\beta=0$, including the gauge mode. In addition to the fact that the solution is gauge-dependent, even in the nonrelativistic limit, the evolution of density perturbations in a pressureless fluid is precluded by the absence of pressure gradients. This is a consequence of the linearized momentum conservation laws. It can thus be concluded that the standard equation (150) with $w=0$ and $\beta=0$ cannot be taken to represent the nonrelativistic limit of a cosmological perturbation theory. Moreover, as indicated in Ref. [6], the constraint equations have not been integrated into the analysis. It is of crucial importance to acknowledge that the constraint equations represent relations between initial values, thereby playing a pivotal role in the analysis. In the absence of the constraint equations, the solutions to the system of dynamical equations and conservation laws lack physical significance. Furthermore, it is not possible to achieve a correct Newtonian limit without the constraint equations (for further details, see Sec. IX). In this paper, the complete set of Einstein equations and conservation laws is employed to develop a novel perturbation theory.
In order to interpret the gauge-invariant definition of the density perturbation $\overline{\delta\varepsilon}$, as proposed by Mukhanov et al. [8, 10], it is necessary to set the Hubble function $\cal{H}$ to zero in Eq. (7.47) in Ref. [10]. Nevertheless, this yields a Poisson-like equation that differs from the Newtonian Poisson equation, as the potential is time-dependent in the Poisson-like equation, as evidenced by equations (7.48) and (7.49) in Ref. [10]. Thus, $\overline{\delta\varepsilon}$ does not represent the actual physical density perturbation. Furthermore, the Hubble function cannot be set to zero, as this would violate the Friedmann equation (10a), which is inconsistent with the assumption that the universe is not empty. The present paper introduces a novel perturbation theory that achieves the time-independent Poisson equation and thus the correct nonrelativistic limit by considering zero pressure alone. This obviates the need to set the Hubble function to zero.
As has been stated in Refs. [12, 13], the linearized Einstein equations and conservation laws for density perturbations yield the usual evolution equation (150) derived from Newtonian gravitation when the pressure is zero. The current study will demonstrate that in a pressureless fluid, the linearized Einstein equations and conservation laws are reduced to a single equation: the time-independent Poisson equation of Newtonian gravitation.
The authors of Refs. [14, 15] present a method for the construction of gauge-invariant quantities in the synchronous reference system. However, the definition of gauge invariance presented in these works does not align with the conventional understanding of what constitutes a gauge-invariant quantity. As defined in Refs. [4, 3], a gauge-invariant quantity is independent of the choice of a coordinate system. Consequently, the definition of gauge-invariant quantities is formulated in this paper without the utilization of a coordinate system.
In their study of the evolution of density perturbations in a flat flrw universe in its radiation-dominated phase, Ma & Bertschinger [16] employ two distinct coordinate systems: the synchronous ($g_{0i}=0$) gauge and the conformal Newtonian gauge. The latter gauge is, in fact, a restricted synchronous gauge. Given the existence of gauge transformations between different synchronous reference systems, the solutions are dependent on the selected synchronous coordinates. Consequently, gauge modes emerge in the solutions, rendering them devoid of physical significance even when physical initial values are imposed. Their findings pertaining to the density fluctuation, represented by $\delta$, deviate between the synchronous and conformal Newtonian coordinates. In addition, Eq. (29) in Ref. [16] is flawed in two respects. Firstly, the factor $(1-3w)$ in the expansion term should, in fact, be $(2-3w)$, which is critical in the context of a radiation-dominated universe, where $w=\tfrac{1}{3}$. Secondly, it is crucial to replace the expression $k^{2}$ with $k^{2}/a^{2}$. This substitution is necessary because the background three-space metric of a flat flrw universe is $a^{2}\delta_{ij}$ rather than $\delta_{ij}$. This can be demonstrated by considering the line element (1) in Ref. [16], which illustrates that three-space is expanding. Once these corrections are made, the resulting equation is equal to the covariant divergence of Eq. (43) in this paper. The latter equation is correct, as demonstrated in Appendices B and C.
The aforementioned treatments resulted in a plethora of additional contributions to this field. See, e.g., Refs. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In a review, Ellis [28] discusses the work of Lifshitz and Khalatnikov and other research on the subject. A review of the literature reveals a lack of consensus concerning the accuracy of the quantities employed to represent physical density perturbations.
This paper addresses the issue of gauge dependence in the context of cosmology. It demonstrates the existence of unique gauge-invariant quantities that correspond to the true physical density perturbations. Subsequently, it presents equations for the study of the evolution of density perturbations in universes with an open, flat, or closed flrw geometry. In light of the general covariance of general relativity, the evolution equations for gauge-invariant quantities are mathematically identical in all coordinate systems. Consequently, they yield the same solutions in any given coordinate system.
The novel perturbation theory, when applied to a flat flrw universe, reveals that prior to the decoupling of radiation and matter, Cold Dark Matter (cdm) perturbations were coupled to the radiation energy density perturbations in such a way that they could not clump together. Consequently, cdm proved to be ineffective in facilitating the process of structure formation subsequent to the decoupling of radiation and matter. The formation of the first stars, designated as Population iii stars [29], can be explained by the rapid growth of small density perturbations over a brief period of time. This rapid growth arises from local nonadiabatic random pressure fluctuations in the early universe following the decoupling. The James Webb Space Telescope (jwst), which was launched on December 25, 2021, is designed primarily for near-infrared astronomy and may therefore provide insights regarding the existence and nature of the earliest stars. The data from the jwst [30] suggest that the galaxies observed in the early universe appear to have grown to a size greater than anticipated, forming massive and compact structures at a rate exceeding the expectations of the hierarchical $\Lambda\textsc{cdm}$ structure formation paradigm. This may indicate that the existence of cdm is not supported by the data.
II A Novel Cosmological Perturbation Theory
This paper presents a novel cosmological perturbation theory, one that differs substantially from the perturbation theories in the extant literature, including the treatment of the cosmic fluid. The aim of this section is to provide an overview of the novel approach and to elucidate the method by which the new theory is constructed.
II.1 Cosmic Fluid
It is well established that in the era following the decoupling of matter and radiation, the pressure in the cosmological fluid is negligible in comparison to the energy density. Therefore, it can be concluded that the pressure can be disregarded in the Einstein equations and conservation laws of the background, i.e., the unperturbed, flrw universe. In the context of cosmology, it is commonly assumed that in the perturbed universe, the pressure and, therefore, local pressure gradients can also be neglected. However, it is well known that in the absence of local pressure gradients, local fluid flows cannot occur. This is equally true in the case of the cosmic fluid. As will be demonstrated in Sec. IX, the linearized momentum conservation laws indicate that, in the absence of pressure, local fluid flows will not occur and density perturbations will be static. It is thus crucial to consider pressure gradients in order to examine the evolution of density perturbations. Given that, subsequent to decoupling, the cosmic fluid is to be considered a perfect gas where the pressure is dependent on the mean kinetic energy of its particles, it is essential that the kinetic energy density be incorporated into the analysis. Both the pressure and the kinetic energy density are dependent on the particle number density and temperature. Consequently, in addition to the energy density, it is necessary to include the particle number density and the temperature in the equation of state. See Ref. [31], Sec. 2.10 on relativistic hydrodynamics, and Ref. [32]. As indicated in Ref. [33], the formation of structures in the universe requires a focus on the clumping of a set of particles rather than on the evolution of energy density instabilities. The approach presented in this paper yields novel insights into the evolution of density perturbations in both the pre-decoupling and post-decoupling eras. This is achieved by examining the evolution of energy density and particle number density perturbations, as well as their interactions. A detailed explanation of this can be found in Sec. XII.
II.2 Perturbation Theory
The equations governing the evolution of the unperturbed universe include three scalars: the energy density, the particle number density, and the expansion. As will be demonstrated in Sec. III, the perturbations to these three scalars transform under general linear coordinate transformations in the same way. This ultimately provides a solution to the gauge problem. As will be demonstrated in Sec. IV, there are only three sets of three gauge-invariant quantities that can be constructed from these scalars and their gauge-dependent perturbations. Each of these sets contains precisely one gauge-invariant quantity that is equal to zero. It will be shown that if the gauge-invariant perturbation of the expansion is zero, then it follows from the nonrelativistic limit in Sec. IX that the other two gauge-invariant quantities, denoted by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, are the true physical perturbations of the energy density and the particle number density, respectively.
The subsequent step is to derive the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$. In order to do this, it is necessary to employ a suitable coordinate system. In the context of general relativity, there is no a priori preferred coordinate system. Furthermore, the perturbations of the energy and particle number densities are both gauge-invariant, allowing for the freedom to choose any coordinate system. In order to facilitate the derivation of the aforementioned equations, it is typical to select a coordinate system that aligns with the specific characteristics of the problem in question. In Section Sec. V, a detailed justification is provided for the selection of synchronous coordinates. In Sec. VI, the Einstein equations, conservation laws, and their linearized counterparts with respect to this coordinate system are presented.
In Sec. VII, the synchronous reference system is employed to decompose the coupled system of linearized Einstein equations and conservation laws into three independent systems. These systems describe the evolution of tensor, vector, and scalar perturbations, respectively. In the extant literature, it is commonly assumed that the decomposition of the linearized equations can be achieved by decomposing only the perturbation of the metric tensor into its tensor, vector, and scalar parts. However, this is insufficient. Additionally, the perturbation of the spatial Ricci tensor, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}$, must similarly be decomposed into its tensor, vector, and scalar parts, as demonstrated in Ref. [34].
In the case of tensor and vector perturbations, the perturbation of the spatial Ricci scalar, or the spatial curvature perturbation, denoted by ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, is identically zero. In contrast, for scalar perturbations, the value of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ is not zero. In addition, it will be demonstrated that the Newtonian potential is encapsulated within the expression for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$. This implies that only scalar perturbations are associated with density perturbations, specifically the physical density perturbations, represented by $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
In Sec. VIII, the system of equations (37) is derived which describes the evolution of scalar perturbations. This system incorporates not only the usual conservation laws for energy density and momentum, but also a conservation law for the particle number density. Furthermore, it is found that this system comprises an algebraic energy density constraint equation (37a) that incorporates the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, in addition to an evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$. The latter equation is the covariant divergence of the three linearized momentum constraint equations (12b). In the nonrelativistic limit (i.e., the limit in which the pressure vanishes), the algebraic energy density constraint equation (37a) and the evolution equation (37b) for ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, when combined, yield the time-independent Poisson equation of Newtonian gravitation with source term $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x})$, as detailed in Sec. IX. Consequently, the constraint equations (37a)–(37b) are of great importance in the evolution of density perturbations and in the derivation of the nonrelativistic limit. The constraint equations presented in this specific formulation are notably absent from the extant literature on cosmological perturbation theories. This is the precise reason why an exact and correct nonrelativistic limit for a cosmological perturbation theory has not yet been discovered. The absence of a clearly defined nonrelativistic limit makes it impossible to interpret gauge-invariant quantities in previous perturbation theories. The issue is further discussed in Sec. IX, which addresses the nonrelativistic limit.
In Sec. X, the primary result is presented, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/n_{{\scriptscriptstyle(0)}}$ in closed, flat, and open flrw universes. As the quantities $\delta_{\varepsilon}$ and $\delta_{n}$ are gauge-invariant, the evolution equations will have the same mathematical form in all coordinate systems and yield solutions that are independent of the choice of a coordinate system.
In Sec. XII, the recently developed perturbation theory is applied to a flat flrw universe in its phases preceding and following the decoupling of radiation and matter. In order to proceed with the analysis, it is necessary to derive gauge-invariant thermodynamic quantities. These quantities will be presented in Sec. XI. Finally, in Sec. XIII, a summary of the results is provided.
The paper is accompanied by three appendices. In Appendix A, the primary result, Eqs. (60), is derived using the background equations (10), and the equations for scalar perturbations (37). The latter system represents a novel result and is indispensable for the proposed methodology. A proof of its correctness is provided in Appendices B and C. In Appendix C, the systems of equations (10) and (37) are employed once more. In this instance, however, the evolution equations (154) and (156) are recovered for the gauge-dependent quantity defined by $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, in a flat flrw universe. From these equations, the well-known physical solutions for large-scale perturbations in a flat flrw universe are derived. These include the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe following the decoupling of matter and radiation. A comparison of Eqs. (85) and (155a) and Eqs. (109) and (157a) reveals that the novel perturbation theory encompasses all previously established and well-known physical solutions for large-scale density perturbations. This constitutes a significant result, as the aforementioned solutions are widely documented in the literature. It follows that a comprehensive and accurate perturbation theory must also yield these solutions.
The homogeneous part of Eqs. (154a) and (156a) constitutes the conventional relativistic equation (150). Evidently, the latter equation is incomplete and has the gauge mode $\propto t^{-1}$ as solution. As will be demonstrated in Sec. IX, in the nonrelativistic limit, relativistic gauge transformations converge to Newtonian gauge transformations. Consequently, the quantity $\delta$ is also dependent on the choice of coordinates in Newtonian gravitation. Accordingly, it possesses no physical meaning. It can be concluded that the conventional relativistic evolution equation (150), is inadequate for examining the evolution of density fluctuations in the universe.
In the extant literature, the relativistic equation (150) is derived from Newtonian gravitation under two conditions: first, the expansion of the universe must be taken into account, and second, the equation of state must be nonrelativistic. The presence of the gauge mode in the solution of the aforementioned equation, when derived from the Newtonian theory of gravity, is the reason why Newtonian gravitation is not applicable to the study of density fluctuations in an evolving universe.
III Gauge Problem of Cosmology
Einstein’s equations and conservation laws are invariant under general coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. This signifies that there are no preferred coordinate systems. Consequently, the linearized Einstein equations and conservation laws are invariant under a general linear infinitesimal space-time transformation, which is usually referred to as a gauge transformation. This transformation is given by
$x^{\mu}\rightarrow x^{\prime\mu}=x^{\mu}-\xi^{\mu}(t,\bm{x}),$ (1)
where the gauge functions $\xi^{\mu}(t,\bm{x})$ are four arbitrary infinitesimal functions of time, $x^{0}=ct$ with $c$ the speed of light, and space, $\bm{x}=(x^{1},x^{2},x^{3})$, coordinates.
Consider a closed, flat, or open flrw universe. The evolution equations for these universes contain three scalars: the energy density $\varepsilon$, the particle number density $n$, and the expansion $\theta$. These scalars are defined as follows
$\varepsilon:=T^{\mu\nu}u_{\mu}u_{\nu},\quad n:=N^{\mu}u_{\mu},\quad\theta:=u^{\mu}{}_{;\mu},$ (2)
where $u^{\mu}$ is the fluid four-velocity normalized to unity, ${u^{\mu}u_{\mu}=1}$, and $N^{\mu}:=nu^{\mu}$ represents the cosmological particle current four-vector, which satisfies the particle number conservation law ${N^{\mu}{}_{;\mu}=0}$. A semicolon denotes covariant differentiation.
Let $S_{\scriptscriptstyle(0)}(t)$ denote the quantities $\varepsilon_{\scriptscriptstyle(0)}$, $n_{\scriptscriptstyle(0)}$, and $\theta_{\scriptscriptstyle(0)}$. These quantities satisfy the background, that is to say, the unperturbed Einstein equations and conservation laws. Let $S_{\scriptscriptstyle(1)}(t,\bm{x})$ be their perturbed counterparts $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, which satisfy the linearized equations. As a consequence of the linearity of these equations, new solutions that are physically equivalent can be generated. These solutions are given by (see Weinberg [31], Sec. 10.9 for a detailed explanation)
$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)}.$ (3)
In this expression, the operator $\mathcal{L}_{\xi}$ is the Lie derivative with respect to the infinitesimal four-vector $\xi^{\mu}$. An overdot denotes differentiation with respect to $x^{0}=ct$. The solutions $S_{\scriptscriptstyle(1)}$ and $S^{\prime}_{\scriptscriptstyle(1)}$ contain the so-called gauge modes which are given by $\hat{S}_{\scriptscriptstyle(1)}=\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This indicates that the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are dependent on the choice of a system of reference, rendering them gauge-dependent and, consequently, devoid of physical significance. This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution of the linearized equations. Even when physical initial values are imposed in order to solve the linearized equations, the resulting solution lacks physical meaning. It is therefore imperative to resolve the gauge problem in order to facilitate an investigation into the evolution of density perturbations in the universe.
IV Solving the Gauge Problem of Cosmology
In contrast to the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, their physical counterparts are independent of the choice of a coordinate system. This is to say that they are gauge-invariant. As the physics of the density perturbations is concealed within the general solution of the linearized Einstein equations and conservation laws, it follows that the physical counterparts of $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, must be linear combinations of gauge-dependent solutions to these equations. In order to identify these quantities, it is necessary to consider the fact that a quantity is gauge-invariant if it is independent of the coordinate transformation parameter $\xi^{\mu}$, as outlined by Stewart [3]. This implies that the quantity must be invariant under a linear coordinate transformation (1). As is evident from (3), the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ are independent of the spatial components, $\xi^{i}$, of the transformation, but not of the temporal component, $\xi^{0}$. Given that the transformation of these quantities under the general infinitesimal transformation (1) is identical to that described by (3), it is possible to construct gauge-invariant quantities by combining these gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This results in three distinct sets of linear combinations. In each of these sets, precisely one gauge-invariant quantity is identically zero. Given that the focus is on energy density and particle number density perturbations, the only possible set is
$\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\varepsilon_{\scriptscriptstyle(1)}-\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad n_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=n_{\scriptscriptstyle(1)}-\dfrac{\dot{n}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},$ (4a)
$\theta_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\theta_{\scriptscriptstyle(1)}-\dfrac{\dot{\theta}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}=0.$ (4b)
Local density perturbations exert no influence on the expansion of the universe, as evidenced by (4b). The transformation rule (3), illustrates that the quantities (4) are indeed gauge-invariant.
In Sec. IX, it is demonstrated that in the nonrelativistic limit, where the pressure is considered to be zero, the linearized conservation laws lack physical significance and are decoupled from the linearized energy density constraint equation and momentum constraint equations. The combination of these constraint equations results in the time-independent Poisson equation of Newtonian gravitation with a source term $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$. Furthermore, the special relativistic relation $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}mc^{2}$, where $m$ is the rest mass, is recovered. In light of the aforementioned considerations, it can be concluded that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the actual physical perturbations of the energy density and the particle number density, respectively, in the context of general relativity.
The sole method for constructing gauge-invariant quantities for the energy density and particle number density perturbations is provided by the expressions (4). As these quantities also possess the correct nonrelativistic limit, it can be concluded that the gauge problem of cosmology has been resolved. This outcome is an inevitable result of the mathematical consistency of general relativity. That is to say, the general theory of relativity allows for only one gauge-invariant quantity that accurately represents a physical density perturbation.
V Selection of a Reference System
The quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ are defined without the use of a coordinate system. However, to derive evolution equations for these quantities, it is first necessary to establish a system of reference. Since general relativity is covariant, and the quantities in question are gauge-invariant, it follows that the evolution equations for these quantities can be derived in any chosen coordinate system. The selection of an appropriate coordinate system for this problem is dependent on two criteria, which are outlined below.
Firstly, the interpretation of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ requires taking the nonrelativistic limit. In the context of general relativity, coordinate transformations are in general space-time transformations, specifically $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. In the linear case, space-time transformations are given by (1). In Newtonian gravitation, the space and time transformations are regarded as distinct and independent entities. Consequently, when taking the nonrelativistic limit, the relativistic space-time transformations must be automatically separated into independent space and time transformations. Secondly, it would be advantageous to have a coordinate system that would facilitate the derivation of the evolution equations. It will now be demonstrated that both of these requirements can be met by employing the same coordinate system.
In Newtonian gravitation, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous. In view of the nonrelativistic limit, it can be seen that a synchronous reference system represents the optimal choice. In these coordinates, the metric tensor $g_{\mu\nu}(t,\bm{x})$ of the flrw universes is given by
$g_{00}=1,\quad g_{0i}=0,\quad g_{ij}=-a^{2}(t)\tilde{g}_{ij}(\bm{x}).$ (5)
The scale factor of the universe is designated as $a(t)$. Given that $g_{00}=1$, coordinate time is equal to proper time. The synchronicity condition is $g_{0i}=0$, as explained in Ref. [31], Sec. 11.8 and Ref. [35], § 84. The metric tensor of the three-dimensional maximally symmetric subspaces of constant time is denoted by $\tilde{g}_{ij}$. The Killing equations, $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$, and (5) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (1) become
$\xi^{0}=\psi(\bm{x}),\quad\xi^{i}=\tilde{g}^{ik}(\bm{x})\dfrac{\partial\psi(\bm{x})}{\partial x^{k}}\int\!\!\frac{{\mathrm{d}}t}{a^{2}(t)}+\chi^{i}(\bm{x}),$ (6)
if only transformations between synchronous coordinates are allowed. In expressions (6), $\psi(\bm{x})$ and $\chi^{i}(\bm{x})$ are four arbitrary infinitesimal functions of the spatial coordinates. In Sec. IX, it will be proven that in the nonrelativistic limit, the function $\psi(\bm{x})$ is an arbitrary infinitesimal constant. This means that the transformations of space and time are independent of each other in this limit.
The second requirement will now be considered. Synchronous coordinates possess the property that the space-space components, $R_{ij}$, of the four-dimensional Ricci curvature tensor, $R_{\mu\nu}$, are partitioned into two distinct parts. One part exclusively comprises the time derivatives of the spatial metric tensor, $g_{ij}$, while the second part is the Ricci curvature tensor of the three-dimensional subspaces, denoted by ${{}^{3}\!}R_{ij}$. This is demonstrated in Ref. [35], § 97. As the perturbations of the spatial metric tensor and the spatial Ricci tensor are also tensors (see Sec. VI.2), the covariant decomposition theorems in Refs. [34, 36, 3] may be employed to decompose these perturbed tensors into their scalar, vector, and tensor parts. This decomposition permits the system of linearized Einstein equations and conservation laws to be decomposed into three independent systems, each of which describes the evolution of scalar, vector, and tensor perturbations. As will be demonstrated in Sec. VII, only scalar perturbations are associated with density perturbations. Accordingly, it is sufficient to consider the system for scalar perturbations (37). The aforementioned system is more straightforward to utilize than the original set of equations (12), thus facilitating the derivation of the evolution equations for density fluctuations in synchronous coordinates in comparison to any other coordinate system. In view of the fact that synchronous coordinates are likewise compatible with Newton’s theory of gravitation, they will be employed in the derivation of the primary result (60).
VI Einstein Equations and Conservation Laws
In the preceding section, it was concluded that synchronous coordinates are the optimal choice for addressing the problem at hand. In this section, the background Einstein equations and conservation laws, along with their linearized counterparts, will be expressed in this particular coordinate system.
The background energy density constraint equation and its linearized counterpart are expressed in a manner consistent with the contracted Bianchi identities. These identities state that the Riemann curvature tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$. This can be expressed in an equivalent manner as the four-divergence of the Einstein tensor, i.e., ${G^{\mu\nu}{}_{;\nu}=0}$. This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of the novel perturbation theory. For further details see Ref. [31], Secs. 6.8 and 7.5 and Ref. [35], § 92 and § 95.
It is understood from thermodynamic principles that the energy density, $\varepsilon$, and the pressure, $p$, are functions of both the particle number density, $n$, and the temperature, $T$:
$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (7)
In order to facilitate the calculations, it is advantageous to eliminate the temperature from the given equations of state. This yields the equation of state for the pressure:
$p=p(\varepsilon,n),$ (8)
which will be utilized in the subsequent analysis.
VI.1 Background Equations
The system of background Einstein equations and conservation laws for closed, flat, and open flrw universes filled with a perfect fluid with an energy-momentum tensor as defined by
$T^{\mu\nu}=(\varepsilon+p)u^{\mu}u^{\nu}-pg^{\mu\nu},\quad p=p(n,\varepsilon),$ (9)
is represented by the following equations:
$3H^{2}$ $=\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}+\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4}$ (10a)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(0)}$ $=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(0)},$ (10b)
$\dot{\varepsilon}_{\scriptscriptstyle(0)}$ $=-3H\varepsilon_{\scriptscriptstyle(0)}(1+w),$ $w$ $:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)},$ (10c)
$\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (10d)
It can be demonstrated that the $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically. The $G_{ii}$ dynamical equations are equivalent to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) given by equation (10a). Therefore, the $G_{ij}$ dynamical equations can be omitted. The cosmological constant, $\Lambda$, the gravitational constant, $G_{\text{N}}$, and the speed of light, $c$, are fundamental constants.
The overdot is used to denote differentiation with respect to $x^{0}=ct$. The Hubble function $H$ is defined by $H:=\dot{a}/a$. For flrw universes, the Hubble function is given by $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, where $\theta_{\scriptscriptstyle(0)}$ is the background value of the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$, where $u^{\mu}$ is the four-velocity $u^{\mu}:=c^{-1}U^{\mu}$, normalized to unity, $u^{\mu}u_{\mu}=1$. A semicolon is used to denote covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (5). The equations (10c) and (10d) represent the energy density conservation law and the particle number conservation law, respectively. It is crucial to highlight that the variable $w$ in Eq. (10c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent the equation of state. The spatial parts of the background Riemann tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci curvature tensor, ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction, ${{}^{3}\!}R_{\scriptscriptstyle(0)}$, are given by the following expressions:
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}=\tilde{R}^{i}{}_{jkl}=K(\delta^{i}{}_{k}\tilde{g}_{jl}-\delta^{i}{}_{l}\tilde{g}_{jk}),$ (11a)
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}=-\dfrac{2K}{a^{2}}\delta^{i}{}_{j},\quad{{}^{3}\!}R_{\scriptscriptstyle(0)}=-\dfrac{6K}{a^{2}},$ (11b)
where ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ is the spatial curvature. The value of $K$ determines the nature of the flrw universe. The universe is open for $K=-1$, flat for $K=0$, and closed for $K=+1$.
VI.2 Linearized Equations
The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is succinctly expressed as follows:
$H\dot{h}^{k}{}_{k}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\kappa\varepsilon_{\scriptscriptstyle(1)},$ (12a)
$\dot{h}^{k}{}_{k\mid i}-\dot{h}^{k}{}_{i\mid k}=2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i},$ (12b)
$\ddot{h}^{i}{}_{j}+3H\dot{h}^{i}{}_{j}\,+$
$\qquad\delta^{i}{}_{j}H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}=-\kappa\delta^{i}{}_{j}(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}),$ (12c)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})+\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)}=0,$ (12d)
$\frac{1}{c}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}\right)\,-$
$\qquad g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}=0,$ (12e)
$\dot{n}_{\scriptscriptstyle(1)}+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}=0.$ (12f)
In these equations, the following notation is employed for the perturbed metric tensor: $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$. In light of the use of synchronous coordinates, it follows that $h_{00}=0$ and $h_{0i}=0$. In addition, the raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (5), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$. A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$ or, equivalently, with respect to $\tilde{g}_{ij}$. In the context of flrw universes, where $\Gamma^{k}_{{\scriptscriptstyle(0)}ij}=\tilde{\Gamma}^{k}{}_{ij}$, the operations of taking the time derivative and the covariant derivative are shown to commute. Equation (12a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eqs. (12b) correspond to the linearized momentum constraint equations. The linearized dynamical equations are presented in Eqs. (12c). Equations (12d) and (12e) represent the linearized energy density conservation law and the linearized energy-momentum conservation laws, respectively. Finally, the linearized particle number density conservation law is given by Eq. (12f).
As previously stated in Sec. II, the pressure perturbation, the spatial curvature perturbation, and the covariant divergence of the spatial fluid velocity play a pivotal role in a cosmological perturbation theory. In consequence, the expressions for these quantities are now derived.
VI.2.1 Pressure Perturbation
The gauge-dependent perturbation to the pressure (8) is given by
$p_{{\scriptscriptstyle(1)}}=p_{n}n_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}},$ (13a)
$p_{n}:=\left(\dfrac{\partial p}{\partial n}\right)_{\varepsilon},\quad p_{\varepsilon}:=\left(\dfrac{\partial p}{\partial\varepsilon}\right)_{n}.$ (13b)
The quantity representing the true physical perturbation, designated as $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, will be derived in Sec. XI on thermodynamics.
VI.2.2 Spatial Curvature Perturbation
Lifshitz’ expression for the perturbed connection coefficients presented in Ref. [2], Eq. (I.3) and Ref. [31], Eq. (10.9.1) is a tensor, which is given by
$\Gamma^{k}_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j}+h_{lj\mid i}-h_{ij\mid l}).$ (14)
The contracted Palatini identity as defined in Ref. [2], Eq. (I.5) and Ref. [31], Eq. (10.9.2) is given by
${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=\Gamma^{k}_{{\scriptscriptstyle(1)}ij\mid k}-\Gamma^{k}_{{\scriptscriptstyle(1)}ik\mid j}.$ (15)
By combining (14) and (15), the following expression for the perturbed spatial Ricci tensor may be derived:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j\mid k}+h_{lj\mid i\mid k}-h_{ij\mid l\mid k}-h_{lk\mid i\mid j}).$ (16)
By raising the index $i$, one arrives at the following result, employing (11b):
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}:=$ $\,(g^{ik}\,{{}^{3}\!}R_{kj})_{\scriptscriptstyle(1)}=g^{ik}_{\scriptscriptstyle(0)}\,{{}^{3}\!}R_{{\scriptscriptstyle(1)}kj}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}$
$=$ $-\tfrac{1}{2}g^{il}_{\scriptscriptstyle(0)}(h^{k}{}_{l\mid j\mid k}+h^{k}{}_{j\mid l\mid k}-h^{k}{}_{k\mid l\mid j})$
$+\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}h^{i}{}_{j\mid k\mid l}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}.$ (17)
By employing $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
${{}^{3}\!}R_{\scriptscriptstyle(1)}:={{}^{3}\!}R^{k}_{{\scriptscriptstyle(1)}k}=g_{\scriptscriptstyle(0)}^{ij}(h^{k}{}_{k\mid i\mid j}-h^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{k}{}_{k}.$ (18)
In Sec. VII it will be demonstrated that the expression (18) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$, or the spatial curvature perturbation, induced by local density perturbations. The spatial curvature perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is incorporated into the evolution equations (60), as shown in Eq. (128).
VI.2.3 Covariant Divergence of the Spatial Fluid Velocity
In the background flrw universe, the fluid four-velocity is given by $u^{\mu}_{\scriptscriptstyle(0)}=\delta^{\mu}{}_{0}$, so that the perturbation $\theta_{\scriptscriptstyle(1)}$ to the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$ is
$\theta_{\scriptscriptstyle(1)}=\vartheta_{\scriptscriptstyle(1)}-\tfrac{1}{2}\dot{h}^{k}{}_{k},\quad\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k},$ (19)
where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the spatial part $u^{i}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity $u^{\mu}_{\scriptscriptstyle(1)}$. In deriving (19), the following identity is employed: $(u^{k}{}_{|k})_{{\scriptscriptstyle(1)}}=u^{k}_{{\scriptscriptstyle(1)}|k}$, which follows from $u^{i}_{{\scriptscriptstyle(0)}}=0$. The covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the evolution equations (60), as shown in Eq. (133).
VII Decomposition of Spatial Tensors
The following important theorem is used to decompose the linearized Einstein equations and conservation laws into three independent systems. In Refs. [34, 36, 3], it is demonstrated that any symmetric spatial tensor ${\cal{T}}_{ij}$ of rank two can be unambiguously decomposed in a covariant manner into three irreducible components:
${\cal{T}}^{i}{}_{j}={\cal{T}}^{i}_{\parallel\;j}+{\cal{T}}^{i}_{\perp\;j}+{\cal{T}}^{i}_{\ast\;j}.$ (20)
The constituents possess the following properties:
${\cal{T}}^{k}_{\perp\;\,k}=0,\quad{\cal{T}}^{k}_{\ast\;\;k}=0,\quad{\cal{T}}^{k}_{\ast\;\;i\mid k}=0,\quad{\cal{T}}^{i}_{\parallel\;j}=\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j},$ (21)
where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials. Consequently, the component $h^{i}_{\parallel\,j}$ can be expressed in terms of the two independent potentials as follows:
$h^{i}_{\parallel\,j}=\frac{2}{c^{2}}\bigl{(}\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j}\bigr{)}.$ (22)
The factor $2/c^{2}$ is included in view of the nonrelativistic limit. Evidently, from (21), it can be inferred that an expression analogous to (22) exists for the tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\parallel\,j}$. In the context of cosmological perturbation theory, only the contraction ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ of the latter tensor is required. Upon substituting (22) into (18), the following result is obtained:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}=\dfrac{2}{c^{2}}\Bigl{[}2\phi^{|i}{}_{|i}$ $+\zeta^{|k}{}_{|k}{}^{|i}{}_{|i}-\zeta^{|k|i}{}_{|k|i}$
$+\tfrac{1}{3}\,{{}^{3}\!}R_{{\scriptscriptstyle(0)}}(3\phi+\zeta^{|k}{}_{|k})\Bigr{]},$ (23)
and for the perturbed expansion (19) it is found
$\theta_{{\scriptscriptstyle(1)}}=\vartheta_{{\scriptscriptstyle(1)}}-\dfrac{1}{c^{2}}\bigl{(}3\dot{\phi}+\dot{\zeta}^{|k}{}_{|k}\bigr{)}.$ (24)
As evidenced by (VII), the perturbed Ricci scalar ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}$ encompasses two distinct potentials $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$. In the nonrelativistic limit, as will be discussed in detail in Sec. IX, the potential $\phi$ becomes independent of time and equal to the Newtonian potential, while the potential $\zeta$ becomes inconsequential.
The spatial component $\bm{u}_{\scriptscriptstyle(1)}$ of the perturbed fluid four-velocity can be decomposed as follows
$\bm{u}_{{\scriptscriptstyle(1)}}=\bm{u}_{{\scriptscriptstyle(1)}\parallel}+\bm{u}_{{\scriptscriptstyle(1)}\perp},$ (25)
where the components $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ and $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ have the properties
~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (26)
In these expressions, the generalized vector differential operator, denoted by $\bm{\tilde{\nabla}}$, is defined by the relation ~${\tilde{\nabla}_{i}v_{k}:=v_{k\mid i}}$. The expressions (25)–(26) are referred to as the Helmholtz decomposition. The following section will demonstrate that the decomposition of ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}$ and $h^{i}{}_{j}$ as given by (20) and the properties (21) are consistent with the the Riemann tensor (11a) and the momentum constraint equations (12b).
It is important to note that the velocity vector $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is not uniquely determined by the system of equations (12). Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it is possible to supplement it with the gradient of an arbitrary function. Given that the system of equations (12) is invariant under the gauge transformation (1), where the gauge function $\xi^{\mu}(t,\bm{x})$ is given by (6), it follows that
$u^{\prime}_{{\scriptscriptstyle(1)}\parallel i}=u_{{\scriptscriptstyle(1)}\parallel i}+\psi_{|i},$ (27)
is also a solution of the system (12). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
$\hat{u}_{{\scriptscriptstyle(1)}\parallel i}=\psi_{|i},\quad\hat{u}^{i}_{{\scriptscriptstyle(1)}\parallel}=g_{{\scriptscriptstyle(0)}}^{ik}\psi_{|k}=-\dfrac{1}{a^{2}}\tilde{g}^{ik}\psi_{|k}.$ (28)
This result is of particular significance in the context of the derivation of the nonrelativistic limit, as presented in Sec. IX.
VII.1 Decomposition of the Linearized Equations
The system of equations (12) is split into three independent systems by using the decomposition (20) with the properties (21) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as (25)–(26). The solutions to these systems are conventionally designated as tensor perturbations $\ast$, vector perturbations $\perp$, and scalar perturbations $\parallel$.
In the following three subsections, it is demonstrated that scalar perturbations are the only ones associated with density perturbations $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
VII.1.1 Tensor Perturbations
From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$. This, in combination with Eq. (12a), yields ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$. Eqs. (12b) yield the result that ${\bm{u}_{\scriptscriptstyle(1)}=\bm{0}}$. This, in turn, implies, with (19) that $\theta_{{\scriptscriptstyle(1)}}=0$. From (4a), it follows that ${\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0}$. Given that $\theta_{{\scriptscriptstyle(1)}}=0$, Eqs. (10d) and (12f) are identical. This implies that $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$. In consequence, the evolution equations for tensor perturbations are as follows:
$\ddot{h}^{i}_{\ast\,j}+3H\dot{h}^{i}_{\ast\,j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\ast j}=0.$ (29)
Given the form of these equations, tensor perturbations are typically referred to as gravitational waves.
VII.1.2 Vector Perturbations
From the expressions (18) and (21), it can be deduced that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$. This implies that
$h^{kl}_{\perp\mid k\mid l}=0.$ (30)
Moreover, it is found from Eq. (12a), that ${\varepsilon_{{\scriptscriptstyle(1)}}=0}$. Raising the index $i$ of Eqs. (12b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\mid j}.$ (31)
The combination of (30) and (31) results in ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=0$. This indicates that only the component $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ remains. From (19) one finds that $\theta_{{\scriptscriptstyle(1)}}=0$. This, combined with $\varepsilon_{{\scriptscriptstyle(1)}}=0$, yields $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. From $\theta_{{\scriptscriptstyle(1)}}=0$ one finds that Eq. (12f) is identical to the background equation (10d), implying that $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies with (4a) that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (12c) or (12d) is that $p_{\scriptscriptstyle(1)}=0$. Therefore, the system of equations for vector perturbations is given by
$\!\dot{h}^{k}_{\perp i\mid k}+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}\perp i}=0,$ (32a)
$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
$\!\frac{1}{c}\frac{{\text{d}}}{{\text{d}}t}\Bigl{(}\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}\Bigr{)}+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}=0.$ (32c)
According to (26), vector perturbations are also referred to as rotational perturbations.
VII.1.3 Scalar Perturbations
In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$. By taking the covariant derivative of (12b) with respect to the index $j$ and subsequently substituting the expression (22), the following result is obtained:
$2\dot{\phi}_{\mid i\mid j}+\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}=\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i\mid j}.$ (33)
Upon interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
$\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}=-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}),$ (34)
where it is used that $\dot{\phi}_{\mid i\mid j}=\dot{\phi}_{\mid j\mid i}$ and $\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}=\dot{\zeta}^{\mid k}{}_{\mid k\mid j\mid i}$. By rearranging the covariant derivatives, Eqs. (34) can be transformed in the following form:
$(\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid j\mid k})-(\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i\mid k})\,+$
$\qquad(\dot{\zeta}^{\mid k}{}_{\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i})_{\mid k}\,=$
$\qquad-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}).$ (35)
The expressions for the commutator of second-order covariant derivatives (see Ref. [31], Sec. 6.5) are given by
$A^{i}{}_{j\mid p\mid q}-A^{i}{}_{j\mid q\mid p}=A^{i}{}_{k}\,{{}^{3}\!}R^{k}_{{\scriptscriptstyle(0)}jpq}-A^{k}{}_{j}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq},$ (36a)
$B^{i}{}_{\mid p\mid q}-B^{i}{}_{\mid q\mid p}=B^{k}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq}.$ (36b)
Upon substituting the background Riemann tensor (11a), it is found that the left-hand sides of Eqs. (35) vanished identically. This implies that ~×$\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains. Since ~$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
VIII Evolution Equations for Scalar Perturbations
As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (10) and the equations that govern the evolution of scalar perturbations. In this section, the findings of the preceding section are utilized to reformulate the system (12) into a novel set of equations that exclusively describe the evolution of scalar perturbations. Given the subsequent focus on scalar perturbations, the subscript $\parallel$ is omitted. The evolution equations for scalar perturbations are as follows:
$2H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+\kappa\varepsilon_{\scriptscriptstyle(1)},$ (37a)
$\!\!{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$
$\qquad\;\,+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}-\tfrac{2}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)}),$ (37b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}=-3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})-\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)},$ (37c)
$\dot{\vartheta}_{\scriptscriptstyle(1)}=-H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}-\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}p_{\scriptscriptstyle(1)}}{a^{2}},$ (37d)
$\dot{n}_{\scriptscriptstyle(1)}=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (37e)
The symbol ~$\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
~$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
The parameter $\beta$ is defined as follows:
$\beta^{2}:=\dfrac{\dot{p}_{{\scriptscriptstyle(0)}}}{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}.$ (39)
By employing the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (10c) and (10d), the following result is obtained:
$\beta^{2}=p_{\varepsilon}+\dfrac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (40)
where $p_{\varepsilon}$ and $p_{n}$ are given by  (13b).
The derivation of Eqs. (37) will now be conducted. To initiate this process, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq. (12a) using the expression (19), which results in the algebraic energy density constraint equation (37a).
Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equations (12b) and the momentum conservation laws (12e) by employing the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$, (19), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$. By multiplying both sides of Eqs. (12b) by $g^{ij}_{\scriptscriptstyle(0)}$ and taking the covariant derivative with respect to the index $j$, one finds
$g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})=2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}.$ (41)
The left-hand side of Eq. (41) will appear as a component of the time derivative of the local spatial curvature perturbation, ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Indeed, differentiating (18) with respect to time, and using the relation $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$ and (10b), yields the following result:
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\dot{h}^{k}{}_{k}.$ (42)
By combining (41) and (42) and using (19) to eliminate $\dot{h}^{k}{}_{k}$, one obtains Eq. (37b). In consequence, the three $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equations (12b) have been reformulated in the form of a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (18).
The momentum conservation laws (12e) are rewritten by performing the differentiation with respect to time and using (10c), then dividing by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$. The result is as follows:
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}+H(2-3w)u^{i}_{{\scriptscriptstyle(1)}}+\dfrac{\dot{w}}{1+w}u^{i}_{{\scriptscriptstyle(1)}}-\dfrac{g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}}{\varepsilon_{{\scriptscriptstyle(0)}}(1+w)}=0.$ (43)
This equation can be simplified by eliminating the time derivative of $w$. From the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and the energy conservation law (10c) one obtains
$\dot{w}=3H(1+w)(w-\beta^{2}).$ (44)
Upon substituting this expression into Eq. (43), it can be seen that the momentum conservation laws can be reformulated as
$\dot{u}^{i}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})u^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}=0.$ (45)
By taking the covariant divergence of (45) with respect to the background metric tensor (5), and employing (19) and (38), one obtains Eq. (37d).
As demonstrated in Appendix B, the dynamical equations (12c) are superfluous since the system (37) comprises the constraint equations and conservation laws, and thus its solution automatically satisfies the dynamical equations. Consequently, the derivation of the equations (37) has been successfully completed.
In Appendix A, the primary result, namely the evolution equations (60) for the density fluctuations $\delta_{\varepsilon}:=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\delta_{n}:=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}/n_{{\scriptscriptstyle(0)}}$, will be derived using the systems (10) and (37). Prior to proceeding, it is crucial to demonstrate that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent the true physical energy density and particle number density perturbations, respectively. This necessitates the nonrelativistic limit being employed for the interpretation of these quantities.
IX Nonrelativistic Limit
The standard nonrelativistic limit is defined by four fundamental criteria (see, e.g., Ref. [37], Sec. 4.1):
  1. 1.
    The subspaces of constant time are flat.
  2. 2.
    When the pressure is reduced to zero, there are no local fluid flows.
  3. 3.
    Relativistic gauge transformations are reduced to Newtonian gauge transformations.
  4. 4.
    The gravitational field of a density perturbation is required to be constant.
The third requirement is of considerable importance, as it permits the selection of a coordinate system in Newtonian gravitation. This fact is not referenced in the existing literature. The fourth requirement is a direct consequence of the first and second requirement.
In the case of flat three-spaces, it is evident that ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$. In a flat flrw universe, covariant derivatives are equivalent to ordinary derivatives, thereby simplifying the expression (VII) to the following form:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}}=\dfrac{4}{c^{2}}\phi^{\mid k}{}_{\mid k}=-\dfrac{4}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}.$ (46)
The symbol $\nabla^{2}$ represents the conventional Laplace operator. Upon substituting the expression (46) into the perturbation equations (37), and employing the relation $H:=\dot{a}/a$, the following result is obtained:
$-H\vartheta_{\scriptscriptstyle(1)}$ $=-\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)},\quad\kappa=8\pi G_{\text{N}}/c^{4},$ (47a)
$\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\dot{\phi}}{a^{2}}$ $=-\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)},\quad w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}},$ (47b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $=-3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})-\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)},$ (47c)
$\dot{\vartheta}_{\scriptscriptstyle(1)}$ $=-H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}-\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\nabla^{2}p_{\scriptscriptstyle(1)}}{a^{2}},$ (47d)
$\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},\quad\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (47e)
In order to derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq. (37a), it is first necessary to eliminate the function $\varepsilon_{{\scriptscriptstyle(1)}}$ using the expression (4a). The result is
$2H(\theta_{{\scriptscriptstyle(1)}}-\vartheta_{{\scriptscriptstyle(1)}})=-\dfrac{2}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\kappa\left(\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\dfrac{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}}\right).$ (48)
Subsequently, the Friedmann equation (10a) with the Ricci scalar ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ set to zero, the energy conservation law (10c), and the relation between the Hubble parameter and the expansion scalar, namely, $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, are employed to obtain the result $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$. Upon substituting this expression into (48), the desired result, Eq. (47a), is obtained. With the condition of ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, the first requirement is satisfied.
Subsequently, the second requirement will be implemented. Upon substituting $p=0$, i.e., $w=0$ and $\beta=0$, into the momentum conservation laws (45), the result is
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}=-2Hu^{i}_{{\scriptscriptstyle(1)}}.$ (49)
By lowering the index $i$ with the background metric $g_{{\scriptscriptstyle(0)}ij}$ given by (5), and using the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$, one finds
$\dot{u}_{{\scriptscriptstyle(1)}i}=0.$ (50)
These equations admit only functions of the spatial coordinates as solutions. As these equations are part of the system (12), it can be concluded that the gauge modes, $\psi_{|i}$, given by (28), are the only solutions of equations (50). Consequently, it can be deduced that
$p\rightarrow 0\quad\Rightarrow\quad u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
In the absence of pressure, local pressure gradients are nonexistent. Consequently, the emergence of local fluid flows is precluded. This implies that
$u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}(t,\bm{x})\rightarrow 0\quad\Rightarrow\quad\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})\rightarrow\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(\bm{x}).$ (52)
Therefore, in the event of a zero fluid flow, the density perturbations are observed to remain static. It can be concluded that the second requirement is fulfilled by the limits (51)–(52).
Given that the physical parts of the spatial components of the fluid four-velocity are identically zero, only the gauge modes $\psi_{|i}$ remain. It can thus be concluded with confidence that the gauge modes may be taken to be equal to zero without any loss of physical information. Taking $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}=0$ implies that $\psi$ is an arbitrary infinitesimal constant. Upon substituting $\psi=C$ into the expressions (6), the result is that the relativistic gauge transformation between synchronous coordinates reduces in the limits given in (51) to the gauge transformation of Newtonian gravitation:
$t\rightarrow t^{\prime}=t-C,\quad x^{i}\rightarrow x^{\prime i}=x^{i}-\chi^{i}(\bm{x}).$ (53)
This implies that within the Newtonian theory of gravitation, there is the possibility of shifting the time coordinate and selecting the spatial coordinates at will. In accordance with the expressions (53), the third requisite is thereby fulfilled.
The fourth and final requirement will now be discussed. Given that $u^{i}_{{\scriptscriptstyle(1)}{\textnormal{phys}}}=0$ and $\hat{u}^{i}_{{\scriptscriptstyle(1)}}=0$, it follows that $\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k}=0$. By combining this result with (52), the linearized Einstein equations and conservation laws (47) are reduced to
$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}),$ (54a)
$\nabla^{2}\dot{\phi}(t,\bm{x})$ $=0,$ (54b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $=-3H\varepsilon_{\scriptscriptstyle(1)}-\varepsilon_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},$ (54c)
$\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (54d)
The constraint equations (54a) and (54b) are combined to yield the following result:
$\nabla^{2}\phi(\bm{x})=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
The potential $\varphi(\bm{x})$ is defined as $\varphi(\bm{x}):=\phi(\bm{x})/a^{2}(t_{0})$, which results in the Poisson equation of Newtonian gravitation:
$\nabla^{2}\varphi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}}.$ (56)
With (56), the fourth and final requirement has been met.
In the nonrelativistic limit, the background equations (10) are given by
$3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad\dot{\varepsilon}_{\scriptscriptstyle(0)}=-3H\varepsilon_{\scriptscriptstyle(0)},\quad\dot{n}_{\scriptscriptstyle(0)}=-3Hn_{\scriptscriptstyle(0)}.$ (57)
Consequently, the conservation laws imply that $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$. As may be observed, Eqs. (54c)–(54d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (53) of Newtonian gravitation. This, combined with $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$ yields $\varepsilon_{{\scriptscriptstyle(1)}}=n_{{\scriptscriptstyle(1)}}mc^{2}$. Combining the latter two relations with the definitions (4a) one finds the well-known special relativistic relation between the rest mass density and energy density:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})mc^{2}.$ (58)
The equations (54c) and (54d) represent the residual form of the conservation laws (47c) and (47e) in the nonrelativistic limit. It is evident that equations Eqs. (54c) and (54d) have no physical significance. These equations are decoupled from the physical constraint equations (54a) and (54b) and thus do not form part of Newtonian gravitation. Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$, (24), is no longer relevant. In view of the aforementioned considerations, it becomes evident that solely the Poisson equation (56) and the special relativistic relation (58) have survived.
Following an examination of the four points regarding the nonrelativistic limit, the following conclusion is reached. It has been demonstrated that the system of linearized Einstein equations and conservation laws for scalar perturbations (37), when combined with the gauge-invariant quantities (4a), reduces in the nonrelativistic limit to the Poisson equation (56) of Newtonian gravitation and the special relativistic relation (58). This indicates that the gauge-invariant quantities $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, defined by expressions (4a), represent the local perturbations of the energy density and the particle number density, respectively, in Newtonian gravitation. Therefore, it can be stated that $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ represent the actual local density perturbations of the energy density and the particle number density in general relativity. In conclusion, since the gauge-invariant perturbation of the expansion, $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ (4b), is identically zero, it can be stated that density perturbations have no influence on the expansion of the universe.
The subsequent section of this study will proceed to derive the evolutionary equations for $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$. These equations permit the investigation of the evolution of density fluctuations in closed, flat, and open flrw universes independently of the choice of coordinate system.
X Evolution Equations for Density Fluctuations
In light of the preceding analysis, it can be concluded that the quantities represented by $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$, are the real physical perturbations to the energy density and particle number density, respectively. Furthermore, it has been demonstrated that the evolution equations for scalar perturbations (37), when combined with the background equations (10), govern the evolution of the energy density perturbation and the particle number density perturbation. In Appendix A, the derivation of evolution equations for their corresponding density fluctuations will be presented. These density fluctuations are defined as follows:
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)}.$ (59)
The final result is a system of evolution equations for the density fluctuations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes. These equations are given by
$\ddot{\delta}_{\varepsilon}+b_{1}\dot{\delta}_{\varepsilon}+b_{2}\delta_{\varepsilon}$ $=b_{3}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right),$ (60a)
$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right)$ $=\frac{3Hn_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right).$ (60b)
The coefficients $b_{1}$, $b_{2}$ and $b_{3}$ are as follows:
$b_{1}=$ $\,\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{H}-2\dfrac{\dot{\beta}}{\beta}-H(2+6w+3\beta^{2})$
$+{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\dfrac{1}{3H}+\dfrac{2H(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right),$ (61a)
$b_{2}=$ $-\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)$
$+H^{2}\left(1-3w+6\beta^{2}(2+3w)\right)$
$+6H\dfrac{\dot{\beta}}{\beta}\left(w+\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
$-{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\tfrac{1}{2}w+\dfrac{H^{2}(1+6w)(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
$-\beta^{2}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right),$ (61b)
$b_{3}=$ $\,\Biggl{[}\dfrac{-18H^{2}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\Biggl{(}\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
$+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr{)}+p_{n}\Biggr{]}$
×$\times\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\left(\frac{\tilde{\nabla}^{2}}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\right).$ (61c)
In these expressions the partial derivatives of the pressure, $p_{\varepsilon}$ and $p_{n}$, are defined by (13b). The second-order partial derivatives are defined by $p_{nn}:={\partial^{2}p}/{\partial n^{2}}$ and $p_{\varepsilon n}:={\partial^{2}p}/{\partial\varepsilon\,\partial n}$.
It is essential to recognize that, with regard to the nonrelativistic limit, the aforementioned equations apply solely in the case of $p\not\equiv 0$.
XI Thermodynamic Quantities
The combined first and second laws of thermodynamics are expressed in terms of density fluctuations, denoted by $\delta_{\varepsilon}$ and $\delta_{n}$ and gauge-invariant expressions are provided for the local pressure and temperature perturbations.
XI.1 Thermodynamic Laws
The combined first and second laws of thermodynamics for a simple, single-species system is given by (see, for example, Ref. [38], Sec. 2.1)
${\mathrm{d}}E=T{\mathrm{d}}S-p{\mathrm{d}}V+\mu{\mathrm{d}}N,$ (62)
where $E$, $S$, and $N$ are the energy, the entropy and the number of particles of a system with volume $V$, temperature $T$, and pressure $p$. The thermal—or chemical—potential $\mu$, is the energy required to add one particle to the system. In terms of the particle number density $n=N/V$, the energy per particle $E/N=\varepsilon/n$ and the entropy per particle $s=S/N$ the law (62) can be rewritten in the form
${\mathrm{d}}\left(\dfrac{\varepsilon}{n}N\right)=T{\mathrm{d}}(sN)-p{\mathrm{d}}\left(\dfrac{N}{n}\right)+\mu{\mathrm{d}}N,$ (63)
where $\varepsilon$ is the energy density. The system is extensive, i.e., $S(\lambda E,\lambda V,\lambda N)=\lambda S(E,V,N)$, which implies that the entropy of the gas is given by $S=(E+pV-\mu N)/T$. Upon dividing this relation by $N$ the Euler relation is obtained:
$\mu=\dfrac{\varepsilon+p}{n}-Ts.$ (64)
Eliminating $\mu$ in (63) with the aid of (64) reveals that the combined first and second laws of thermodynamics  (62) can be expressed in a form that does not include $\mu$ or $N$ (see Ref. [33]):
$T{\mathrm{d}}s={\mathrm{d}}\Bigl{(}\dfrac{\varepsilon}{n}\Bigr{)}+p{\mathrm{d}}\Bigl{(}\dfrac{1}{n}\Bigr{)}.$ (65)
From the background equations (10) and the thermodynamic law (65) it can be demonstrated that $\dot{s}_{\scriptscriptstyle(0)}=0$, indicating that the expansion of the universe takes place without generating entropy. From the transformation (3) it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant. By employing (138) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (65) can be rewritten as
$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n^{2}_{\scriptscriptstyle(0)}}\left(n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\right).$ (66)
By employing the definitions (59), the following is obtained:
$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n_{\scriptscriptstyle(0)}}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (67)
Therefore, $s_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}$ represents the local perturbation to the entropy per particle.
XI.2 Temperature and Pressure Perturbations
Given that both the background temperature, denoted by $T_{{\scriptscriptstyle(0)}}$, and the pressure, represented by $p_{{\scriptscriptstyle(0)}}$, are scalars, their gauge-invariant temperature and pressure perturbations can be defined in a manner analogous to that of the expressions (4):
$T^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=T_{\scriptscriptstyle(1)}-\dfrac{\dot{T}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad p^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=p_{\scriptscriptstyle(1)}-\dfrac{\dot{p}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}.$ (68)
In order to arrive at the gauge-invariant counterpart of (13a), it is first necessary to eliminate $\varepsilon_{{\scriptscriptstyle(1)}}$ and $n_{{\scriptscriptstyle(1)}}$ from (13a) using (4a). The utilization of $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (68), results in
$p^{\textnormal{phys}}_{\scriptscriptstyle(1)}=p_{n}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (69)
Eliminating $p_{\varepsilon}$ with the aid of (40) and utilizing the definitions provided in (59) for density fluctuations results in the following expression:
$p_{\scriptscriptstyle(1)}^{\textnormal{phys}}=\beta^{2}\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}+n_{\scriptscriptstyle(0)}p_{n}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (70)
The first term represents the adiabatic component of the pressure perturbation, while the second term denotes the nonadiabatic component.
XII Application: Structure Formation in a Flat FLRW Universe
The evolution equations presented in (60), are applicable to open, flat, and closed flrw universes. Henceforth, the discussion will be focused on the flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, universe. Given that the initial density fluctuations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant, $\Lambda$, will be neglected. Consequently, the background equations, as given in Eq. (10), are reduced to
$3H^{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)},\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4},$ (71a)
$\dot{\varepsilon}_{\scriptscriptstyle(0)}$ $=-3H\varepsilon_{\scriptscriptstyle(0)}(1+w),$ $\quad w$ $:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)},$ (71b)
$\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (71c)
By employing the Friedmann equation (71a), the coefficients (61) of Eq. (60a) result in
$b_{1}$ $=H(1-3w-3\beta^{2})-2\dfrac{\dot{\beta}}{\beta},$ (72a)
$b_{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl{(}2\beta^{2}(2+3w)-\tfrac{1}{6}(1+18w+9w^{2})\Bigr{)}$
$+2H\dfrac{\dot{\beta}}{\beta}(1+3w)-\beta^{2}\dfrac{\nabla^{2}}{a^{2}},$ (72b)
$b_{3}$ $=\Biggl{[}\dfrac{-2}{1+w}\Biggl{(}\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
$+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr{)}+p_{n}\Biggr{]}\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\dfrac{\nabla^{2}}{a^{2}},$ (72c)
where $\nabla^{2}$ represents the conventional Laplace operator.
XII.1 Era before Decoupling of Matter and Radiation
In this era the primordial fluid is a mixture of radiation and matter, wherein the contribution of matter to the pressure is negligible. Consequently, the equations of state (see, for example, Ref. [8], Eq. (5.49), Ref. [10], Eq. (1.78), and Ref. [17], § V-1) are given by:
$\varepsilon=a_{\text{B}}T_{\gamma}^{4}+nmc^{2},\quad p=\tfrac{1}{3}a_{\text{B}}T_{\gamma}^{4}.$ (73)
The black body constant is represented by $a_{\text{B}}$, the radiation temperature is denoted by $T_{\gamma}$ and the particle number density of ordinary matter or cdm is given by $n$. Upon eliminating $T_{\gamma}$, the following result is obtained [see Ref. [31], Eq. (2.10.27)]:
$p=\tfrac{1}{3}(\varepsilon-nmc^{2}).$ (74)
By making use of (13b), one has
$p_{n}=-\tfrac{1}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{1}{3}.$ (75)
For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), the following expressions are derived:
$w=\frac{\tfrac{1}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}},\quad\beta^{2}=\frac{1}{3}-\frac{\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{\tfrac{4}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}}.$ (76)
The ensuing subsections provide a concise overview of the radiation-dominated and matter-dominated eras preceding decoupling.
XII.1.1 Radiation-dominated Era
The universe was radiation-dominated when
$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}\gg n_{{\scriptscriptstyle(0)}}mc^{2}.$ (77)
In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$ and $p_{\varepsilon}\approx w$. Upon substituting the aforementioned values and expressions (75) into the coefficients (72), the following result is obtained:
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}$
$\qquad\qquad\qquad\;\,\approx\dfrac{-\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\dfrac{\nabla^{2}}{a^{2}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)},$ (78a)
$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}\approx-\tfrac{3}{4}H\dfrac{n_{{\scriptscriptstyle(0)}}mc^{2}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}\bigl{(}\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\bigr{)}.$ (78b)
From (77), it can be deduced that the pressure term on the left-hand side of equation (78a) is, in absolute value, significantly larger than the source term. Consequently, the source term may be neglected. In the radiation-dominated phase, the system of equations (60) can be expressed as follows:
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}\approx 0,$ (79a)
$\left|\delta_{n}-\tfrac{3}{4}\delta_{\varepsilon}\right|\rightarrow 0.$ (79b)
The limit in Eq. (79b) is a consequence of the fact that the coefficient in Eq. (78b) is negative, since $p_{n}<0$.
In order to solve Eq. (79a), it is necessary to employ the solutions of the background equations (71). These solutions are provided by:
$H=\tfrac{1}{2}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}(ct)^{-2},\quad a\propto t^{1/2}.$ (80)
The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{0}\geq 1$. This definition implies that
$\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{0}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl{(}2H(t_{0})\bigr{)}^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (81)
By employing the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (81), it can be shown that the equation (79a) can be rewritten as
$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0.$ (82)
In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{r}}$ is defined as follows:
$\mu_{\text{r}}:=\dfrac{2\pi}{\lambda_{0}}\dfrac{1}{H(t_{0})}\dfrac{1}{\sqrt{3}}.$ (83)
In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a fluctuation at $t=t_{0}$, with the magnitude of the fluctuation given by $\lvert\bm{q}\rvert=2\pi/\lambda_{0}$.
In order to solve Eq. (82), it is first necessary to replace the variable $\tau$ with the variable defined by $x:=\mu_{\text{r}}\sqrt{\tau}$. Upon transforming back to $\tau$, the general solution of Eq. (82) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained, namely
$\delta_{\varepsilon}(\tau,\bm{q})=\bigl{[}A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr{]}\sqrt{\tau},$ (84)
It can be inferred from this solution that, during the radiation-dominated era of the universe, density fluctuations oscillated with an amplitude that grew in a manner proportional to the square root of time.
In the case of large-scale fluctuations, that is, ${\lambda_{0}\rightarrow\infty}$, which leads to $\mu_{\text{r}}\rightarrow 0$, (83), the general solution of Eq. (82) is given by
$\delta_{\varepsilon}(\tau)=-\bigl{[}\delta_{\varepsilon}(1)-2\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau+2\bigl{[}\delta_{\varepsilon}(1)-\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau^{1/2}.$ (85)
This solution, with the exception of the precise factors of proportionality, has been derived by a significant number of authors, as evidenced in Refs. [2, 39, 40, 41, 42].
A comparison of the solutions (85) and (155a) reveals that the solution ${\propto\tau}$ is a consequence of the homogeneous part of equation (154a). The solution $\propto\tau^{1/2}$ is the particular solution of the inhomogeneous equation. This solution can only be obtained if the covariant divergence, $\vartheta_{{\scriptscriptstyle(1)}}$, is taken into account, as demonstrated in Appendix C.1. The minus sign preceding the solution $\propto\tau$ is a direct consequence of the fact that ${\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0}$, as follows from (155b). The presence of the gauge function $\psi$ precludes the derivation of the solution (85) from (154). As demonstrated in Appendix A, $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into the left-hand side of (78a), as evidenced by the derivation of (60a).
XII.1.2 Matter-dominated Era
The radiation-dominated phase came to an end and the matter-dominated phase commenced at $z(t_{\text{eq}})=3387$, as indicated in Table 1, well before decoupling [43], when
$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}=n_{{\scriptscriptstyle(0)}}mc^{2}.$ (86)
Given that $p_{n}<0$, it can be shown from Eq. (60b) that
$\left|\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right|\rightarrow 0.$ (87)
It can thus be seen that fluctuations in the energy density are coupled to those in the particle number density. In light of the fact that the evolution equations (60) are general relativistic, it follows that the weak equivalence principle is valid. This implies that the evolution of fluctuations in the particle number density is independent of the composition of matter. Therefore, if cdm consists of particles that interact via gravitation, the fluctuations in the cdm are also linked to the fluctuations in the energy density. This implies that it is impossible for cdm to have contracted before decoupling. Consequently, cdm cannot have initiated the formation of structure after decoupling.
XII.2 Era after Decoupling of Matter and Radiation
Once protons and electrons have combined to form hydrogen, the radiation pressure will be negligible, and the equations of state will be those of a nonrelativistic monatomic perfect gas with three degrees of freedom. These are given by the equations of state [see Ref. [31], Eqs. (15.8.20)–(15.8.21), and Ref. [33], Eq. (13)]:
$\varepsilon(n,T)=nmc^{2}+\tfrac{3}{2}nk_{\text{B}}T,\quad p(n,T)=nk_{\text{B}}T.$ (88)
In these expressions, the symbol $k_{\text{B}}$ represents Boltzmann’s constant, the quantity $m$ denotes the proton rest mass, and the temperature of the matter is indicated by the symbol $T$. The rest mass energy density is $nmc^{2}$ and the kinetic energy density is $\tfrac{3}{2}nk_{\text{B}}T$. The elimination of the temperature, $T$, from the equations of state (88), yields the equation of state for the pressure [see Ref. [31], Eq. (2.10.27)]:
$p(n,\varepsilon)=\tfrac{2}{3}(\varepsilon-nmc^{2}).$ (89)
The partial derivatives are determined by the utilization of the expressions (13b) and are given by
$p_{n}=-\tfrac{2}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{2}{3}.$ (90)
The parameter $w$ is defined by
$w:=\dfrac{p_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}=\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}+\tfrac{3}{2}k_{\text{B}}T_{\scriptscriptstyle(0)}}\approx\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}\ll 1.$ (91)
Upon substituting (88), (90) and (91) into (40), one arrives at the well-known result [see Ref. [31], Eq. (15.8.22)]:
$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}}.$ (92)
In this expression, $v_{\text{s}}$ represents the adiabatic speed of sound. From $\beta^{2}\approx\tfrac{5}{3}w$, it is found that Eq. (44) results in $\dot{w}\approx-2Hw$. Therefore, with $H:=\dot{a}/a$, it can be concluded that $w\propto a^{-2}$. Given that $w$ is proportional to $T_{{\scriptscriptstyle(0)}}$, the well-known result (see Ref. [31], Eq. (15.5.16) with $\gamma=\tfrac{5}{3}$, where ${\gamma:=c_{\text{p}}/c_{\text{v}}}$ is the adiabatic index) is obtained:
$T_{\scriptscriptstyle(0)}\propto a^{-2}.$ (93)
In light of the relatively minor values of both $w$ and $\beta^{2}$, it is now possible to justify the equations of state (88). From the proportionality (93) and the fact that $n_{{\scriptscriptstyle(0)}}\propto a^{-3}$, as can be seen from Eq. (71c), it follows from (88) that ${p_{{\scriptscriptstyle(0)}}\propto a^{-5}}$. Given that pressure decayed rapidly and was already negligible in comparison to the energy density at the time of decoupling, it can be concluded that the unperturbed pressure had a minimal impact on structure formation. It is therefore reasonable to neglect the quantities $w$ and $\beta^{2}$ with respect to the constants of order one that appear in the momentum conservation laws (45). In doing so, one finds
$\dot{u}^{i}_{\scriptscriptstyle(1)}+2Hu^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}-\dfrac{p_{{\scriptscriptstyle(1)}|j}}{\varepsilon_{\scriptscriptstyle(0)}}=0.$ (94)
The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$. Accordingly, the evolution of density fluctuations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$. In the event that the pressure gradients are identically zero, Eqs. (94) reduce to the form of Eqs. (49) and (50) in the nonrelativistic limit. This implies that the physical component of the fluid velocity is zero, and consequently, density perturbations cease to evolve. Given that structure formation in the universe was predominantly determined by pressure gradients, it is necessary to consider the kinetic energy density and pressure in addition to the energy density. This is the reason why the particle number density must be taken into account in the equations of state. In conclusion, it can be inferred from the equations (94) that the gauge modes (28) are zero. This result demonstrates that the relativistic gauge transformations (1), where $\xi^{\mu}$ is given by (6), are equivalent to the Newtonian gauge transformations (53). This conclusion is consistent with expectations, given that the equations of state (88) are nonrelativistic.
The subsequent step is to derive the evolution equations for density perturbations. The proportionality (93) implies with (92) that $\dot{\beta}/\beta=-H$. The system of equations (60) with coefficients (72) can now be rewritten in the following form:
$\ddot{\delta}_{\varepsilon}+3H\dot{\delta}_{\varepsilon}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{5}{6}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}=-\dfrac{2}{3}\dfrac{\nabla^{2}}{a^{2}}\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95a)
$\dfrac{1}{c}\dfrac{\mathrm{d}}{{\mathrm{d}}t}\left(\delta_{n}-\delta_{\varepsilon}\right)=-2H\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95b)
In this calculation, the approximation is made that $w$ and $\beta^{2}$ are negligible with respect to the constants of order one.
Given that the term $(\delta_{n}-\delta_{\varepsilon})$ occurs within the source term of Eq. (95a), it is necessary to first solve for the solution to Eq. (95b). By making the substitution $H:=\dot{a}/a$, the following result is obtained:
$\delta_{n}-\delta_{\varepsilon}\propto a^{-2}.$ (96)
In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of density fluctuations. It can be deduced from $\varepsilon=\varepsilon(n,T)$ that
$\dot{\varepsilon}_{\scriptscriptstyle(0)}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}\dot{n}_{\scriptscriptstyle(0)}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}\dot{T}_{\scriptscriptstyle(0)},$ (97a)
$\varepsilon_{\scriptscriptstyle(1)}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}.$ (97b)
By multiplying (97a) by $\theta_{\scriptscriptstyle(1)}/\dot{\theta}_{\scriptscriptstyle(0)}$ and subtracting the result from (97b), one obtains the following result:
$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}.$ (98)
This result is based on the use of expressions (4a) and (68). By employing the expression for $\varepsilon$ in (88) to eliminate the partial derivatives, one obtains the perturbed equation of state:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}mc^{2}+\tfrac{3}{2}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}k_{\text{B}}T_{\scriptscriptstyle(0)}+\tfrac{3}{2}n_{\scriptscriptstyle(0)}k_{\text{B}}T^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (99)
By dividing expression (99) by $\varepsilon_{\scriptscriptstyle(0)}$, and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, the perturbed equation of state for the energy density expressed in fluctuations (59) is obtained:
$\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})=-\tfrac{3}{2}w(t)\delta_{T}(t,\bm{x}).$ (100)
In this context, the quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$. The solution to Eq. (95b) is obtained by combining (96) and (100) and using the approximate value of $w$, (91), which implies that $w\propto a^{-2}$. The result is as follows:
$\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x}).$ (101)
The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the temperature fluctuation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation. Since $w\ll 1$, (101) is in alignment with (87).
According to (67), (91) and (101) the entropy per particle is as follows:
$s^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})\approx\tfrac{3}{2}k_{\text{B}}\delta_{T}(t_{\text{dec}},\bm{x}).$ (102)
Therefore, it can be concluded that the quantity $\delta_{T}$ is random and that density fluctuations are nonadiabatic if $\delta_{T}$ is nonzero.
Upon substituting (90), (92), (101), and the approximation ${w\ll 1}$, the local pressure fluctuation, defined by $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$, can be calculated from expression (70). One obtains
$\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x}).$ (103)
In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and nonadiabatic pressure fluctuations, respectively.
The solution to Eq. (95b) has been determined. This allows for the reformulation of the second-order equation (95a) to facilitate analysis of the evolution of density fluctuations. To that end, the solutions to the background equations (71) are required. Given that the pressure with respect to the rest mass energy density can be disregarded ($w\ll 1$), the following solutions are obtained:
$H=\tfrac{2}{3}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}(ct)^{-2},\quad a\propto t^{2/3}.$ (104)
The dimensionless time, denoted by the symbol $\tau$, is defined as follows: $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
$\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{\text{dec}}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl{(}\tfrac{3}{2}H(t_{\text{dec}})\bigr{)}^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (105)
By employing Eqs. (71a) and (101), as well as the expressions (92), (104) (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq. (95a) can be expressed in the following manner:
$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}).$ (106)
In this equation, the prime denotes differentiation with respect to the variable $\tau$. The parameter $\mu_{\text{m}}$ is defined as follows:
$\mu_{\text{m}}:=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{H(t_{\text{dec}})}\dfrac{v_{\text{s}}(t_{\text{dec}})}{c}.$ (107)
In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a fluctuation at $t=t_{\text{dec}}$. The magnitude of the fluctuation is given by $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$. Furthermore, one has $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$, as stated in expression (92).
Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of density fluctuations:
$\delta^{\prime\prime}_{\varepsilon}+\frac{2}{\tau}\delta^{\prime}_{\varepsilon}+\dfrac{4}{15}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}\delta_{p}-\dfrac{10}{9\tau^{2}}\delta_{\varepsilon}=0,\quad\tau:=\dfrac{t}{t_{\text{dec}}}\geq 1.$ (108)
The second term in this equation represents the expansion, the third term is the pressure term, where $\delta_{p}$ is given by (103), and the fourth term represents gravitation.
In the limit of large scales, that is, $\lambda_{\text{dec}}\rightarrow\infty$, which leads to $\mu_{\text{m}}\rightarrow 0$, the solution of Eq. (108) is
$\delta_{\varepsilon}(\tau)=\bigl{[}\tfrac{5}{7}\delta_{\varepsilon}(1)+\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau^{2/3}+\bigl{[}\tfrac{2}{7}\delta_{\varepsilon}(1)-\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr{]}\tau^{-5/3}.$ (109)
Therefore, pressure fluctuations $\delta_{p}$ played a negligible role in the evolution of the large-scale fluctuations. In other words, these fluctuations were almost adiabatic and evolved solely under the influence of gravitation and expansion.
A comparison of the solutions (109) and (157a) reveals that the solution ${\propto\tau^{2/3}}$ is a consequence of the homogeneous part of Eq. (156a). The solution $\propto\tau^{-5/3}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.2. In contrast to the standard equation (150), the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq. (95a). The presence of the gauge constant $C$, precludes the derivation of solution (109) from the system (156).
XII.2.1 Cosmological Quantities
To investigate the growth of density fluctuations, it is first necessary to express the parameter $\mu_{\text{m}}$ (107) in observable quantities.
The redshift $z(t)$ is defined by the expression
$z(t):=\dfrac{a(t_{\text{p}})}{a(t)}-1,$ (110)
where $a(t_{\text{p}})$ is the present value of the scale factor. In the case of a flat flrw universe, it is permissible to take $a(t_{\text{p}})=1$. Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift. The result is
$H(t)$ $=H(t_{\text{p}})\bigl{(}z(t)+1\bigr{)}^{3/2},$ (111a)
$t$ $=t_{\text{p}}\bigl{(}z(t)+1\bigr{)}^{-3/2},$ (111b)
$T_{{\scriptscriptstyle(0)}\gamma}(t)$ $=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})\bigl{(}z(t)+1\bigr{)},$ (111c)
$n_{{\scriptscriptstyle(0)}}(t)$ $=n_{{\scriptscriptstyle(0)}}(t_{\text{p}})\bigl{(}z(t)+1\bigr{)}^{3}.$ (111d)
In deriving relation (111c), it is utilised that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$. The dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is given by
$\tau=\left(\dfrac{z(t_{\text{dec}})+1}{z(t)+1}\right)^{3/2},$ (112)
where relation (111b) with $z(t_{\text{p}})=0$ is used. Upon substituting the result (92) into expression (107) and utilizing the relations (111), the following is obtained:
$\mu_{\text{m}}=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{cH(t_{\text{p}})}\dfrac{1}{z(t_{\text{dec}})+1}\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})}{m}},$ (113)
where it is assumed that at decoupling, the matter and radiation temperatures were equal, that is, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. In this manner, the parameter $\mu_{\text{m}}$ has been expressed in observable quantities. Upon substituting the numerical values from Tables 1 and 2, one obtains the result:
$\mu_{\text{m}}=\dfrac{16.48}{\lambda_{\text{dec}}},\quad\lambda_{\text{dec}}\text{ in pc}.$ (114)
It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density fluctuation.
Table 1: Planck satellite results
$z(t_{\text{dec}})$ = 1090
$z(t_{\text{eq}})$ = 3387
$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
[…]
$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert$ $\lesssim$ $10^{-5}$
XII.2.2 Initial Values
In order to solve Eq. (108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$. Furthermore, it is imperative to ascertain the values of the local nonadiabatic random pressure fluctuation $\delta_{T}$ in expression (103).
Planck Satellite
The Planck observations [43] of the fluctuations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$. In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$. Therefore, the initial values for Eq. (108) are as follows:
$\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\lesssim 10^{-5},\quad\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0.$ (115)
As demonstrated in the calculations presented in Sec. XII.2.3, the outcome of Eq. (108) is largely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition of having a maximum value of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
Nonadiabatic Pressure Fluctuations
The values of the quantities $\delta_{\varepsilon}(t,\bm{q})$ and $\delta_{n}(t,\bm{q})$ are related to the local nonadiabatic random pressure fluctuation $\delta_{T}(t_{\text{dec}},\bm{q})$ by Eq. (101). Immediately following the decoupling, this equation reads:
$\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx-4.1\times 10^{-10}\delta_{T}(t_{\text{dec}},\bm{q}),$ (116)
where it has been assumed that at decoupling, the matter temperature was equal to the radiation temperature, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. Expressions (91) and (111c) were employed in conjunction with the values presented in Tables 1 and 2.
The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic. A notable decline was observed in the mean particle velocity, and the pressure. Furthermore, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering. Therefore, it is probable that after the transition, the initial values of the difference between fluctuations, represented by ${(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})}$, were randomly distributed among all density fluctuations. From expression (116), it can be seen that very small differences between the initial energy density and particle number density fluctuations result in relatively large positive or negative fluctuations in the nonadiabatic component of the pressure fluctuations in expression (103). Therefore, the nonadiabatic pressure fluctuations, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, are randomly (102) distributed among the multitude of density fluctuations.
Table 2: Physical constants
$m$ = $1.6726\times 10^{-27}\,\text{kg}$
pc = $3.0857\times 10^{16}\,\text{m}=3.2616\,\text{ly}$
$c$ = $2.9979\times 10^{8}\,\text{m}\,\text{s}^{-1}$
$k_{\text{B}}$ = $1.3806\times 10^{-23}\,\text{J}\,\text{K}^{-1}$
$G_{\text{N}}$ = $6.6743\times 10^{-11}\,\text{m}^{3}\,\text{kg}^{-1}\,\text{s}^{-2}$
$\text{M}_{\odot}$ = $1.9889\times 10^{30}\,\text{kg}$
$a_{\text{B}}$ = $7.5657\times 10^{-16}\,\text{J}\,\text{m}^{-3}\,\text{K}^{-4}$
XII.2.3 Structure Formation
The subsequent analysis will employ Eq. (108) to examine the evolution of density fluctuations. This equation demonstrates that the evolution of a fluctuation $\delta_{\varepsilon}(t,\bm{q})$ with an initial scale $\lambda_{\text{dec}}$, was influenced by three key factors: the expansion of the universe, fluctuations in pressure, and gravitation. Fluctuations in pressure, as indicated by (103), were found to consist of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115) and the nonadiabatic component was observed to be random.
[…]
Fig. 1 provides a clear summary of the evolution of density fluctuations. It was created using the following methodology. For each value of ${\delta_{T}=-0.01}$, $-0.05$, $-0.10$, $-0.15$, and $-0.20$, Eq. (108) is numerically solved for a large number of values for the initial fluctuation scale $\lambda_{\text{dec}}$ using the initial values (115).
[…]
$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc}.$ (117)
This value has been calculated using Eq. (111b) and Tables 1 and 2.
Figure 2: The graphs show the growth rates $\delta^{\prime}_{\varepsilon}$, with initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$, as function of the redshift $z$, or time in million of years. The initial scales $\lambda_{\text{dec}}$ of the fluctuations are measured in parsec. The evolution of density fluctuations started at $z=1090$.
[…]
$M(t_{\text{dec}})=\dfrac{4\pi}{3}\bigl{(}\tfrac{1}{2}\lambda_{\text{dec}}\bigr{)}^{3}n_{\scriptscriptstyle(0)}(t_{\text{dec}})m.$ (118)
It is first necessary to determine the particle number density at the time of decoupling denoted by $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, in order to proceed. This quantity is derived from the Friedmann equation (71a). For $w\ll 1$ and at time $t_{\text{p}}$ this equation is given by:
$3H^{2}(t_{\text{p}})=\kappa n_{{\scriptscriptstyle(0)}}(t_{\text{p}})mc^{2},\quad\kappa=8\pi G_{\text{N}}/c^{4}.$ (119)
By employing the relation (111d), one obtains
$n_{\scriptscriptstyle(0)}(t_{\text{dec}})=\frac{3\bigl{[}cH(t_{\text{p}})\bigr{]}^{2}\bigl{[}z(t_{\text{dec}})+1\bigr{]}^{3}}{8\pi mG_{\text{N}}}.$ (120)
[…]
$\lambda_{\text{Jeans}}(t_{\text{dec}}):=\lambda_{\text{Jeans}}a(t_{\text{dec}})\approx 6.4\,\text{pc}.$ (121)
Accordingly, the Jeans mass at decoupling is thus given by
$M_{\text{Jeans}}(t_{\text{dec}})\approx 2.2\times 10^{4}\,\text{M}_{\odot}.$ (122)
As evidenced by the preceding analysis, fluctuations in density, with initial values of $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$, result in the formation of structures in a flat flrw universe, provided that pressure fluctuations are taken into account. This marks the conclusion of the analysis concerning the evolution of density fluctuations in a flat flrw universe.
XIII Summary and Conclusion
This paper addresses four issues in the field of cosmological perturbation theory for Friedmann-Lemaître-Robertson-Walker (flrw) universes. The first and most significant issue concerns the gauge problem. It is well-established that the linearized Einstein equations and conservation laws yield solutions that are both physical and nonphysical. Consequently, the imposition of physical initial conditions does not result in solutions with physical significance. To resolve this issue, it is crucial to consider gauge-invariant quantities, that is, quantities that are independent of the choice of a coordinate system. It has been demonstrated that perturbations in the energy density and the particle number density can be constructed in a unique manner. In the nonrelativistic limit, it is demonstrated that these quantities represent the actual physical perturbations of the energy density and particle number density in the general theory of relativity. Similarly, as has been done for density perturbations, gauge-invariant perturbations in temperature, pressure, and entropy can be constructed. This addresses the gauge issue in the field of cosmology.
The second issue is that of the growth of minor density fluctuations subsequent to the decoupling of matter and radiation. In a fluid with zero pressure, density fluctuations do not evolve. To address this issue, it is necessary to incorporate the particle number density in addition to the energy density in the linearized Einstein equations and conservation laws. This allows for the kinetic energy density and pressure to be accounted for, resulting in the emergence of local pressure gradients. These gradients give rise to local fluid flows that are instrumental in the formation of structures after decoupling.
A third result is the novel perturbation theory (60), which provides a clear and unambiguous description of the evolution of fluctuations in the particle number density the energy density, as well as their interaction in a perfect fluid where the pressure depends on both the particle number density and the energy density. The aforementioned equations are not only applicable to a flat flrw universe, but also to open and closed universes. Furthermore, the cosmological constant need not be zero. It has been demonstrated that Eqs. (60) yield all previously established physical solutions for large-scale density perturbations in a flat flrw universe in its radiation-dominated phase and in the era following the decoupling of matter and radiation. In the development of the new cosmological perturbation theory, it has been demonstrated that in the nonrelativistic limit, relativistic gauge transformations reduce to Newtonian gauge transformations. Consequently, the conventional relativistic evolution equation—whether derived from general relativity or Newtonian gravitation modified to account for the expansion of the universe—also exhibits gauge-dependent solutions. This finding indicates that Newtonian gravitation is inadequate in accurately describing the evolution of density perturbations.
[…]
The investigation into cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic monatomic perfect gas with three degrees of freedom accurately represents the cosmic fluid in our universe following the decoupling of matter and radiation, then a perturbation theory based on general relativity provides an explanation for the origin of the first structures in the universe that does not rely on cdm. In contrast, the theory considers the local nonadiabatic random pressure fluctuations that emerged during the turbulent transition to the era that followed the decoupling of matter and radiation. Subsequent observations with the James Webb Space Telescope are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe.
Appendix A Derivation of the Evolution Equations
Table 3: The coefficients $\alpha_{ij}$ figuring in the equations (125)
$3H(1+p_{\varepsilon})+\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{2H}$ $3Hp_{n}$ $\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{4H}$
$\dfrac{\kappa n_{\scriptscriptstyle(0)}}{2H}$ $3H$ $n_{\scriptscriptstyle(0)}$ $\dfrac{n_{\scriptscriptstyle(0)}}{4H}$
$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 00
$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 00 $-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 00 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
In this appendix, the perturbation equations (60) are derived. Prior to undertaking the subsequent analysis, it is necessary to carry out two preliminary steps.
Firstly, as $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$, (4b), it is unnecessary to consider the gauge-dependent quantity $\theta_{{\scriptscriptstyle(1)}}$. As a result, the system (37) can be rewritten in a more suitable form. By eliminating the variable $\theta_{\scriptscriptstyle(1)}$ from the differential equations  (37b)–(37e) using the algebraic constraint equation (37a), a system of four first-order ordinary differential equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})$
$+\varepsilon_{\scriptscriptstyle(0)}(1+w)\Bigl{[}\vartheta_{\scriptscriptstyle(1)}+\frac{1}{2H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)\Bigr{]}=0,$ (123a)
$\dot{n}_{\scriptscriptstyle(1)}$ $+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\Bigl{[}\vartheta_{\scriptscriptstyle(1)}+\frac{1}{2H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)\Bigr{]}=0,$ (123b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}$ $+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}p_{\scriptscriptstyle(1)}}{a^{2}}=0,$ (123c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$ $+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$
$-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}+\frac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)=0.$ (123d)
Secondly, the energy density and particle number density perturbations (4a) are expressed in the four quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Employing the background equations (10) to eliminate all time derivatives and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\dfrac{\varepsilon_{\scriptscriptstyle(1)}{{}^{3}\!}R_{\scriptscriptstyle(0)}-3\varepsilon_{\scriptscriptstyle(0)}(1+w)(2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (124a)
$n^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n_{\scriptscriptstyle(1)}-\dfrac{3n_{\scriptscriptstyle(0)}(\kappa\varepsilon_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (124b)
From an algebraic standpoint, it is more straightforward to begin by deriving equations for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. This will be accomplished in Sec. A.1. Subsequently, Eqs. (60) for the fluctuations (59) will be derived in Sec. A.2. The expression (44) will be utilized to eliminate the time derivative of $w$.
A.1 Evolution Equation for Energy Density Perturbations
The system (123) and the expression (124a) are reformulated using (13a) in a form suitable for implementation in a computer algebra program:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+\alpha_{11}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{12}n_{\scriptscriptstyle(1)}+\alpha_{13}\vartheta_{\scriptscriptstyle(1)}+\alpha_{14}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125a)
$\dot{n}_{\scriptscriptstyle(1)}+\alpha_{21}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{22}n_{\scriptscriptstyle(1)}+\alpha_{23}\vartheta_{\scriptscriptstyle(1)}+\alpha_{24}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+\alpha_{31}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{32}n_{\scriptscriptstyle(1)}+\alpha_{33}\vartheta_{\scriptscriptstyle(1)}+\alpha_{34}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\alpha_{41}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{42}n_{\scriptscriptstyle(1)}+\alpha_{43}\vartheta_{\scriptscriptstyle(1)}+\alpha_{44}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (125d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\alpha_{51}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{52}n_{\scriptscriptstyle(1)}+\alpha_{53}\vartheta_{\scriptscriptstyle(1)}+\alpha_{54}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0.\!$ (125e)
The coefficients $\alpha_{ij}$ are given in Table 3.
The quantities in question (124) do not include the gauge function $\psi(\bm{x})$. Consequently, the gauge modes (147) will naturally disappear as a result of deriving the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$. This process will be carried out in three steps.
Step 1.
The first step is to remove the explicit occurrence of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ from Eqs. (125). By differentiating Eq. (125e) with respect to time and eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{{\scriptscriptstyle(1)}}$, $\vartheta_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ with the help of Eqs. (125a)–(125d), one obtains the following equation:
$\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{1}\varepsilon_{\scriptscriptstyle(1)}+p_{2}n_{\scriptscriptstyle(1)}+p_{3}\vartheta_{\scriptscriptstyle(1)}+p_{4}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=0,$ (126)
where the coefficients $p_{1},\ldots,p_{4}$ are given by
$p_{i}=\dot{\alpha}_{5i}-\alpha_{51}\alpha_{1i}-\alpha_{52}\alpha_{2i}-\alpha_{53}\alpha_{3i}-\alpha_{54}\alpha_{4i}.$ (127)
As can be seen from Eq. (126), it follows that
${{}^{3}\!}R_{\scriptscriptstyle(1)}=-\dfrac{1}{p_{4}}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{p_{1}}{p_{4}}\varepsilon_{\scriptscriptstyle(1)}-\dfrac{p_{2}}{p_{4}}n_{\scriptscriptstyle(1)}-\dfrac{p_{3}}{p_{4}}\vartheta_{\scriptscriptstyle(1)}.$ (128)
In this manner, ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is expressed as a linear combination of the quantities $\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\vartheta_{\scriptscriptstyle(1)}$. Upon replacing ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ in Eqs. (125) by the right-hand side of (128), the following system of equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+q_{1}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{11}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{12}n_{\scriptscriptstyle(1)}+\gamma_{13}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129a)
$\dot{n}_{\scriptscriptstyle(1)}+q_{2}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{21}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{22}n_{\scriptscriptstyle(1)}+\gamma_{23}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+q_{3}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{31}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{32}n_{\scriptscriptstyle(1)}+\gamma_{33}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+q_{4}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\gamma_{41}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{42}n_{\scriptscriptstyle(1)}+\gamma_{43}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+q_{5}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\gamma_{51}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{52}n_{\scriptscriptstyle(1)}+\gamma_{53}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (129e)
where the coefficients $q_{i}$ and $\gamma_{ij}$ are given by
$q_{i}=-\dfrac{\alpha_{i4}}{p_{4}},\quad\gamma_{ij}=\alpha_{ij}+q_{i}p_{j}.$ (130)
It has been achieved that ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occurs explicitly only in Eq. (129d), whereas ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occurs implicitly in the remaining equations. Therefore, Eq. (129d) is no longer required. Accordingly, the remaining four ordinary differential equations are as follows: (129a)–(129c) and (129e) for the four unknown quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
Step 2.
In a manner analogous to the approach undertaken in Step 1, the explicit appearance of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (129). Differentiating Eq. (129e) with respect to time and subsequently eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{\scriptscriptstyle(1)}$ and $\vartheta_{\scriptscriptstyle(1)}$ with the help of Eqs. (129a)–(129c), results in the following equation:
$q_{5}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+r\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+s_{1}\varepsilon_{\scriptscriptstyle(1)}+s_{2}n_{\scriptscriptstyle(1)}+s_{3}\vartheta_{\scriptscriptstyle(1)}=0,$ (131)
where the coefficients $r$ and $s_{i}$ are given by
$s_{i}$ $=\dot{\gamma}_{5i}-\gamma_{51}\gamma_{1i}-\gamma_{52}\gamma_{2i}-\gamma_{53}\gamma_{3i},$ (132a)
$r$ $=1+\dot{q}_{5}-\gamma_{51}q_{1}-\gamma_{52}q_{2}-\gamma_{53}q_{3}.$ (132b)
As can be seen from Eq. (131), it follows that
$\vartheta{\scriptscriptstyle(1)}=-\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{r}{s_{3}}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{s_{1}}{s_{3}}\varepsilon_{\scriptscriptstyle(1)}-\dfrac{s_{2}}{s_{3}}n_{\scriptscriptstyle(1)}.$ (133)
In this manner, $\vartheta_{\scriptscriptstyle(1)}$ is expressed as a linear combination of $\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$. Upon replacing $\vartheta_{\scriptscriptstyle(1)}$ in Eqs. (129) by the right-hand side of (133), the follwing system of equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{13}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{1}-\gamma_{13}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{11}-\gamma_{13}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{12}-\gamma_{13}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134a)
$\dot{n}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{23}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{2}-\gamma_{23}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{21}-\gamma_{23}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{22}-\gamma_{23}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{33}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{3}-\gamma_{33}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{31}-\gamma_{33}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{32}-\gamma_{33}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{43}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{4}-\gamma_{43}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{41}-\gamma_{43}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{42}-\gamma_{43}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (134d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\,-$ $\;\gamma_{53}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{5}-\gamma_{53}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{51}-\gamma_{53}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{52}-\gamma_{53}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0.$ (134e)
It has been achieved that the quantities $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occur explicitly only in Eqs. (134c) and (134d), whereas they occur implicitly in the remaining equations. Consequently, Eqs. (134c) and (134d) are no longer required. The remaining equations (134a), (134b) and (134e) are three ordinary differential equations for the three unknown quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
Step 3.
The subsequent steps would be to eliminate, in sequence, $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$ from Eq. (134e) with the aid of Eqs. (134a) and (134b). This would result in a fourth-order differential equation for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. It must be noted, however, that this is not a possibility, given that the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$ do not occur explicitly in Eq. (134e). To demonstrate this, Eq. (134e) will be rewritten in the subsequent form:
$\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{1}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{2}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=a_{3}\left(n_{\scriptscriptstyle(1)}+\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}\varepsilon_{\scriptscriptstyle(1)}\right),$ (135)
where the coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by
$a_{1}=-\dfrac{s_{3}}{\gamma_{53}}+\dfrac{r}{q_{5}},\quad a_{2}=-\dfrac{s_{3}}{\gamma_{53}q_{5}},\quad a_{3}=\dfrac{\gamma_{52}s_{3}}{\gamma_{53}q_{5}}-\dfrac{s_{2}}{q_{5}}.$ (136)
As a consequence of the gauge-invariance of the quantity $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, the left-hand side of Eq. (135) is also gauge-invariant. It thus follows that the right-hand side of this equation is also gauge-invariant, as will be demonstrated subsequently. Indeed, the result is as follows:
$\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}=-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (137)
It can be seen that the conservation laws (10c) and (10d) and the expressions (4a) lead to the following conclusion:
$n_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (138)
It follows from expressions (137) and (138) that the right-hand side of Eq. (135) does not explicitly contain the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$. Consequently, it is gauge-invariant. Therefore, equation (135) governs the evolution of $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
The proof of the equality (137) is straightforward but requires significant computational effort. The algebraic task was performed using the computer algebra system Maxima [44].
The final step in this process will be deriving an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. The linearized equations (123), from which the evolution equations are derived, are of fourth order. From this system, a second-order equation (135) for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ has been derived. It thus follows that the remaining system from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order. As the gauge-invariant quantities, namely, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been employed, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated. Consequently, only a first-order evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived. In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the expression (138), which contains $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. By differentiating the left-hand side of equation (138) with respect to time and employing the background equations (10c) and (10d), the first-order equations (123a) and (123b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
$\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right)$
$=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right).$ (139)
The algebraic task was performed using the Maxima [44] computer algebra software. Due to the equality in Eq. (138), it is possible to replace $n_{\scriptscriptstyle(1)}$ and $\varepsilon_{\scriptscriptstyle(1)}$ in Eq. (A.1) by $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
A.2 Evolution Equations for Density Fluctuations
First Eq. (60a) will be derived. Upon substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ in Eq. (135) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$, the result is
$b_{1}=2\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1},\quad b_{2}=\dfrac{\ddot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1}\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{2},\quad b_{3}=a_{3}\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}.$ (140)
The coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by (136). With the aid of the computational software Maxima [44], the coefficients (61) of the main text have been calculated.
Finally, Eq. (60b) will be derived. From the definitions (59), it can be deduced that
$n_{\scriptscriptstyle(1)}^{\textnormal{phys}}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}=n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (141)
By employing the equality (138) and substituting the expression (141) into Eq. (A.1), the following is obtained:
$\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right]$
$=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right].$ (142)
Derivation of Eq. (60b) of the main text was accomplished through the utilization of Eq. (10d).
Appendix B Verifying the Evolution Equations for Scalar Perturbations
The present paper does not provide a derivation of the linearized equations (12). A demonstration of the validity of the system of equations (37), which is derived from the aforementioned linearized equations and is fundamental to the new perturbation theory, will therefore be given via three different methods.
The initial step in this process is to utilize the well-established property of the Einstein equations and conservation laws. This implies that if the constraint equations are satisfied at an initial time, then the solutions of the dynamical equations and conservation laws will satisfy the constraint equations for all subsequent times. A more detailed discussion can be found in Weinberg’s book [31], in Sec. 7.5 on the Cauchy problem. This is also true in reverse: if solutions of the constraint equations and conservation laws can be found for all times, they must necessarily satisfy the dynamical equations. The system (37) comprises the conservation laws and constraint equations, and consists of four first-order ordinary differential equations and one algebraic equation. It thus follows that the system (37), can be solved without the use of the dynamical equations, since the latter are inherently satisfied by the solution of the aforementioned system. Accordingly, the latter of the two observations will be employed to substantiate the validity of equations (37).
The off-diagonal ($i\neq j$) equations of (12c) are not associated with scalar perturbations. For the diagonal ($i=j$) equations, it is sufficient to consider the contraction of the dynamical equations (12c):
$\ddot{h}^{k}{}_{k}+6H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-3\kappa(\varepsilon_{{\scriptscriptstyle(1)}}-p_{{\scriptscriptstyle(1)}}).$ (143)
Eliminating the quantity $\dot{h}^{k}{}_{k}$ with the help of (19), yields the following result:
$(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})+6H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})-{{}^{3}\!}R_{\scriptscriptstyle(1)}=\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}).$ (144)
Using the energy density constraint equation (37a) to eliminate the second term of (144), results in the following equation:
$\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)}).$ (145)
This equation is identical to the time derivative of the energy density constraint equation (37a), as will now be demonstrated. Differentiation of (37a) with respect to time yields
$2\dot{H}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})+2H(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\kappa\dot{\varepsilon}_{\scriptscriptstyle(1)}.$ (146)
Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (10a)–(10c), the momentum constraint equation (37b) and the energy density conservation law (37c), respectively, yields the dynamical equation (145). Consequently, Eqs. (37a)–(37c) are correct.
As indicated in the aforementioned verification, the correctness of the equations (37d) and (37e) has not yet been verified. This will now be addressed by employing the gauge modes. It is observed that Eqs. (37) remain invariant under the gauge transformations (1) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (6). Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are known to be gauge-dependent. It is therefore evident that the gauge modes
$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta,p,$ (147a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}$ =~$=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H\left(\frac{\tilde{\nabla}^{2}\psi}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\psi\right),$ (147c)
must be solutions to Eqs. (37). The gauge modes (147a) are derived directly from (3), with the understanding that $\xi^{0}=\psi$. The gauge mode (147b) is the covariant divergence of (28). The gauge mode (147c) is derived by substituting the gauge modes (147a)–(147b) into the perturbed Friedmann equation (37a):
${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H(\hat{\theta}_{{\scriptscriptstyle(1)}}-\hat{\vartheta}_{{\scriptscriptstyle(1)}})-2\kappa\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$
$=4H\left(\psi\dot{\theta}_{{\scriptscriptstyle(0)}}+\frac{\tilde{\nabla}^{2}\psi}{a^{2}}\right)-2\kappa\psi\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (148)
Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (10a), and Eq. (10b), one arrives at the gauge mode (147c). It will now be demonstrated that the gauge modes given by (147) are indeed solutions of Eqs. (37b)–(37e). In order to execute this task, the first step entails deriving the time derivatives of the gauge modes:
$\dot{\hat{\vartheta}}_{{\scriptscriptstyle(1)}}$ $=-2H\hat{\vartheta}_{{\scriptscriptstyle(1)}},$ (149a)
${{}^{3}\!}\dot{\hat{R}}_{{\scriptscriptstyle(1)}}$ $=-4(\dot{H}-2H^{2})\bigl{[}\hat{\vartheta}_{{\scriptscriptstyle(1)}}+\tfrac{1}{2}\,{{}^{3}\!}R_{{\scriptscriptstyle(0)}}\psi\bigr{]},$ (149b)
$\dot{\hat{\varepsilon}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{\varepsilon}_{{\scriptscriptstyle(0)}}=-3\varepsilon_{{\scriptscriptstyle(0)}}(1+w)\bigl{[}\dot{H}-3H^{2}(1+\beta^{2})\bigr{]}\psi,$ (149c)
$\dot{\hat{n}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{n}_{{\scriptscriptstyle(0)}}=-3n_{{\scriptscriptstyle(0)}}\bigl{[}\dot{H}-3H^{2}\bigr{]}\psi,$ (149d)
where Eqs. (10b)–(10d) and Eq. (44) have been used. Subsequent to the aforementioned preparation, the verification of the accuracy of the equations (37b)–(37e) can be undertaken. First, by substituting the gauge modes, namely $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq. (37b), the result is an equation from which the variable $\psi$ disappears. Next, dividing the resulting equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying by $\tfrac{3}{2}H$, one arrives at the time derivative of Eq. (10a). Second, the substitution of the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37c) results in the elimination of the term with $\dot{H}$. Next, dividing, the resulting equation by $\psi$ and employing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by utilizing Eq. (10c) yields an identity. Third, by substituting the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37d), and utilizing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq. (10c) one arrives at an identity. Finally, upon substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq. (37e) the term with $\dot{H}$ vanishes. Using Eq. (10d), one arrives at an identity. It has been demonstrated that the gauge modes (147) satisfy Eqs. (37b)–(37e). Consequently, it can be concluded that these equations are accurate. Given the established validity of Eq. (37a), it can be deduced that the system (37) is also accurate.
As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (152) for the gauge-dependent density fluctuation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs. (37). Of particular note is the fact that the well-known evolution equations for $\delta$ are recovered in the cases of a radiation-dominated universe (154) and a universe after decoupling of matter and radiation (156).
Consequently, it can be concluded that the system of equations (37) has been verified in three distinct ways.
Appendix C Conventional Evolution Equation for Cosmological Density Fluctuations
The usual evolution equation for density perturbations in a flat, ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}=0$, flrw universe is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\frac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)\right)\delta=0,$ (150)
where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$. The derivation of this equation has four primary objectives. Firstly, it demonstrates the correctness of the system (37), at least in the context of a flat flrw universe. Secondly, it is demonstrated that the standard equation (150) is incomplete. That is, its solution is a linear combination of only one physical mode and one gauge mode, independent of the scale of a fluctuation. Consequently, Eq. (150) does not give an adequate description of cosmological density perturbations. Thirdly, it is shown that the novel perturbation theory (60) provides all well-known physical solutions for large-scale perturbations in a flat flrw universe. Specifically, it produces the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe after decoupling of matter and radiation. Finally, it will be demonstrated that Eq. (150), when derived from the Newtonian theory of gravitation, yields solutions that remain dependent on the selection of the coordinates.
In order to derive equation (150), it is necessary to utilize the background equations (71) and the equations (123) for scalar perturbations. These two systems of equations are complete, since they contain all conservation laws and constraint equations. In this appendix, it is assumed that $c=1$.
Equation (150) has been derived from an equation of state $p=p(\varepsilon)$ in the existing literature. Therefore, the particle number conservation law (71c) and its linearized counterpart (123b) are not required. Substituting $p_{n}=0$ into (13) and (40) yields $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$. According to the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs. (123) can be rewritten as follows:
$\dot{\delta}+3H\delta\bigl{(}\beta^{2}+\tfrac{1}{2}(1-w)\bigr{)}+(1+w)\left(\vartheta_{\scriptscriptstyle(1)}+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(1)}}{4H}\right)=0,$ (151a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\frac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0,$ (151b)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}=0,$ (151c)
where Eq. (71b) has been used.
In order to arrive at a generalization of (150), the following steps must be taken. First, it is necessary to differentiate equation (151a) with respect to time. Subsequently, the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ must be eliminated. This can be achieved by utilizing the equations (44), (71a) and (71b) and the perturbation equations (151b) and (151c). The final step in the process is the elimination of ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ through the application of (151a). This results in the following system of equations:
$\ddot{\delta}+2H\dot{\delta}\bigl{(}1+3\beta^{2}-3w\bigr{)}$
$\qquad-\biggl{[}\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl{(}(1+w)(1+3w)\Bigr{.}\Bigr{.}$
+ $\qquad+\,4w-6w^{2}+12\beta^{2}w-4\beta^{2}-6\beta^{4}\Bigr{)}-6\beta\dot{\beta}H\biggr{]}\delta$
$\quad=-3H\beta^{2}(1+w)\vartheta_{\scriptscriptstyle(1)},$ (152a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\dfrac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (152b)
In this procedure the computer algebra program Maxima [44] has been used.
The gauge modes (147) are for the system (152) given by
$\hat{\delta}(t,\bm{x}):=\dfrac{\psi(\bm{x})\dot{\varepsilon}_{\scriptscriptstyle(0)}(t)}{\varepsilon_{\scriptscriptstyle(0)}(t)}=-3H(t)\psi(\bm{x})\bigl{(}1+w(t)\bigr{)},$ (153a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}(t)}.$ (153b)
The fact that the gauge modes (153) satisfy the system (152) has been shown in Appendix B.
It can be inferred from the set of equations (152) and the gauge modes (153) that:
  • The solution of the homogeneous part of Eq. (152a) is a linear combination of a physical mode and the gauge mode (153a). The particular solution is obtained by solving Eq. (152b), and contains the gauge mode (153b).
  • It is impossible to determine the arbitrary infinitesimal gauge function $\psi(\bm{x})$ through physical means. Therefore, the imposition of physical initial values on the system (152) does not yield a physical solution.
  • In the small-scale limit, the gauge modes (153) do not vanish. This means that in the small-scale limit gauge-dependent quantities do not become gauge-invariant. Consequently, $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ does not become equal to its Newtonian counterpart $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, given by (56) and (58). This is consistent with the gauge transformation (53) of the Newtonian theory of gravitation.
  • A notable distinction between the conventional equation (150) and Eq. (152a) is that the former is homogeneous, whereas Eq. (152a) is inhomogeneous. This is because the system (152) is derived from the complete set of constraint equations and conservation laws. In the case where the right-hand side of Eq. (152a) is absent, as is the case in Eq. (150), then (153a) is a gauge mode if $\nabla^{2}\psi=0$. Consequently, $\psi=C$ is an arbitrary infinitesimal constant, so that the gauge mode (153a) does not vanish.
  • In the universe following decoupling, the values of $w\ll 1$ and $\beta^{2}\ll 1$ are negligible, allowing for the subsequent neglect of $w$ in relation to the constants of order one in Eq. (150). This is equivalent to deriving the usual equation (150) from the Newtonian theory of gravitation adapted to the expansion of the universe. See, for example, Ref. [10], Sec. 6.2, Ref. [31], Sec. 15.9, and Ref. [45], Eq. (3.17). However, the gauge transformation (53) in the nonrelativistic limit ensures that the gauge mode (153a) does not vanish.
From these points it is clear that the conventional equation (150) is unsuitable to study the evolution of density perturbations in the universe. Given that Eq. (150) can be derived from Newtonian gravitation for $w\ll 1$, it follows that the Newtonian theory is unable to provide an adequate description of density perturbations in the universe.
In the ensuing two sections, a comparison is presented between Eqs. (79) and (95), with the system of equations (152). This comparison will be conducted in the case of large-scale fluctuations. It will become evident that the gauge modes preclude the attainment of physical solutions.
C.1 Radiation-Dominated Era
In this era, the pressure is given by a linear barotropic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and ${w=\tfrac{1}{3}}$. In this case Eqs. (152) result in the system
$\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-\tfrac{4}{3}H\vartheta_{\scriptscriptstyle(1)},$ (154a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H\vartheta_{\scriptscriptstyle(1)}+\dfrac{1}{4}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (154b)
The gauge modes (153) are solutions of the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
For large-scale fluctuations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution of the system (154) is
$\delta(t,\bm{x})=(c_{1}t-2\psi(\bm{x})t^{-1})+\tfrac{9}{8}t^{1/2},$ (155a)
$\vartheta_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}}+\tfrac{9}{8}t^{-1/2},$ (155b)
where it is used that $H=1/2t$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=3/4t^{2}$. The expression between parentheses in (155a) is the solution of the homogeneous part of Eq. (154a). The particular solution $\tfrac{9}{8}t^{1/2}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\tfrac{9}{8}t^{-1/2}$. The solution (85) derived from the novel perturbation equation (82) is not derivable from the standard equations (154) due to the presence of the gauge function $\psi$ in the solution (155a).
Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale fluctuations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$. That is why (79a) yields oscillating density fluctuations (84) with an increasing amplitude, instead of a constant amplitude which follows from (150).
C.2 Era after Decoupling of Matter and Radiation
In this era, the equation of state for the pressure is according to thermodynamics given by (88). Therefore, (91) and (92) imply that $w\approx\tfrac{3}{5}\beta^{2}\ll 1$, so that with $T_{\scriptscriptstyle(0)}\propto a^{-2}$ one obtains $\dot{\beta}/\beta=-H$. Using Eq. (71a) one can derive the following result: $6\beta\dot{\beta}H=-2\kappa\varepsilon_{\scriptscriptstyle(0)}\beta^{2}$. Upon substituting the latter expression into (152a) and neglecting $w$ and $\beta^{2}$ with respect to the constants of order one, the system (152) results in
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-3H\beta^{2}\vartheta_{\scriptscriptstyle(1)},$ (156a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\beta^{2}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (156b)
The gauge modes (153) are solutions of the system (156) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution. Consequently, for the system (156) $\psi$ is an arbitrary infinitesimal constant $C$. This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (153), vanishes identically. Consequently, for large-scale fluctuations the solution of the system (156) is
$\delta(t)=(c_{1}t^{2/3}-2Ct^{-1})+\tfrac{7}{9}t^{-5/3},$ (157a)
$\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(t)=-\tfrac{7}{9}t^{-4/3},$ (157b)
where $H=2/3t$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=4/3t^{2}$, and $\beta^{2}\propto t^{-4/3}$ has been used. The latter proportionality follows from (91) and (92), $T_{\scriptscriptstyle(0)}\propto a^{-2}$ and (104). The expression between parentheses in (157a) is the solution of the homogeneous part of Eq. (156a). The particular solution $\tfrac{7}{9}t^{-5/3}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\tfrac{7}{9}t^{-4/3}$. The solution (109) derived from the novel perturbation equation (108) is not derivable from the standard equations (156) due to the presence of the gauge constant $C$ in the solution (157a).
Since $\psi=C$, the gauge transformation (1) with $\xi^{\mu}(t,\bm{x})$ given by (6) reduces to the Newtonian gauge transformation (53). This is to be expected, since a cosmological fluid for which ${w\ll 1}$ and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (88).