### v3

# Earliest Structures in the Universe can be explained by a Relativistic Cosmological Perturbation Theory

Abstract
A relativistic cosmological perturbation theory for the Friedmann-Lemaître-Robertson-Walker universe is presented that explains the masses and formation times of the first structures in our universe. First, it is shown that, without a coordinate system being used, quantities intended to represent energy density and particle number density perturbations can be defined in only one way. The Newtonian limit, where the pressure becomes zero, proves that these quantities are indeed the perturbations of the energy density and the particle number density. Then, after selecting a reference frame, a perturbation theory will be formulated based on these quantities. This formulation considers the local perturbation to the spatial curvature resulting from a density perturbation, the local fluid velocity due to pressure gradients caused by the self-gravity of the density perturbation, and entropy perturbations, all of which are necessary for structure formation. Pressure perturbations consist of two components: an adiabatic component caused by the density perturbation itself, and a random, nonadiabatic component resulting from the rapid, chaotic transition to the era when matter and radiation were decoupled. Immediately after decoupling, negative nonadiabatic pressure perturbations in various density perturbations enabled their rapid growth over a short period of time. This brief period ended when the total pressure perturbation became positive. Subsequently, the density perturbations gradually grew toward their nonlinear phase, which was reached early on.
Cosmology; Early Universe; Structure Formation; Perturbation Theory; Gauge Problem.
pacs:
98.80.-k; 98.80.Cq; 97.10.Bt; 04.25.Nx.
I Introduction
Structures in our universe, such as stars, galaxies, and black holes, form through the gravitational growth of small density perturbations. Therefore, it is reasonable to expect that a cosmological perturbation theory based on general relativity — the most accurate theory of gravity to date — would explain the masses and formation times of the first structures. There are two main reasons why such a theory does not yet exist. First, the gauge problem of cosmology has not yet been solved. Second, particle number density has not been considered, so the pressure gradients necessary for fluid flow — which is required for structure formation after matter and radiation decoupled — cannot be taken into account. This article resolves both issues.
The study of the evolution of density perturbations by way of linearization of the Einstein equations and conservation laws with respect to a flat Friedmann-Lemaître-Robertson-Walker (flrw) universe was pioneered by Lifshitz [1] and Lifshitz & Khalatnikov [2]. They observed that solutions to the linearized Einstein equations and conservation laws are dependent on the choice of a coordinate system, i.e., the solutions are gauge-dependent. Consequently, these solutions can be changed by linear coordinate transformations, also known as gauge transformations. This means that these solutions have no physical meaning, which is a problem known as the gauge problem of cosmology. Solving this long-standing problem is the only way to obtain physically significant solutions from the linearized Einstein equations and conservation laws. Doing so is paramount to explaining how the earliest structures [3, 4] in the universe formed from small density perturbations.
Physical phenomena such as density, temperature, pressure, and entropy perturbations do not depend on the choice of a coordinate system. Accordingly, the corresponding mathematical quantities describing these phenomena must be defined without using a reference frame. In other words, these quantities must be gauge-invariant.
Two different pioneering methods to address the gauge problem of cosmology have been proposed in the literature. In his seminal article, Bardeen [5] introduced gauge-invariant quantities intended to represent energy density perturbations as linear combinations of gauge-dependent quantities. Building on Bardeen’s work, Mukhanov et al.[6, 7] made significant contributions to the subject. Ellis et al.[8, 9, 10] present a second method that takes advantage of the fact that the gradient of a gauge-dependent quantity is gauge-invariant. For a comprehensive overview of this approach and its further extension, see Tsagas et al.[11].
Unfortunately, these approaches have some disadvantages. Firstly, they allow for different definitions of gauge-invariant quantities. Since these approaches lack a Newtonian limit, it is impossible to determine which gauge-invariant quantity corresponds to an energy density perturbation. Secondly, these approaches do not account for the particle number density. Thirdly, metric perturbations are explicitly present in the evolution equations for density perturbations. Finally, these approaches have not been extended to open and closed flrw universes.
Many new ideas have been added to this field. See Ellis’s $2017$ review article [12] and other publications, e.g., Refs. [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. These researchers proposed alternative perturbation theories using different gauge-invariant quantities that differ from those used by Bardeen and Mukhanov et al. However, the most fundamental question regarding which gauge-invariant quantity accurately and unambiguously represents a density perturbation remains unanswered.
This article presents a relativistic cosmological perturbation theory that does not have the aforementioned disadvantages. Firstly, it demonstrates that a gauge-invariant quantity describing a density perturbation can be defined in precisely one way. Secondly, in addition to energy density, particle number density [32] is considered. Both are crucial for structure formation. Finally, it is shown that perturbations in the spatial metric and its derivatives are included in the perturbations of the spatial Ricci scalar and expansion scalar. The use of these quantities significantly simplifies the system of linearized Einstein equations and conservation laws, leading not only to a perturbation theory that encompasses closed, flat, and open flrw universes but also allowing the Newtonian limit to be derived.
Unlike the approaches in Refs. [33, 34, 35, 36, 37, 38], this article does not introduce any ad hoc modifications to the theory of relativity, free parameters, Cold Dark Matter (cdm), or the peculiar velocities [39, 40, 41] of matter. However, the linearized Einstein equations and conservation laws naturally encompass the velocity of matter caused by pressure gradients, as well as the perturbed Ricci scalar caused by a density perturbation. These must be taken into account.
II Development and Overview of the Perturbation Theory
Sec. III explains why, in addition to energy density, particle number density must also be considered.
Sec. IV explains the gauge problem in cosmology and provides a solution by demonstrating that a gauge-invariant quantity that describes energy density perturbations can be defined in only one way. The same applies to the gauge-invariant perturbation of the particle number density. These definitions imply that the gauge-invariant perturbation to the expansion of the universe is equal to zero. In Sec. X, the Newtonian limit will be used to demonstrate that these quantities represent density perturbations.
To formulate a perturbation theory based on these quantities, a coordinate system is necessary. The reasons why a synchronous reference frame is the best choice for this task are explained in Sec. V.
In Sec. VI, the Einstein equations and conservation laws for flrw universes, as well as their linearized counterparts, are presented with respect to synchronous coordinates. Given that the evolution of density perturbations is considerably influenced by pressure perturbations, the covariant divergence of the fluid velocity, and the perturbation of the Ricci scalar, expressions are derived for these quantities.
Since we are working in synchronous coordinates, we can apply the decomposition theorem for symmetric, second-rank spatial tensors, such as the perturbed metric tensor and the perturbed spatial Ricci tensor, as well as the Helmholtz decomposition for the fluid velocity vector. These theorems are discussed in Sec. VII. According to these theorems, the linearized Einstein equations and conservation laws decompose into three independent systems of equations: one for tensor perturbations, one for vector perturbations, and one for scalar perturbations. In the literature, the decomposition theorem is only applied to the spatial metric tensor perturbation. However, this is insufficient. Since the perturbed Ricci scalar plays an important role in the evolution of density perturbations and in the Newtonian limit, the decomposition theorem must also be applied to the perturbed Ricci tensor.
Sec. VIII uses these theorems to show that only scalar perturbations are associated with density perturbations. In this case, the perturbed Ricci scalar contains the spatial derivatives of two independent potentials, while the perturbed expansion scalar contains their time derivatives. In the Newtonian limit, one of the potentials becomes the Newtonian potential while the other becomes irrelevant.
Since we are interested in density perturbations, we derive in Sec. IX a system of equations describing the evolution of scalar perturbations. This system contains three conservation laws: one for the energy density perturbation, one for the particle number density perturbation, and one for the momentum. It also contains two constraint equations. The algebraic energy density constraint equation contains the perturbed Ricci scalar, and the momentum constraint equation describes how the perturbed Ricci scalar evolves.
The Newtonian limit is reached when the velocities of the particles become negligible compared to the speed of light — that is, when the pressure approaches zero. Using this limit it is shown in Sec. X that the newly defined gauge-invariant quantities in Sec. IV represent perturbations of the energy and particle number densities. In other words, in the Newtonian limit, the momentum constraint equation ensures that the Newtonian potential and the energy density perturbation are independent of time. It also makes sure that the energy density constraint equation is identical to the Poisson equation of Newtonian gravity. Additionally, the well-known special relativistic equation relating rest mass density to energy density is obtained.
The Newtonian limit can only be obtained using the reformulated evolution equations for scalar perturbations given in Sec. IX. Since these equations are not presented in this form in the literature, previous cosmological perturbation theories have no Newtonian limit. Without a clearly defined Newtonian limit, it is impossible to demonstrate which gauge-invariant quantity represents a density perturbation. This article resolves this issue.
The Newtonian limit demonstrates furthermore that relativistic gauge transformations become Newtonian gauge transformations. Consequently, the conventional evolution equation for relative perturbations in a flat flrw universe filled with a nonrelativistic fluid — whether derived from general relativity or Newtonian gravity adapted to the expansion of the universe — also has gauge-dependent solutions devoid of physical meaning. Consequently, Newtonian gravity is not suitable for studying density perturbations in the universe, regardless of their scale. This means that no distinction can be made between density perturbations inside and outside the horizon.
The perturbation theory is presented in Sec. XI. It describes the evolution of density perturbations in closed, flat, and open flrw universes. The theory incorporates the perturbed Ricci scalar, the pressure, the entropy, and the covariant divergence of the fluid velocity. This is demonstrated by the algorithm in Appendix A.
In Sec. XII, expressions are derived for the perturbations of entropy, temperature, and pressure to connect the perturbation theory to an equation of state for a cosmological fluid. It will be shown that pressure perturbations have two components: an adiabatic component caused by the density perturbation itself, and a nonadiabatic component.
In Sec. XIII, we apply the perturbation theory to a flat flrw universe in its radiation-dominated phase and after the decoupling of matter and radiation.
Finally, we summarize the results and draw conclusions in Sec. XIV.
The article is accompanied by three appendices. Appendix A presents an algorithm that, when executed with the assistance of a computer algebra program, derives the perturbation theory from the background Einstein equations and the system of equations for scalar perturbations. Since the latter system is a new result, its correctness is demonstrated in Appendices B and C. Finally, Appendix C compares the conventional perturbation equations with the perturbation theory developed in this article.
III Equations of State
We distinguish between two main eras of the universe: the period before the decoupling of matter and radiation and the post-decoupling era.
Before decoupling, the cosmological fluid can be treated as a relativistic fluid consisting of radiation, ordinary matter, and cdm, where the contributions of matter and cdm to the pressure are negligible and the radiation pressure is of the same order of magnitude as the energy density. See, e.g., Ref. [6], Eq. (5.49), Ref. [7], Eq. (1.78), and Ref. [13], § V-1.
The conditions in the universe became significantly different after decoupling of matter and radiation. This is because the pressure in the cosmological fluid has become negligible compared to its energy density. Therefore, pressure can be neglected in the Einstein equations and conservation laws describing the evolution of a flrw universe. However, pressure perturbations cannot be ignored when studying structure formation. The self-gravity of a local density perturbation creates pressure gradients. These gradients then cause particle flows, thereby increasing a local density perturbation. The linearized momentum conservation law shows that the absence of pressure perturbations implies an absence of fluid flows. This results in static density perturbations. Therefore, pressure perturbations are crucial to understanding how structures in the universe formed. After decoupling, the cosmological fluid can be treated as a perfect nonrelativistic gas. In this case, pressure depends on the average kinetic energy density of the particles. Therefore, it is essential to include, next to the rest energy density, also the kinetic energy density in the analysis. Both pressure and kinetic energy density depend on the particle number density and temperature. Because of this, particle number density and temperature must be included in the equations of state. See Ref. [42], Sec. 2.10 on relativistic hydrodynamics, and Refs. [32, 43]. This results in random, nonadiabatic density perturbations, which are crucial for the formation of structures in the universe.
The aforementioned considerations are consistent with thermodynamic principles, namely that energy density $\varepsilon$ and pressure $p$ are functions of particle number density $n$ and temperature $T$, i.e.,
$\varepsilon=\varepsilon(n,T),\quad p=p(n,T).$ (1)
In order to facilitate the calculations, it is advantageous to eliminate the temperature from the given equations of state. This yields the equation of state for the pressure:
$p=p(n,\varepsilon),$ (2)
which will be used in the subsequent analysis.
IV Gauge Problem and its Solution
Einstein’s equations and conservation laws are invariant under coordinate transformations $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. This signifies that there are no preferred coordinate systems. Consequently, the linearized Einstein equations and conservation laws are gauge-invariant, meaning they are invariant under linear, infinitesimal space-time transformations, also known as gauge transformations, which are given by
$x^{\mu}\rightarrow x^{\prime\mu}=x^{\mu}-\xi^{\mu}(t,\bm{x}),$ (3)
where the gauge functions $\xi^{\mu}(t,\bm{x})$ are four arbitrary infinitesimal functions of time, $x^{0}=ct$ with $c$ the speed of light, and space, $\bm{x}=(x^{1},x^{2},x^{3})$, coordinates. The gauge problem of cosmology stems from the linear transformation (3) and the linearity of the linearized Einstein equations and conservation laws. This problem will now be examined.
Consider a closed, flat, or open flrw universe. The evolution equations for these universes contain three scalars: the energy density $\varepsilon$, the particle number density $n$, and the expansion $\theta$. These scalars are defined as follows
$\varepsilon:=T^{\mu\nu}u_{\mu}u_{\nu},\quad n:=N^{\mu}u_{\mu},\quad\theta:=u^{\mu}{}_{;\mu}$, (4)
where $u^{\mu}$ is the fluid four-velocity normalized to unity, $u^{\mu}u_{\mu}=1$}, and $N^{\mu}:=nu^{\mu}$ represents the cosmological particle current four-vector, which satisfies the particle number conservation law $N^{\mu}{}_{;\mu}=0$}. A semicolon denotes covariant differentiation. Let $S_{\scriptscriptstyle(0)}(t)$ denote the quantities $\varepsilon_{\scriptscriptstyle(0)}$, $n_{\scriptscriptstyle(0)}$, and $\theta_{\scriptscriptstyle(0)}$. These quantities satisfy the background equations, that is to say, the unperturbed Einstein equations and conservation laws. Let $S_{\scriptscriptstyle(1)}(t,\bm{x})$ be their perturbed counterparts $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$, which satisfy the linearized equations. As a consequence of the linearity of these equations, new solutions that have precisely the same physical content can be generated. These new solutions are given by (see Weinberg [42], Sec. 10.9 for a detailed explanation)
$S^{\prime}_{\scriptscriptstyle(1)}=S_{\scriptscriptstyle(1)}+\mathcal{L}_{\xi}S_{\scriptscriptstyle(0)}=S_{\scriptscriptstyle(1)}+\xi^{0}\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta.$ (5)
In this expression, the operator $\mathcal{L}_{\xi}$ is the Lie derivative with respect to the infinitesimal four-vector $\xi^{\mu}$. An overdot denotes differentiation with respect to $x^{0}=ct$. The solutions $S_{\scriptscriptstyle(1)}$ and $S^{\prime}_{\scriptscriptstyle(1)}$ contain the so-called gauge modes which are given by $\hat{S}_{\scriptscriptstyle(1)}=\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This indicates that the solutions $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ to the linearized equations are gauge-dependent. Since these solutions depend on the choice of a reference frame, they are devoid of physical significance. Therefore, $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ do not represent physical phenomena. This is the gauge problem of cosmology, whereby the coordinate artifacts and the underlying physics are inextricably linked in the solution to the linearized equations. Now that the gauge problem of cosmology has been understood, a solution can be found.
The physics of the perturbations is hidden in the solutions to the linearized Einstein equations and conservation laws. Consequently, quantities that describe physical phenomena must be linear combinations of gauge-dependent solutions to these equations. Since the transformation of the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\theta_{\scriptscriptstyle(1)}$ under the infinitesimal transformation (3) is identical to that described by (5), we can construct gauge-invariant quantities by combining these three gauge-dependent quantities in a way that eliminates the gauge modes $\xi^{0}\dot{S}_{\scriptscriptstyle(0)}$. This results in three distinct sets of linear combinations. In each of these sets, precisely one gauge-invariant quantity is identically zero. Since we are only interested in energy density and particle number density perturbations, the only possible set is
$\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\varepsilon_{\scriptscriptstyle(1)}-\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad n_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=n_{\scriptscriptstyle(1)}-\dfrac{\dot{n}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},$ (6a)
$\theta_{\scriptscriptstyle(1)}^{\textnormal{phys}}:=\theta_{\scriptscriptstyle(1)}-\dfrac{\dot{\theta}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}=0.$ (6b)
From these definitions, it can be inferred that the general form of a gauge-invariant quantity describing a physical phenomenon is given by
$q^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}:=q_{{\scriptscriptstyle(1)}}-\dfrac{\dot{q}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}},$ (7)
where $q$ stands for energy density $\varepsilon$, particle number density $n$, temperature $T$, pressure $p$, or the entropy per particle $s$.
The quantities (6a) represent density perturbations and (6b) represents the perturbation to the expansion of the universe. This will be demonstrated in Sec. X on the Newtonian limit.
V Selection of a Reference Frame
The quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ are defined without the use of a coordinate system. However, to derive evolution equations for these quantities, it is first necessary to choose a suitable reference frame. Since general relativity is covariant, and the quantities in question are gauge-invariant, it follows that the evolution equations for these quantities can be derived in any chosen coordinate system. The selection of an appropriate coordinate system for this problem is dependent on two criteria, which are outlined below.
Firstly, to show that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent density perturbations requires taking the Newtonian limit. In the context of general relativity, coordinate transformations are in general space-time transformations, specifically $x^{\mu}\rightarrow x^{\prime\mu}(x^{\nu})$. In the linear case, space-time transformations are given by (3). In Newtonian gravity, space and time are different and independent entities. Consequently, in the Newtonian limit, the relativistic space-time transformations (3) must be automatically separated into independent space and time transformations. Secondly, it would be preferable to have a coordinate system that would facilitate the derivation of the perturbation theory. It will now be demonstrated that both of these requirements can be met by employing the same coordinate system.
Firstly, in Newtonian gravity, where space and time are treated as distinct and independent entities, all coordinate systems are inherently synchronous. In view of the Newtonian limit, it can be seen that a synchronous reference frame represents the optimal choice. In these coordinates, the metric tensor $g_{\mu\nu}(t,\bm{x})$ of the flrw universes is given by
$g_{00}=1,\quad g_{0i}=0,\quad g_{ij}=-a^{2}(t)\tilde{g}_{ij}(\bm{x}).$ (8)
The scale factor of the universe is designated as $a(t)$. Given that $g_{00}=1$, coordinate time is equal to proper time. The synchronicity condition is $g_{0i}=0$, as explained in Ref. [42], Sec. 11.8 and Ref. [44], § 84. The metric tensor of the three-dimensional maximally symmetric subspaces of constant time is denoted by $\tilde{g}_{ij}$. The Killing equations $\xi_{\mu;\nu}+\xi_{\nu;\mu}=0$ and (8) demonstrate that the gauge functions $\xi^{\mu}(t,\bm{x})$ in the transformation (3) become
$\xi^{0}=\psi(\bm{x}),\quad\xi^{i}=\tilde{g}^{ik}(\bm{x})\dfrac{\partial\psi(\bm{x})}{\partial x^{k}}\int\!\!\frac{{\mathrm{d}}t}{a^{2}(t)}+\chi^{i}(\bm{x}),$ (9)
if only transformations between synchronous coordinates are allowed (see Ref. [42], Sec. 10.9). In expressions (9), $\psi(\bm{x})$ and $\chi^{i}(\bm{x})$ are four arbitrary infinitesimal functions of the spatial coordinates.
Secondly, in synchronous coordinates the space-space components $R_{ij}$ of the four-dimensional Ricci tensor $R_{\mu\nu}$ are partitioned into two distinct parts. One part exclusively comprises the time derivatives of the spatial metric tensor $g_{ij}$ while the second part is the Ricci tensor ${{}^{3}\!}R_{ij}$ of the three-dimensional subspaces. This is demonstrated in Ref. [44], § 97. This implies that the spatial metric and Ricci tensor perturbations are both tensors. Furthermore, the spatial part of the fluid four-vector is a vector in synchronous coordinates. This allows us to apply the decomposition theorem for second-rank, symmetric three-tensors and the Helmholtz theorem to decompose the fluid vector. Using these theorems, it is shown that the linearized Einstein equations and conservation laws can be decomposed into three mutually exclusive systems: one for tensor perturbations, one for vector perturbations, and one for scalar perturbations. Next, it will be demonstrated that only scalar perturbations are associated with density perturbations. Therefore, when studying density perturbations, it is sufficient to consider the evolution equations for scalar perturbations. This system is easier to use than the original set of equations, thus facilitating the derivation of a cosmological perturbation theory. Because synchronous coordinates are also compatible with Newtonian gravity, they will be used from this point onward.
VI Einstein Equations and Conservation Laws
As was concluded in the previous section, synchronous coordinates are the optimal choice for deriving a perturbation theory. This section presents the background Einstein equations and conservation laws, as well as their linearized counterparts, in this coordinate system.
The background energy density constraint equation and its linearized counterpart are expressed in a manner consistent with the contracted Bianchi identities. These identities state that the Riemann tensor satisfies the equation $R^{\mu}{}_{\nu;\mu}=\tfrac{1}{2}R_{,\nu}$, where a comma denotes ordinary differentiation. This can be expressed in an equivalent manner as the four-divergence of the Einstein tensor, i.e., $G^{\mu\nu}{}_{;\nu}=0$}. This result demonstrates that the constraint equations contain at most first-order time derivatives of the metric, which is precisely what is required for the development of a cosmological perturbation theory. For further details see Ref. [42], Secs. 6.8 and 7.5 and Ref. [44], § 92 and § 95.
VI.1 Background Equations
The system of background Einstein equations and conservation laws for closed, flat, and open flrw universes filled with a perfect fluid with an energy-momentum tensor as defined by
$T^{\mu\nu}=(\varepsilon+p)u^{\mu}u^{\nu}-pg^{\mu\nu},\quad p=p(n,\varepsilon),$ (10)
is represented by the following equations:
$3H^{2}$ $=\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}+\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4}$ (11a)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(0)}$ $=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(0)},$ (11b)
$\dot{\varepsilon}_{\scriptscriptstyle(0)}$ $=-3H\varepsilon_{\scriptscriptstyle(0)}(1+w),$ $w$ $:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)},$ (11c)
$\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (11d)
The $G_{0i}$ constraint equations and the $G_{ij}$ dynamical equations with $i\neq j$ are satisfied identically. The $G_{ii}$ dynamical equations are identical to the time derivative of the $G_{00}$ constraint equation (Friedmann equation) (11a). Therefore, the $G_{0i}$ and $G_{ij}$ equations can be omitted. The cosmological constant $\Lambda$, the gravitational constant $G_{\text{N}}$, and the speed of light $c$ are fundamental constants.
The overdot denotes differentiation with respect to $x^{0}=ct$. The Hubble function $H$ is defined by $H:=\dot{a}/a$. For flrw universes, the Hubble function is given by $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, where $\theta_{\scriptscriptstyle(0)}$ is the background value of the expansion scalar $\theta:=u^{\mu}{}_{;\mu}$, where $u^{\mu}$ is the four-velocity $u^{\mu}:=c^{-1}U^{\mu}$, normalized to unity, $u^{\mu}u_{\mu}=1$. A semicolon denotes covariant differentiation with respect to the background metric tensor $g_{{\scriptscriptstyle(0)}\mu\nu}$ (8). The equations (11c) and (11d) represent the energy density conservation law and the particle number conservation law, respectively. The variable $w$ in Eq. (11c) is an abbreviation for $p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, and it does not represent an equation of state.
The spatial parts of the background Riemann tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}$, the Ricci tensor ${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}$, and its contraction ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ are given by
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}jkl}=\tilde{R}^{i}{}_{jkl}=K(\delta^{i}{}_{k}\tilde{g}_{jl}-\delta^{i}{}_{l}\tilde{g}_{jk}),$ (12a)
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}j}=-\dfrac{2K}{a^{2}}\delta^{i}{}_{j},\quad{{}^{3}\!}R_{\scriptscriptstyle(0)}=-\dfrac{6K}{a^{2}},$ (12b)
where ${{}^{3}\!}R_{\scriptscriptstyle(0)}$ is the spatial curvature. The value of $K$ determines the nature of the flrw universe. The universe is open for K=1K=-1, flat for $K=0$, and closed for $K=+1$.
VI.2 Linearized Equations
The system of linearized Einstein equations and conservation laws for closed, flat, and open flrw universes is given by
$H\dot{h}^{k}{}_{k}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\kappa\varepsilon_{\scriptscriptstyle(1)},$ (13a)
$\dot{h}^{k}{}_{k\mid i}-\dot{h}^{k}{}_{i\mid k}=2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i},$ (13b)
$\ddot{h}^{i}{}_{j}+3H\dot{h}^{i}{}_{j}\,$+
[…]
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})+\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)}=0,$ (13d)
[…]
$\qquad g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{\scriptscriptstyle(1)}=0,$ (13e)
$\dot{n}_{\scriptscriptstyle(1)}+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}=0.$ (13f)
In these equations, the perturbed metric tensor is given by $h_{\mu\nu}:=-g_{{\scriptscriptstyle(1)}\mu\nu}$ and $h^{\mu\nu}:=g^{\mu\nu}_{{\scriptscriptstyle(1)}}$. In light of the use of synchronous coordinates, it follows that $h_{00}=0$ and $h_{0i}=0$. Raising and lowering of indices is performed by the spatial background metric tensor as defined in expression (8), namely, $h^{i}{}_{j}=g^{ik}_{\scriptscriptstyle(0)}h_{kj}$ where $g^{ik}_{\scriptscriptstyle(0)}=-\tilde{g}^{ik}/a^{2}$. A vertical bar denotes covariant differentiation with respect to the spatial background metric tensor $g_{{\scriptscriptstyle(0)}ij}=-a^{2}\tilde{g}_{ij}$. In the context of flrw universes, where $\Gamma^{k}_{{\scriptscriptstyle(0)}ij}=\tilde{\Gamma}^{k}{}_{ij}$, the operations of taking the time derivative and the covariant derivative are shown to commute. Equation (13a) represents the linearized energy density constraint equation (linearized Friedmann equation), while Eq. (13b) corresponds to the linearized momentum constraint equation. The linearized dynamical equations are presented in Eqs. (13c). Equations (13d) and (13e) represent the linearized energy density conservation law and the linearized momentum conservation law, respectively. Finally, the linearized particle number density conservation law is given by Eq. (13f).
Pressure perturbations, spatial curvature perturbations, and the covariant divergence of the fluid velocity play pivotal roles in the evolution of density perturbations. Therefore, these quantities must be included in the perturbation theory. The next three subsections will derive expressions for these quantities.
VI.2.1 Pressure Perturbation
The gauge-dependent quantity $p_{{\scriptscriptstyle(1)}}$, which is associated with pressure (2), is given by
$p_{{\scriptscriptstyle(1)}}=p_{n}n_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}},$ (14a)
$p_{n}:=\left(\dfrac{\partial p}{\partial n}\right)_{\varepsilon},\quad p_{\varepsilon}:=\left(\dfrac{\partial p}{\partial\varepsilon}\right)_{n}.$ (14b)
The quantity $p^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ representing the pressure perturbation will be derived in Sec. XII on thermodynamics.
VI.2.2 Spatial Curvature Perturbation
Lifshitz’s expression for the perturbed connection coefficient presented in Ref. [2], Eq. (I.3) and Ref. [42], Eq. (10.9.1) is a tensor, which is given by
$\Gamma^{k}_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j}+h_{lj\mid i}-h_{ij\mid l}).$ (15)
The contracted Palatini identity as defined in Ref. [2], Eq. (I.5) and Ref. [42], Eq. (10.9.2) is given by
${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=\Gamma^{k}_{{\scriptscriptstyle(1)}ij\mid k}-\Gamma^{k}_{{\scriptscriptstyle(1)}ik\mid j}.$ (16)
By combining (15) and (16), the following expression for the perturbed spatial Ricci tensor can be derived (see Ref. [42], Eq. (10.9.3):
${{}^{3}\!}R_{{\scriptscriptstyle(1)}ij}=-\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}(h_{li\mid j\mid k}+h_{lj\mid i\mid k}-h_{ij\mid l\mid k}-h_{lk\mid i\mid j}).$ (17)
By raising the index $i$ and using (12b), one arrives at
${{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}j}:=$ $\,(g^{ik}\,{{}^{3}\!}R_{kj})_{\scriptscriptstyle(1)}=g^{ik}_{\scriptscriptstyle(0)}\,{{}^{3}\!}R_{{\scriptscriptstyle(1)}kj}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}$
[…]
$+\tfrac{1}{2}g^{kl}_{\scriptscriptstyle(0)}h^{i}{}_{j\mid k\mid l}+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{i}{}_{j}$. (18)
By using $g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid j\mid k}=g_{\scriptscriptstyle(0)}^{ij}h^{k}{}_{i\mid k\mid j}$, one arrives at the contraction of the perturbed spatial Ricci tensor:
${{}^{3}\!}R_{\scriptscriptstyle(1)}:={{}^{3}\!}R^{k}_{{\scriptscriptstyle(1)}k}=g_{\scriptscriptstyle(0)}^{ij}(h^{k}{}_{k\mid i\mid j}-h^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}h^{k}{}_{k}$. (19)
In Sec. VII it will be demonstrated that the expression (19) represents the local perturbation to the spatial curvature ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}$ induced by local density perturbations. The quantity ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ will be used in Sec. IX to eliminate the metric perturbations and their spatial derivatives from the momentum constraint equation (13b). It is incorporated into the perturbation theory using Eq. (127).
VI.2.3 Covariant Divergence of the Fluid Velocity
[…]
$\theta_{\scriptscriptstyle(1)}=\vartheta_{\scriptscriptstyle(1)}-\tfrac{1}{2}\dot{h}^{k}{}_{k},\quad\vartheta_{\scriptscriptstyle(1)}:=u^{k}_{{\scriptscriptstyle(1)}\mid k},$ (20)
where $\vartheta_{\scriptscriptstyle(1)}$ is the covariant divergence of the fluid vector. The quantity $\theta_{{\scriptscriptstyle(1)}}$ will be used in Sec. IX to eliminate the time derivative of the metric perturbation from the energy density constraint equation (13a) in favor of $\vartheta_{{\scriptscriptstyle(1)}}$. The latter quantity is incorporated into the perturbation theory using Eq. (132).
VII Decomposition of Spatial Tensors and Vectors
The system of linearized Einstein equations and conservation laws (13) is not suitable for deriving the perturbation theory because it contains solutions related to and unrelated to density perturbations. Therefore, our goal is to transform the system (13) into one that is exclusively linked to density perturbations and excludes all other solutions. To that end, we will use two key theorems to break the linearized equations down into three independent systems. These theorems are presented in the next two subsections.
VII.1 Tensor Decomposition
Refs. [45, 46, 47] demonstrate that any rank-two symmetric spatial tensor ${\cal{T}}_{ij}$ can be decomposed into three irreducible components in a covariant manner:
${\cal{T}}^{i}{}_{j}={\cal{T}}^{i}_{\parallel\,j}+{\cal{T}}^{i}_{\perp\,j}+{\cal{T}}^{i}_{\ast\,j}.$ (21)
The components have the following properties:
${\cal{T}}^{k}_{\perp\,k}=0,\quad{\cal{T}}^{k}_{\ast\,\,k}=0,\quad{\cal{T}}^{k}_{\ast\;i\mid k}=0,\quad{\cal{T}}^{i}_{\parallel\,j}=\phi\delta^{i}{}_{j}+\zeta^{\mid i}{}_{\mid j}$, (22)
where $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$ are two independent potentials. Consequently, the component $h^{i}_{\parallel\,j}$ of the perturbed metric can be expressed in terms of the two independent potentials as follows:
[…]
The factor $2/c^{2}$ is included in view of the Newtonian limit. By substituting (23) into (19), the following result is obtained:
${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}=\dfrac{2}{c^{2}}\Bigl[2\phi^{|i}{}_{|i}$ $+\zeta^{|k}{}_{|k}{}^{|i}{}_{|i}-\zeta^{|k|i}{}_{|k|i}$
[…]
and for the perturbed expansion (20) it is found
[…]
The expression (VII.1) reveals that two potentials, $\phi(t,\bm{x})$ and $\zeta(t,\bm{x})$, are associated with cosmological density perturbations in open and closed flrw universes. For a flat flrw universe, however, density perturbations are characterized by one potential, $\phi(t,\bm{x})$. This is because, for scalar perturbations, the quantity $\theta_{{\scriptscriptstyle(1)}}$ can be eliminated from expressions (6a) and Eqs. (37), as was done in (122) and (123).
As will be demonstrated in Sec. X, in the Newtonian limit, the potential $\phi(t,\bm{x})$ becomes identical to the Newtonian potential, $\phi(\bm{x})$.
VII.2 Vector Decomposition
Since $\bm{u}_{\scriptscriptstyle(1)}$ is a vector in three-space, it can be decomposed using the Helmholtz decomposition as follows:
$\bm{u}_{{\scriptscriptstyle(1)}}=\bm{u}_{{\scriptscriptstyle(1)}\parallel}+\bm{u}_{{\scriptscriptstyle(1)}\perp},$ (26)
where the components $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ and $\bm{u}_{{\scriptscriptstyle(1)}\perp}$ have the properties
$\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel},\quad\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{{\scriptscriptstyle(1)}\perp}.$ (27)
[…]
Since $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is irrotational, it can be supplemented with the gradient of an arbitrary function. Given that the system of equations (13) is invariant under the gauge transformation (3), where the gauge function $\xi^{\mu}(t,\bm{x})$ is given by (9), it follows that
$u^{\prime}_{{\scriptscriptstyle(1)}\parallel i}=u_{{\scriptscriptstyle(1)}\parallel i}+\psi_{|i},$ (28)
is also a solution to the system (13). As a result, the gauge modes associated with $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ are given by
$\hat{u}_{{\scriptscriptstyle(1)}\parallel i}=\psi_{|i},\quad\hat{u}^{i}_{{\scriptscriptstyle(1)}\parallel}=g_{{\scriptscriptstyle(0)}}^{ik}\psi_{|k}=-\dfrac{1}{a^{2}}\tilde{g}^{ik}\psi_{|k}.$ (29)
This result will be used in Sec. X on the Newtonian limit.
VIII Decomposition of the Linearized Equations
The system of equations (13) is split into three independent systems by using the decomposition (21) with the properties (22) for $h_{ij}$ and ${{}^{3}\!}R_{ij}$, as well as the decomposition (26) with the properties (27) for $\bm{u}_{{\scriptscriptstyle(1)}}$. The solutions to these systems are conventionally designated as tensor perturbations $\ast$, vector perturbations $\perp$, and scalar perturbations $\parallel$.
In the following three subsections, it is demonstrated that scalar perturbations are the only ones associated with density perturbations $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
VIII.1 Tensor Perturbations
From the expressions (19) and (22), it follows that ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\ast}=0$, and Eq. (13b) yields $\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$}, so that, with (20), $\theta_{{\scriptscriptstyle(1)}}=0$. Equation (13a) yields $\varepsilon_{{\scriptscriptstyle(1)}}=0$}. From (6a), it follows that $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$}. Given that $\theta_{{\scriptscriptstyle(1)}}=0$, Eqs. (11d) and (13f) are identical, so $n_{{\scriptscriptstyle(1)}}=0$. This, in turn, implies with (6a), that $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, the result obtained from either (13d) or the contraction of (13c) is that $p_{{\scriptscriptstyle(1)}}=0$. Consequently, the evolution equations for tensor perturbations are as follows:
$\ddot{h}^{i}_{\ast\,j}+3H\dot{h}^{i}_{\ast\,j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\ast j}=0.$ (30)
Given the form of these equations, tensor perturbations are typically referred to as gravitational waves.
VIII.2 Vector Perturbations
Expressions (19) and (22), imply that, for vector perturbations, we have ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=-h^{kj}_{\perp|k|j}$. Raising the index $i$ of the momentum constraint equation (13b) with $g^{ij}_{\scriptscriptstyle(0)}$, and subsequently taking the covariant derivative with respect to the index $j$, one finds, using that $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$,
$\dot{h}^{kj}_{\perp{\mid k\mid j}}+2Hh^{kj}_{\perp{\mid k\mid j}}=-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{j}_{{\scriptscriptstyle(1)}\perp\mid j}.$ (31)
The right-hand side of this equation vanishes, as can be seen from (27). Consequently, the covariant divergence of the momentum constraint equation (31) can only be satisfied if ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\perp}=0$. The consequences are as follows.
Equation (13a) implies that $\varepsilon_{{\scriptscriptstyle(1)}}=0$}, and (20) yields $\theta_{{\scriptscriptstyle(1)}}=0$. The latter identity shows that Eq. (13f) is identical to the background equation (11d), so $n_{{\scriptscriptstyle(1)}}=0$}. From (6a) we find that $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$. Finally, either (13d) or the contraction of (13c) implies that $p_{{\scriptscriptstyle(1)}}=0$. Therefore, the system of equations for vector perturbations is given by
$\!\dot{h}^{k}_{\perp i\mid k}+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}\perp i}=0,$ (32a)
$\!\ddot{h}^{i}_{\perp j}+3H\dot{h}^{i}_{\perp j}+2\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(1)}\perp j}=0,$ (32b)
$\!\frac{1}{c}\frac{{\text{d}}}{{\text{d}}t}\Bigl(\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}\Bigr)+5H\varepsilon_{\scriptscriptstyle(0)}(1+w)u^{i}_{{\scriptscriptstyle(1)}\perp}=0.$ (32c)
According to (27), vector perturbations are also referred to as rotational perturbations.
VIII.3 Scalar Perturbations
In this particular instance one has ${{}^{3}\!}R_{{\scriptscriptstyle(1)}\parallel}\neq 0$, as follows from expression (VII.1). By taking the covariant derivative of the momentum constraint equation (13b) with respect to the index $j$ and subsequently substituting the expression (23), the following result is obtained:
$2\dot{\phi}_{\mid i\mid j}+\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}=\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)u_{{\scriptscriptstyle(1)}i\mid j}.$ (33)
By interchanging the indices $i$ and $j$ and subtracting the result from Eqs. (33) one finds
$\dot{\zeta}^{\mid k}{}_{\mid i\mid k\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid k\mid i}=-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}),$ (34)
where it is used that $\dot{\phi}_{\mid i\mid j}=\dot{\phi}_{\mid j\mid i}$ and $\dot{\zeta}^{\mid k}{}_{\mid k\mid i\mid j}=\dot{\zeta}^{\mid k}{}_{\mid k\mid j\mid i}$. Rearranging the second-order covariant derivatives converts Eq. (34) into a form with only commutators on the left-hand side:
[…]
$\qquad(\dot{\zeta}^{\mid k}{}_{\mid i\mid j}-\dot{\zeta}^{\mid k}{}_{\mid j\mid i})_{\mid k}\,=$
$\qquad-\kappa c^{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)(u_{{\scriptscriptstyle(1)}i\mid j}-u_{{\scriptscriptstyle(1)}j\mid i}).$ (35)
The commutators of second-order covariant derivatives (see Ref. [42], Sec. 6.5) are given by
$A^{i}{}_{j\mid p\mid q}-A^{i}{}_{j\mid q\mid p}=A^{i}{}_{k}\,{{}^{3}\!}R^{k}_{{\scriptscriptstyle(0)}jpq}-A^{k}{}_{j}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq},$ (36a)
$B^{i}{}_{\mid p\mid q}-B^{i}{}_{\mid q\mid p}=B^{k}\,{{}^{3}\!}R^{i}_{{\scriptscriptstyle(0)}kpq}.$ (36b)
Upon substituting the background Riemann tensor (12a), the left-hand side of Eq. (35) vanishes identically. This implies that $\bm{\tilde{\nabla}}\bm{\times}\bm{u}_{\scriptscriptstyle(1)}=\bm{0}$, so that only the component $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ remains. Since $\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}\neq 0$ implies that $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}\neq 0$, it can thus be concluded that only scalar perturbations are coupled to density perturbations.
IX Evolution Equations for Scalar Perturbations
As demonstrated in the preceding section, the evolution of $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is determined by the background equations (11) and the equations that describe the evolution of scalar perturbations. In this section, the findings from the previous section are used to reformulate the system (13) into a new system of equations that exclusively describes the evolution of scalar perturbations. Given the subsequent focus on density perturbations, the subscript $\parallel$ is omitted. The evolution equations for scalar perturbations are as follows:
$2H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+\kappa\varepsilon_{\scriptscriptstyle(1)},$ (37a)
$\!\!{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$
$\qquad\;\,+2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}-\tfrac{2}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)}),$ (37b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}=-3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})-\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)},$ (37c)
$\dot{\vartheta}_{\scriptscriptstyle(1)}=-H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}-\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}p_{\scriptscriptstyle(1)}}{a^{2}},$ (37d)
$\dot{n}_{\scriptscriptstyle(1)}=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (37e)
[…]
The symbol $\tilde{\nabla}^{2}$ denotes the Laplace-Beltrami operator with respect to the metric $\tilde{g}_{ij}(\bm{x})$ of three-dimensional subspaces of constant time:
$\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}:=\tilde{g}^{ij}p_{{\scriptscriptstyle(1)}\mid i\mid j},\quad g^{ij}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|i|j}=-\dfrac{\tilde{\nabla}^{2}p_{{\scriptscriptstyle(1)}}}{a^{2}}.$ (38)
The parameter $\beta$ is defined as follows:
$\beta^{2}:=\dfrac{\dot{p}_{{\scriptscriptstyle(0)}}}{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}.$ (39)
Using the expression $\dot{p}_{\scriptscriptstyle(0)}=p_{n}\dot{n}_{\scriptscriptstyle(0)}+p_{\varepsilon}\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and eliminating the time derivatives of $\varepsilon_{\scriptscriptstyle(0)}$ and $n_{\scriptscriptstyle(0)}$ with the aid of the conservation laws (11c) and (11d), the following result is obtained:
$\beta^{2}=p_{\varepsilon}+\dfrac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (40)
where $p_{\varepsilon}$ and $p_{n}$ are given by (14b).
We will now derive Eqs. (37). To begin, the term $\dot{h}^{k}{}_{k}$ is eliminated from Eq. (13a) using the expression (20). This results in the algebraic energy density constraint equation (37a).
Given that only $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$ is associated with scalar perturbations, it is possible to simplify the momentum constraint equation (13b) and the momentum conservation law (13e) by using the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}:=\bm{\tilde{\nabla}}\bm{\cdot}\bm{u}_{{\scriptscriptstyle(1)}\parallel}$}, (20), as an alternative to $\bm{u}_{{\scriptscriptstyle(1)}\parallel}$. By multiplying both sides of Eq. (13b) by $g^{ij}_{\scriptscriptstyle(0)}$ and taking the covariant derivative with respect to the index $j$, one finds
$g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})=2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}.$ (41)
The left-hand side of Eq. (41) will appear as a component of the time derivative of the local spatial curvature perturbation, ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Indeed, differentiating (19) with respect to time, and using the relation $\dot{g}^{ij}_{\scriptscriptstyle(0)}=-2Hg^{ij}_{\scriptscriptstyle(0)}$ and (11b), yields the following result:
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}=-2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}+g_{\scriptscriptstyle(0)}^{ij}(\dot{h}^{k}{}_{k\mid i\mid j}-\dot{h}^{k}{}_{i\mid k\mid j})+\tfrac{1}{3}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\dot{h}^{k}{}_{k}$. (42)
By combining (41) and (42) and using (20) to eliminate $\dot{h}^{k}{}_{k}$, one obtains Eq. (37b). Consequently, the $G^{0}_{{\scriptscriptstyle(1)}i}$ momentum constraint equation (13b) has been reformulated as a first-order ordinary differential equation (37b) for the perturbed spatial Ricci scalar (VII.1) expressed in the potentials.
The momentum conservation law (13e) is reformulated by first performing the differentiation with respect to time. Then, we use Eq. (11c) to eliminate $\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$. Next, dividing the result by $\varepsilon_{{\scriptscriptstyle(0)}}(1+w)$ yields
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}+H(2-3w)u^{i}_{{\scriptscriptstyle(1)}}+\dfrac{\dot{w}}{1+w}u^{i}_{{\scriptscriptstyle(1)}}-\dfrac{g^{ik}_{\scriptscriptstyle(0)}p_{{\scriptscriptstyle(1)}\mid k}}{\varepsilon_{{\scriptscriptstyle(0)}}(1+w)}=0.$ (43)
This equation can be simplified by eliminating the time derivative of $w$. From the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$ and the energy conservation law (11c) one obtains
$\dot{w}=3H(1+w)(w-\beta^{2}).$ (44)
Upon substituting this expression into Eq. (43), the momentum conservation law becomes
$\dot{u}^{i}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})u^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}=0.$ (45)
Taking the covariant divergence of (45) with respect to the spatial background metric tensor (8), and using (20) and (38) yields Eq. (37d).
As demonstrated in Appendix B, the dynamical equation (13c) is redundant since the system (37) comprises the constraint equations and conservation laws. Thus, the derivation of Eqs. (37) has been completed.
X Newtonian Limit
As demonstrated in Sec. IV, the gauge-invariant quantities, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, which are intended to describe density perturbations, can only be defined in one way. This section proves that the quantities $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ represent perturbations to energy and particle number densities, respectively.
In Newtonian gravity, space is Euclidean. Consequently, the Newtonian limit can only be realized in a flat flrw universe. In this case, the Newtonian limit is reached when gravity is weak and the velocities of the particles are negligible compared to the speed of light. These conditions are met in a linear perturbation theory and when pressure approaches zero.
In a flat flrw universe where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, covariant derivatives are equal to ordinary derivatives. This simplifies the expression (VII.1) to
${{}^{3}\!}R_{{\scriptscriptstyle(1)}}=\dfrac{4}{c^{2}}\phi^{\mid k}{}_{\mid k}=-\dfrac{4}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}.$ (46)
The symbol $\nabla^{2}$ represents the conventional Laplace operator. By substituting the expression (46) into the perturbation equations (37) and using that $H:=\dot{a}/a$, we obtain the result:
$-H\vartheta_{\scriptscriptstyle(1)}$ $=-\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)},\quad\kappa=8\pi G_{\text{N}}/c^{4},$ (47a)
$\dfrac{1}{c^{2}}\dfrac{\nabla^{2}\dot{\phi}}{a^{2}}$ $=-\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)},\quad w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}},$ (47b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $=-3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})-\varepsilon_{\scriptscriptstyle(0)}(1+w)\theta_{\scriptscriptstyle(1)},$ (47c)
$\dot{\vartheta}_{\scriptscriptstyle(1)}$ $=-H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}-\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\nabla^{2}p_{\scriptscriptstyle(1)}}{a^{2}},$ (47d)
$\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},\quad\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (47e)
To derive the $G^{0}_{{\scriptscriptstyle(1)}0}$ constraint equation (47a) from Eq. (37a), the function $\varepsilon_{{\scriptscriptstyle(1)}}$ must first be eliminated using the expression (6a). The result is
$2H(\theta_{{\scriptscriptstyle(1)}}-\vartheta_{{\scriptscriptstyle(1)}})=-\dfrac{2}{c^{2}}\dfrac{\nabla^{2}\phi}{a^{2}}+\kappa\left(\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\dfrac{\dot{\varepsilon}_{{\scriptscriptstyle(0)}}}{\dot{\theta}_{{\scriptscriptstyle(0)}}}\theta_{{\scriptscriptstyle(1)}}\right).$ (48)
Setting ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$ in the Friedmann equation (11a) and differentiating the result with respect to time, and using the relation $H=\tfrac{1}{3}\theta_{\scriptscriptstyle(0)}$, we obtain $\dot{\varepsilon}_{\scriptscriptstyle(0)}/\dot{\theta}_{\scriptscriptstyle(0)}=2H/\kappa$. Substituting this expression into Eq. (48) yields the desired result, Eq. (47a). The Newtonian limit can now be derived using the system (47).
In a universe filled with a nonrelativistic fluid, we have $w\ll 1$ and $\beta^{2}\ll 1$. Neglecting the quantities $w$ and $\beta^{2}$ with respect to the constants of order one in the momentum conservation law (45), yields
$\dot{u}^{i}_{\scriptscriptstyle(1)}+2Hu^{i}_{\scriptscriptstyle(1)}-\frac{g^{ik}_{{\scriptscriptstyle(0)}}p_{{\scriptscriptstyle(1)}|k}}{\varepsilon_{\scriptscriptstyle(0)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}-\dfrac{p_{{\scriptscriptstyle(1)}|j}}{\varepsilon_{\scriptscriptstyle(0)}}=0.$ (49)
The second equation is derived by lowering the index $i$ in the first equation using the background metric $g_{{\scriptscriptstyle(0)}ij}$ and the relation $\dot{g}_{{\scriptscriptstyle(0)}ij}=2Hg_{{\scriptscriptstyle(0)}ij}$. Accordingly, the evolution of density perturbations is highly dependent upon the presence of pressure gradients, represented by $p_{{\scriptscriptstyle(1)}|j}$, where $p_{{\scriptscriptstyle(1)}}$ is given by (14). In a pressureless fluid, i.e., when $w=0$ and $\beta=0$, however, the result is
$\dot{u}^{i}_{{\scriptscriptstyle(1)}}+2Hu^{i}_{{\scriptscriptstyle(1)}}=0,\quad\dot{u}_{{\scriptscriptstyle(1)}j}=0.$ (50)
The second equation has solutions that are only functions of the spatial coordinates. Since these equations are part of the system (13) that is invariant under the gauge transformation (3) with $\xi^{\mu}(t,\bm{x})$ given by (9), it can be inferred that the gauge modes given by (29), are the only solutions to equations (50). Therefore, it can be concluded that
$p\rightarrow 0\quad\Rightarrow\quad u_{{\scriptscriptstyle(1)}i}^{{\textnormal{phys}}}(t,\bm{x})\rightarrow 0.$ (51)
In other words, in the absence of pressure, local pressure gradients $p_{{\scriptscriptstyle(1)}|i}$ are nonexistent, so local fluid flows are impossible. Thus, in the Newtonian limit, density perturbations are independent of time, that is, they are static. The remaining part of this section examines the consequences of the limit (51).
Since the components (51) of the fluid velocity are zero, only the gauge modes $\hat{u}_{{\scriptscriptstyle(1)}i}=\psi_{|i}$ remain. We can set these gauge modes to zero without losing any physical information. Substituting $\psi_{|i}=0$ into the expressions (9) yields $\xi^{0}=C$ and $\xi^{i}=\chi^{i}$, implying that the relativistic gauge transformation (3) reduces to the gauge transformation of Newtonian gravity:
$t\rightarrow t^{\prime}=t-C,\quad x^{i}\rightarrow x^{\prime i}=x^{i}-\chi^{i}(\bm{x}).$ (52)
Thus, we arrive at the well-known fact that in Newtonian gravity, we can freely select the spatial coordinates and shift the time coordinate at will, and that space and time are independent of each other.
Since both $\hat{u}_{{\scriptscriptstyle(1)}i}=0$ and $u_{{\scriptscriptstyle(1)}i}^{{\textnormal{phys}}}=0$ we find from (20) that $\vartheta_{\scriptscriptstyle(1)}=0$. By substituting the latter into the system (47) we obtain
$\nabla^{2}\phi(t,\bm{x})$ $=\dfrac{4\pi G_{\text{N}}}{c^{2}}a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x}),$ (53a)
$\nabla^{2}\dot{\phi}(t,\bm{x})$ $=0,$ (53b)
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $=-3H\varepsilon_{\scriptscriptstyle(1)}-\varepsilon_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)},$ (53c)
$\dot{n}_{\scriptscriptstyle(1)}$ $=-3Hn_{\scriptscriptstyle(1)}-n_{\scriptscriptstyle(0)}\theta_{\scriptscriptstyle(1)}.$ (53d)
The constraint equations (53a) and (53b) imply
$a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=a^{2}(t_{0})\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t_{0},\bm{x}).$ (54)
This equation applies at any given time $t_{0}$. From the Friedmann equation (11a), it follows that, for ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, one may take $a(t_{0})=1$, so that
$a^{2}(t)\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x}).$ (55)
This expression is consistent with the limit (51) that results in static density perturbations. Combining (53a), (53b), and (55) yields
$\nabla^{2}\phi(\bm{x})=4\pi G_{\text{N}}\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{c^{2}},$ (56)
which is the Poisson equation of Newtonian gravity.
Finally, the meaning of $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is established. For a universe filled with a pressureless fluid, the background equations (11) are given by
$3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}+\Lambda,\quad\dot{\varepsilon}_{\scriptscriptstyle(0)}=-3H\varepsilon_{\scriptscriptstyle(0)},\quad\dot{n}_{\scriptscriptstyle(0)}=-3Hn_{\scriptscriptstyle(0)}.$ (57)
Consequently, the conservation laws imply that $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$. As can be verified by substitution, Eqs. (53c)–(53d) yield solely gauge modes as solutions: specifically, $\hat{\varepsilon}_{\scriptscriptstyle(1)}=C\dot{\varepsilon}_{\scriptscriptstyle(0)}$, $\hat{n}_{{\scriptscriptstyle(1)}}=C\dot{n}_{{\scriptscriptstyle(0)}}$, and $\hat{\theta}_{{\scriptscriptstyle(1)}}=C\dot{\theta}_{{\scriptscriptstyle(0)}}$, where $C$ is an infinitesimal constant of the gauge transformation (52) of Newtonian gravity. This, combined with $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}$ yields $\varepsilon_{{\scriptscriptstyle(1)}}=n_{{\scriptscriptstyle(1)}}mc^{2}$. Combining the latter two relations with the definitions (6a) one finds
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})mc^{2}.$ (58)
This is the well-known special relativistic relation between rest mass and energy.
The equations (53c) and (53d) represent the residual form of the conservation laws (47c) and (47e) in the Newtonian limit. Since Eqs. (53c) and (53d) only have gauge modes as solutions, they clearly have no physical significance. These equations are decoupled from the physical constraint equations (53a) and (53b) and thus do not form part of Newtonian gravity. Therefore, the potential $\zeta$ which appears in $\theta_{{\scriptscriptstyle(1)}}$ (25) is no longer relevant. This is consistent with the comment just below (25). Thus, in the Newtonian limit, only the Poisson equation (56) and the special relativistic relation (58) remain.
We now come to a final conclusion. The analysis above demonstrates that, in the Newtonian limit, only the Poisson equation (56) and the relation (58) remain. Therefore, $\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ and $n_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ (6a) represent the perturbations in energy and particle number density, respectively, in Newtonian gravity. Consequently, they also represent perturbations in energy and particle number density in general relativity. Finally, we note that, since $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$ (6b), density perturbations do not affect the expansion of the universe.
In the next subsections, relative density perturbations in the Newtonian limit are discussed, and the perturbation theory based on Newtonian gravity is examined in the context of relativity theory.
X.1 Relative Density Perturbation in the Newtonian Limit
The expression (55) implies that
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})}{a^{2}(t)}.$ (59)
Since Eq. (11c) with $w=0$ implies $\varepsilon_{{\scriptscriptstyle(0)}}\propto a^{-3}$, it follows that the relative density perturbation $\delta_{\varepsilon}$ evolves according to
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{{\scriptscriptstyle(0)}}(t)}\propto\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})a(t).$ (60)
Thus, due to the expansion of the universe, the relative density perturbation $\delta_{\varepsilon}$ for a static density perturbation $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(\bm{x})$ increases proportionally to the scale factor $a(t)$ of the universe. In other words, a static density perturbation becomes relatively denser when viewed against the decreasing energy density $\varepsilon_{{\scriptscriptstyle(0)}}$ in the background universe.
X.2 Newtonian Gravity
As demonstrated in Appendix C, the relative perturbation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ is gauge-dependent. The gauge mode associated with $\delta$ is given by (152a). For $w\ll 1$, it is a solution to the homogeneous part of Eq. (157a). This gauge mode does not disappear from the solution because $H\neq 0$ in a nonempty universe. Consequently, the quantity $\delta$ depends on the choice of coordinates and therefore has no physical meaning. As can be seen from Eqs. (57), $H\neq 0$ even in the Newtonian limit. Therefore, $\delta$ must also depend on the choice of coordinates in that limit and consequently has no physical significance. This is consistent with the fact that, in the Newtonian limit, the relativistic gauge transformation (3) converges to the Newtonian gauge transformation (52). Given that the system (151) reduces to the system (157) for a nonrelativistic fluid, it can be concluded that the system (157) is inadequate for examining the evolution of density perturbations.
The existing literature derives Eq. (156), from Newtonian gravity under two conditions: the expansion of the universe must be considered and the equation of state must be nonrelativistic, i.e., $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}\ll 1$ and $\beta^{2}\ll 1$. See, for example, Ref. [7], Sec. 6.2, Ref. [42], Sec. 15.9, and Ref. [48], Eq. (3.17). However, the gauge transformation (52) shows that $\delta$ depends on the choice of coordinates in Newtonian gravity. Therefore, it also lacks physical significance in Newtonian gravity. Furthermore, a comparison of the conventional equation (156) and the exact equation (157a) reveals that equation (156) is incomplete. Consequently, the conventional evolution equation (156) — whether derived from general relativity or Newtonian gravity modified to account for the expansion of the universe — also has gauge-dependent solutions. These findings indicate that Newtonian gravity is inadequate in accurately describing the evolution of cosmological density perturbations. Therefore, there exists no distinction between sub- and super-horizon density perturbations.
XI Relativistic Cosmological Perturbation Theory
As demonstrated in Sec. VIII, the evolution equations for scalar perturbations (37) when combined with the background equations (11) describe the evolution of energy density and particle number density perturbations. In Appendix A, the derivation of evolution equations for their corresponding relative density perturbations defined by
$\delta_{\varepsilon}(t,\bm{x}):=\dfrac{\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{\varepsilon_{\scriptscriptstyle(0)}(t)},\quad\delta_{n}(t,\bm{x}):=\dfrac{n^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})}{n_{\scriptscriptstyle(0)}(t)},$ (61)
will be presented. The result is a system of evolution equations for the relative density perturbations $\delta_{\varepsilon}$ and $\delta_{n}$, in closed, flat, and open flrw universes, given by
$\ddot{\delta}_{\varepsilon}+b_{1}\dot{\delta}_{\varepsilon}+b_{2}\delta_{\varepsilon}$ $=b_{3}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right),$ (62a)
$\frac{1}{c}\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right)$ $=\frac{3Hn_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\left(\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}\right).$ (62b)
The background equations (11) and an equation of state (2) for the pressure determine the coefficients $b_{1}$, $b_{2}$, and $b_{3}$. These coefficients are given by
$b_{1}=$ $\,\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{H}-2\dfrac{\dot{\beta}}{\beta}-H(2+6w+3\beta^{2})$
$+{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\dfrac{1}{3H}+\dfrac{2H(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right),$ (63a)
$b_{2}=$ $-\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)$
$+H^{2}\left(1-3w+6\beta^{2}(2+3w)\right)$
$+6H\dfrac{\dot{\beta}}{\beta}\left(w+\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
$-{{}^{3}\!}R_{\scriptscriptstyle(0)}\left(\tfrac{1}{2}w+\dfrac{H^{2}(1+6w)(1+3\beta^{2})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)$
[…]
$b_{3}=$ $\,\Biggl[\dfrac{-18H^{2}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}\Biggl(\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
$+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr)+p_{n}\Biggr]$
[…]
In these expressions the partial derivatives of the pressure, $p_{\varepsilon}$ and $p_{n}$, are defined by (14b). The second-order partial derivatives are defined by $p_{nn}:={\partial^{2}p}/{\partial n^{2}}$ and $p_{\varepsilon n}:={\partial^{2}p}/{\partial\varepsilon\,\partial n}$.
The local curvature perturbation, ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$, caused by a local density perturbation as well as the divergence, ${\vartheta}_{{\scriptscriptstyle(1)}}$, of the local fluid velocity caused by local pressure gradients, are incorporated into system (62) via expressions (127) and (132), respectively.
The system of equations (62) allows us to study the evolution of the relative density perturbations (61) in closed $K=+1$, flat $K=0$, and open K=1K=-1 flrw universes, where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=-6K/a^{2}$.
Sec. XIII will demonstrate that $p_{n}<0$ before and after the decoupling of radiation and matter. For $p_{n}>0$, however, the coefficient in Eq. (62b) is positive, causing $\delta_{n}-\delta_{\varepsilon}/(1+w)$ to grow exponentially, which is not physically plausible.
In the system (62), Eq. (62a) is the evolution equation for energy density perturbations. Since $p_{n}<0$ for an equation of state $p=p(n,\varepsilon)$, the right-hand side of Eq. (62a) is nonzero. This indicates that density perturbations are nonadiabatic. Density perturbations are adiabatic when the source term of Eq. (62a) vanishes, i.e., when
$\delta_{n}-\frac{\delta_{\varepsilon}}{1+w}=0.$ (64)
Therefore, Eq. (62b) can be regarded as the entropy equation. An explanation is provided by Sec. XII.1.
Since $p_{n}<0$, the entropy equation (62b) shows that
$\left|\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right|\rightarrow 0.$ (65)
This means that relative perturbations in energy density are coupled with those in particle number density. Since general relativity is indifferent to the composition of matter, this is true for cdm as well as for ordinary matter, where it is assumed that cdm interacts only via gravity. Therefore, the contraction of cdm before decoupling was impossible. This precludes the formation of potential wells into which ordinary matter could fall after decoupling to form structures. Consequently, the initiation of structure formation by cdm after decoupling is also impossible.
XII Thermodynamic Quantities
To link an equation of state to Eqs. (62), one needs gauge-invariant expressions for the entropy, temperature, and pressure perturbations. These quantities must be expressed in terms of the relative density perturbations $\delta_{\varepsilon}$ and $\delta_{n}$. This section presents the derivation of these expressions.
XII.1 Entropy Perturbations
The combined first and second laws of thermodynamics for a simple, single-species system is given by (see, for example, Ref. [49], Sec. 2.1)
${\mathrm{d}}E=T{\mathrm{d}}S-p\,{\mathrm{d}}V+\mu\,{\mathrm{d}}N,$ (66)
where $E$, $S$, and $N$ are the energy, the entropy and the number of particles of a system with volume $V$, temperature $T$, and pressure $p$. The thermal — or chemical — potential $\mu$, is the energy required to add one particle to the system. In terms of the particle number density $n=N/V$, the energy per particle $E/N=\varepsilon/n$ and the entropy per particle $s=S/N$ the law (66) can be rewritten in the form
${\mathrm{d}}\left(\dfrac{\varepsilon}{n}N\right)=T{\mathrm{d}}(sN)-p\,{\mathrm{d}}\left(\dfrac{N}{n}\right)+\mu\,{\mathrm{d}}N,$ (67)
where $\varepsilon$ is the energy density. The system is extensive, i.e., $S(\lambda E,\lambda V,\lambda N)=\lambda S(E,V,N)$, which implies that the entropy of the gas is given by $S=(E+pV-\mu N)/T$. Upon dividing this relation by $N$ the Euler relation is obtained:
$\mu=\dfrac{\varepsilon+p}{n}-Ts.$ (68)
Eliminating $\mu$ in (67) with the aid of (68) reveals that the combined first and second laws of thermodynamics (66) can be expressed in a form that does not include $\mu$ and $N$ (see Ref. [32]):
$T{\mathrm{d}}s={\mathrm{d}}\Bigl(\dfrac{\varepsilon}{n}\Bigr)+p\,{\mathrm{d}}\Bigl(\dfrac{1}{n}\Bigr).$ (69)
From the background equations (11) and the thermodynamic law (69) it can be demonstrated that $\dot{s}_{\scriptscriptstyle(0)}=0$}, indicating that the expansion of the universe takes place without generating entropy. From the definition of gauge-invariant quantities (7), it can be concluded that $s_{\scriptscriptstyle(1)}=s_{\scriptscriptstyle(1)}^{\textnormal{phys}}$ is automatically gauge-invariant. By using (136) and $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$, the thermodynamic relation (69) can be rewritten as
[…]
By using the definitions (61), the following is obtained:
$T_{\scriptscriptstyle(0)}s^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{n_{\scriptscriptstyle(0)}}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (71)
Therefore, Eq. (62b) can be regarded as the evolution equation for the entropy per particle.
XII.2 Temperature and Pressure Perturbations
Since the background temperature $T_{{\scriptscriptstyle(0)}}$ and pressure $p_{{\scriptscriptstyle(0)}}$ are scalars, their gauge-invariant perturbations are defined using the definition (7):
$T^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=T_{\scriptscriptstyle(1)}-\dfrac{\dot{T}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)},\quad p^{\textnormal{phys}}_{\scriptscriptstyle(1)}:=p_{\scriptscriptstyle(1)}-\dfrac{\dot{p}_{\scriptscriptstyle(0)}}{\dot{\theta}_{\scriptscriptstyle(0)}}\theta_{\scriptscriptstyle(1)}.$ (72)
The expression for $T^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ will be used in Sec. XIII.2. In order to arrive at the gauge-invariant counterpart of (14a), it is first necessary to eliminate $\varepsilon_{{\scriptscriptstyle(1)}}$ and $n_{{\scriptscriptstyle(1)}}$ from (14a) using (6a). Using $\dot{p}_{{\scriptscriptstyle(0)}}=p_{n}\dot{n}_{{\scriptscriptstyle(0)}}+p_{\varepsilon}\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and (72), yields
$p^{\textnormal{phys}}_{\scriptscriptstyle(1)}=p_{n}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{\varepsilon}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (73)
Eliminating $p_{\varepsilon}$ with the aid of (40) and using the definitions provided in (61) for relative density perturbations results in the following expression:
$p_{\scriptscriptstyle(1)}^{\textnormal{phys}}=\beta^{2}\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}+n_{\scriptscriptstyle(0)}p_{n}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (74)
The first term represents the adiabatic component of the pressure perturbation, while the second term denotes the nonadiabatic component.
XIII Application: Flat FLRW Universe
The evolution equations presented in (62), are applicable to open, flat, and closed flrw universes. We will now focus on studying density perturbations in a flat flrw universe, for which ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$.
Given that the initial density perturbations occurred in the early universe when $\Lambda\ll\kappa\varepsilon_{\scriptscriptstyle(0)}$, the cosmological constant $\Lambda$ will be neglected. Consequently, the background equations (11), are reduced to
$3H^{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)},\quad$ $\kappa$ $=8\pi G_{\text{N}}/c^{4},$ (75a)
$\dot{\varepsilon}_{\scriptscriptstyle(0)}$ $=-3H\varepsilon_{\scriptscriptstyle(0)}(1+w),$ $\quad w$ $:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)},$ (75b)
$\dot{n}_{\scriptscriptstyle(0)}$ $=-3Hn_{\scriptscriptstyle(0)}.$ (75c)
By using the Friedmann equation (75a), the coefficients (63) of Eq. (62a) result in
$b_{1}$ $=H(1-3w-3\beta^{2})-2\dfrac{\dot{\beta}}{\beta},$ (76a)
$b_{2}$ $=\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl(2\beta^{2}(2+3w)-\tfrac{1}{6}(1+18w+9w^{2})\Bigr)$
$+2H\dfrac{\dot{\beta}}{\beta}(1+3w)-\beta^{2}\dfrac{\nabla^{2}}{a^{2}},$ (76b)
$b_{3}$ $=\Biggl[\dfrac{-2}{1+w}\Biggl(\varepsilon_{\scriptscriptstyle(0)}p_{\varepsilon n}(1+w)+\dfrac{2p_{n}}{3H}\dfrac{\dot{\beta}}{\beta}$
$+p_{n}(p_{\varepsilon}-\beta^{2})+n_{\scriptscriptstyle(0)}p_{nn}\Biggr)+p_{n}\Biggr]\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}\dfrac{\nabla^{2}}{a^{2}},$ (76c)
where $\nabla^{2}$ represents the conventional Laplace operator.
XIII.1 Era before Decoupling of Matter and Radiation
In this era the primordial fluid is a mixture of radiation and matter, wherein the contribution of matter to the pressure is negligible. Consequently, the equations of state are given by [see Ref. [13], § V-1]:
$\varepsilon=a_{\text{B}}T_{\gamma}^{4}+nmc^{2},\quad p=\tfrac{1}{3}a_{\text{B}}T_{\gamma}^{4}.$ (77)
The black body constant is represented by $a_{\text{B}}$, the radiation temperature is denoted by $T_{\gamma}$ and the particle number density of ordinary matter or cdm is given by $n$. Upon eliminating $T_{\gamma}$, the following result is obtained [see Ref. [42], Eq. (2.10.27)]:
$p=\tfrac{1}{3}(\varepsilon-nmc^{2}).$ (78)
By making use of (14b), one has
$p_{n}=-\tfrac{1}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{1}{3}.$ (79)
Since $p_{n}<0$, energy density perturbations are coupled to particle number density perturbations according to (65).
For the parameter $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and the nonadiabatic speed of sound $\beta$, (40), it is found
$w=\frac{\tfrac{1}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}}{a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}},\quad\beta^{2}=\frac{1}{3}-\frac{\tfrac{1}{3}n_{{\scriptscriptstyle(0)}}mc^{2}}{\tfrac{4}{3}a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}+n_{{\scriptscriptstyle(0)}}mc^{2}}.$ (80)
The universe was radiation-dominated when
$a_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}^{4}\gg n_{{\scriptscriptstyle(0)}}mc^{2}.$ (81)
In this case one has $\beta^{2}\approx\tfrac{1}{3}$, and $w\approx\tfrac{1}{3}$, implying that $\dot{\beta}\approx 0$. Upon substituting these values and expressions (79) into the coefficients (76), and using (81) the result is
$\ddot{\delta}_{\varepsilon}-H\dot{\delta}_{\varepsilon}-\left(\dfrac{1}{3}\dfrac{\nabla^{2}}{a^{2}}-\tfrac{2}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}=0.$ (82)
In order to solve Eq. (82), it is necessary to use the solutions to the background equations (75). These solutions are given by
$H=\tfrac{1}{2}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}(ct)^{-2},\quad a\propto t^{1/2}.$ (83)
The dimensionless time, denoted by the symbol $\tau$, is defined by $\tau:=t/t_{0}\geq 1$. This definition implies that
$\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{0}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl(2H(t_{0})\bigr)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (84)
By using the Helmholtz equation $\nabla^{2}\delta_{\varepsilon}=-|\bm{q}|^{2}\delta_{\varepsilon}$ and the result derived in (84), it can be shown that the equation (82) can be rewritten as
$\delta_{\varepsilon}^{\prime\prime}-\dfrac{1}{2\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{\mu_{\text{r}}^{2}}{4\tau}+\dfrac{1}{2\tau^{2}}\right)\delta_{\varepsilon}=0,$ (85)
where the prime denotes differentiation with respect to the $\tau$. With $|\bm{q}|=2\pi/\lambda_{0}$, the parameter $\mu_{\text{r}}$ is defined by
$\mu_{\text{r}}:=\dfrac{2\pi}{\lambda_{0}}\dfrac{1}{H(t_{0})}\dfrac{1}{\sqrt{3}}.$ (86)
In this expression, $\lambda_{0}:=\lambda a(t_{0})$ represents the physical scale of a density perturbation at $t=t_{0}$.
In order to solve Eq. (85), it is first necessary to replace the variable $\tau$ with the variable defined by $x:=\mu_{\text{r}}\sqrt{\tau}$. Upon transforming back to $\tau$, the general solution of Eq. (85) with constants of integration $A_{1}(\bm{q})$ and $A_{2}(\bm{q})$ is obtained:
$\delta_{\varepsilon}(\tau,\bm{q})=\bigl[A_{1}(\bm{q})\sin\left(\mu_{\text{r}}\sqrt{\tau}\right)+A_{2}(\bm{q})\cos\left(\mu_{\text{r}}\sqrt{\tau}\right)\bigr]\sqrt{\tau}.$ (87)
Since $\vartheta_{{\scriptscriptstyle(1)}}$ is included in Eq. (82) it yields oscillating relative density perturbations with an increasing amplitude. In contrast, the standard equation (153), for which $\vartheta_{{\scriptscriptstyle(1)}}=0$ (i.e., no fluid flow) yields oscillating solutions with a constant amplitude.
In the case of large-scale perturbations, that is, $\lambda_{0}\rightarrow\infty$}, which leads to $\mu_{\text{r}}\rightarrow 0$, (86), the general solution to Eq. (85) is given by
$\delta_{\varepsilon}(\tau)=-\bigl[\delta_{\varepsilon}(1)-2\delta^{\prime}_{\varepsilon}(1)\bigr]\tau+2\bigl[\delta_{\varepsilon}(1)-\delta^{\prime}_{\varepsilon}(1)\bigr]\tau^{1/2}.$ (88)
A comparison of the solutions (88) and (155a) reveals that the solution proportional to $\tau$ in (88) is also a solution to the homogeneous part of equation (154a). The solution proportional to $\tau^{1/2}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.1. The minus sign preceding the solution proportional to $\tau$ is a consequence of the fact that $\vartheta_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}>0$}, as follows from (155b). The presence of the gauge function $\psi(\bm{x})$ precludes the derivation of the solution (88) from (154).
XIII.2 Era after Decoupling of Matter and Radiation
Once protons and electrons have combined to form hydrogen, radiation pressure will be negligible and the equations of state will be those of a nonrelativistic, monatomic, perfect gas with three degrees of freedom. Therefore, pressure can be neglected in the background universe. However, the momentum conservation law (49) indicates that, in order to study the evolution of density perturbations, one must consider pressure perturbations, since density perturbations are static (51) in a pressureless fluid. According to the literature [see Ref. [32], Eq. (13), and Ref. [42], Eqs. (15.8.20)–(15.8.21)], we use the following equations of state for a nonrelativistic gas:
$\varepsilon(n,T)=nmc^{2}+\tfrac{3}{2}nk_{\text{B}}T,\quad p(n,T)=nk_{\text{B}}T.$ (89)
In these expressions, the symbol $k_{\text{B}}$ represents Boltzmann’s constant, the quantity $m$ denotes the proton rest mass, and the temperature of the matter is indicated by the symbol $T$. The rest mass energy density is $nmc^{2}$ and the kinetic energy density is $\tfrac{3}{2}nk_{\text{B}}T$.
The elimination of the temperature $T$ from the equations of state (89), yields the equation of state for the pressure [see Ref. [42], Eq. (2.10.27)]:
$p(n,\varepsilon)=\tfrac{2}{3}(\varepsilon-nmc^{2}).$ (90)
The partial derivatives are determined using the expressions (14b) and are given by
$p_{n}=-\tfrac{2}{3}mc^{2},\quad p_{\varepsilon}=\tfrac{2}{3}.$ (91)
The parameter $w$ is defined by
$w:=\dfrac{p_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}=\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}+\tfrac{3}{2}k_{\text{B}}T_{\scriptscriptstyle(0)}}\approx\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}\ll 1.$ (92)
Upon substituting (89), (91) and (92) into (40), one arrives at the well-known result [see Ref. [42], Eq. (15.8.22)]:
$\quad\beta\approx\dfrac{v_{\text{s}}}{c}=\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{\scriptscriptstyle(0)}}{mc^{2}}},$ (93)
where $v_{\text{s}}$ represents the adiabatic speed of sound.
From $\beta^{2}\approx\tfrac{5}{3}w$, we find from Eq. (44) that $\dot{w}\approx-2Hw$. Therefore, with $H:=\dot{a}/a$, it can be concluded that $w\propto a^{-2}$. Given that $w$ is proportional to $T_{{\scriptscriptstyle(0)}}$, the well-known result (see Ref. [42], Eq. (15.5.16) with $\gamma=\tfrac{5}{3}$, where $\gamma:=c_{\text{p}}/c_{\text{v}}$} is the adiabatic index) is obtained:
$T_{\scriptscriptstyle(0)}\propto a^{-2}.$ (94)
The subsequent step is to derive the evolution equations for relative density perturbations. The proportionality (94) implies with (93) that $\dot{\beta}/\beta=-H$. The system of equations (62) with coefficients (76) can now be rewritten in the following form:
$\ddot{\delta}_{\varepsilon}+3H\dot{\delta}_{\varepsilon}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{5}{6}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta_{\varepsilon}=-\dfrac{2}{3}\dfrac{\nabla^{2}}{a^{2}}\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95a)
$\dfrac{1}{c}\dfrac{\mathrm{d}}{{\mathrm{d}}t}\left(\delta_{n}-\delta_{\varepsilon}\right)=-2H\left(\delta_{n}-\delta_{\varepsilon}\right),$ (95b)
where it is used that $w$ and $\beta^{2}$ are negligible with respect to constants of order one.
Given that the term $(\delta_{n}-\delta_{\varepsilon})$ occurs within the source term of Eq. (95a), it is necessary to first solve for the solution to Eq. (95b). By making the substitution $H:=\dot{a}/a$, the following result is obtained:
$(\delta_{n}-\delta_{\varepsilon})\propto a^{-2}.$ (96)
In order to relate this proportionality to thermodynamic quantities, it is necessary to express the perturbed equation of state for the energy density in terms of relative density perturbations. It can be deduced from $\varepsilon=\varepsilon(n,T)$ that
$\dot{\varepsilon}_{\scriptscriptstyle(0)}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}\dot{n}_{\scriptscriptstyle(0)}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}\dot{T}_{\scriptscriptstyle(0)},$ (97a)
$\varepsilon_{\scriptscriptstyle(1)}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}.$ (97b)
By multiplying (97a) by $\theta_{\scriptscriptstyle(1)}/\dot{\theta}_{\scriptscriptstyle(0)}$ and subtracting the result from (97b), one obtains the following result:
$\varepsilon_{{\scriptscriptstyle(1)}}^{{\textnormal{phys}}}=\left(\dfrac{\partial\varepsilon}{\partial n}\right)_{T}n_{\scriptscriptstyle(1)}^{{\textnormal{phys}}}+\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{n}T_{\scriptscriptstyle(1)}^{{\textnormal{phys}}},$ (98)
where expressions (6a) and (72) have been used. By using the expression for $\varepsilon$ in (89) to eliminate the partial derivatives, one obtains the perturbed equation of state:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}mc^{2}+\tfrac{3}{2}n^{\textnormal{phys}}_{\scriptscriptstyle(1)}k_{\text{B}}T_{\scriptscriptstyle(0)}+\tfrac{3}{2}n_{\scriptscriptstyle(0)}k_{\text{B}}T^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (99)
Upon dividing (99) by $\varepsilon_{{\scriptscriptstyle(0)}}=n_{{\scriptscriptstyle(0)}}mc^{2}+\tfrac{3}{2}n_{{\scriptscriptstyle(0)}}k_{\text{B}}T_{{\scriptscriptstyle(0)}}$ and using the exact value of the ratio $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ (92), the perturbed equation of state for the energy density expressed in relative density perturbations (61) is obtained:
$\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})=-\tfrac{3}{2}w(t)\delta_{T}(t,\bm{x}).$ (100)
The quantity $\delta_{T}$ is defined by $\delta_{T}:=T^{\textnormal{phys}}_{\scriptscriptstyle(1)}/T_{\scriptscriptstyle(0)}$. The solution to Eq. (95b) is obtained by combining (96) and (100). Using the approximate value of $w$ implies that $w\propto a^{-2}$ [see (94)], making $\delta_{T}$ independent of time. Therefore, the solution to Eq. (95b) is given by
$\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x}).$ (101)
The quantity $\delta_{T}(t_{\text{dec}},\bm{x})$ represents the relative temperature perturbation of matter at the time $t_{\text{dec}}$, which marks the decoupling of matter from radiation.
According to (71), (92) and (101) the entropy per particle is:
$s^{\textnormal{phys}}_{\scriptscriptstyle(1)}(t,\bm{x})\approx\tfrac{3}{2}k_{\text{B}}\delta_{T}(t_{\text{dec}},\bm{x}).$ (102)
Thus, the quantity $\delta_{T}$ is random. As a result, it is generally not equal to zero, which implies that density perturbations are generally nonadiabatic.
Using (91), (93), and (101), the local relative pressure perturbation $\delta_{p}:=p^{\textnormal{phys}}_{\scriptscriptstyle(1)}/p_{\scriptscriptstyle(0)}$ can be calculated from expression (74). One obtains
$\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x}).$ (103)
In this context, the terms $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ and $\delta_{T}(t_{\text{dec}},\bm{x})$ represent the adiabatic and the random, nonadiabatic relative pressure perturbation, respectively.
The solution (101) to Eq. (95b) allows for the reformulation of the second-order equation (95a). To that end, the solutions to the background equations (75) are required. Given that the pressure with respect to the rest mass energy density can be disregarded ($w\ll 1$), the following solutions are obtained:
$H=\tfrac{2}{3}(ct)^{-1},\quad\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}(ct)^{-2},\quad a\propto t^{2/3}.$ (104)
The dimensionless time, denoted by the symbol $\tau$, is defined by $\tau:=t/t_{\text{dec}}\geq 1$. This definition implies that
$\dfrac{{\mathrm{d}}^{k}}{c^{k}{\mathrm{d}}t^{k}}=\left(\dfrac{1}{ct_{\text{dec}}}\right)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}=\bigl(\tfrac{3}{2}H(t_{\text{dec}})\bigr)^{k}\dfrac{{\mathrm{d}}^{k}}{{\mathrm{d}}\tau^{k}}.$ (105)
By using Eqs. (75a) and (101), as well as the expressions (93), (104), (105), the Helmholtz equation $\nabla^{2}\delta=-|\bm{q}|^{2}\delta$, and $w\propto a^{-2}$, it can be demonstrated that Eq. (95a) can be expressed in the following manner:
$\delta_{\varepsilon}^{\prime\prime}+\dfrac{2}{\tau}\delta_{\varepsilon}^{\prime}+\left(\dfrac{4}{9}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}-\dfrac{10}{9\tau^{2}}\right)\delta_{\varepsilon}=-\dfrac{4}{15}\dfrac{\mu^{2}_{\text{m}}}{\tau^{8/3}}\delta_{T}(t_{\text{dec}},\bm{q}),$ (106)
where the prime denotes differentiation with respect to $\tau$. With $\lvert\bm{q}\rvert=2\pi/\lambda_{\text{dec}}$, the parameter $\mu_{\text{m}}$ is defined by
$\mu_{\text{m}}:=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{H(t_{\text{dec}})}\dfrac{v_{\text{s}}(t_{\text{dec}})}{c}.$ (107)
In this expression, $\lambda_{\text{dec}}:=\lambda a(t_{\text{dec}})$ represents the physical scale of a perturbation at $t=t_{\text{dec}}$, and $v_{\text{s}}(t_{\text{dec}})/c\approx\beta(t_{\text{dec}})$ is given by (93).
Finally, the expression (103) enables the reformulation of Eq. (106) in a form suitable for the study of the evolution of relative density perturbations:
$\delta^{\prime\prime}_{\varepsilon}+\frac{2}{\tau}\delta^{\prime}_{\varepsilon}+\dfrac{4}{15}\dfrac{\mu_{\text{m}}^{2}}{\tau^{8/3}}\delta_{p}-\dfrac{10}{9\tau^{2}}\delta_{\varepsilon}=0,\quad\tau:=\dfrac{t}{t_{\text{dec}}}\geq 1.$ (108)
The second term in this equation represents the expansion, the third term is the pressure term, where $\delta_{p}$ is given by (103), and the fourth term represents gravitation.
For large-scale perturbations, i.e., when $\lambda_{\text{dec}}\rightarrow\infty$, implying that $\mu_{\text{m}}\rightarrow 0$, relative pressure perturbations $\delta_{p}$ played a negligible role in their evolution. Therefore, the solution to Eq. (108) is
$\delta_{\varepsilon}(\tau)=\bigl[\tfrac{5}{7}\delta_{\varepsilon}(1)+\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr]\tau^{2/3}+\bigl[\tfrac{2}{7}\delta_{\varepsilon}(1)-\tfrac{3}{7}\delta^{\prime}_{\varepsilon}(1)\bigr]\tau^{-5/3}.$ (109)
The solution proportional to $\tau^{-5/3}$ dies out quickly. Thus, large-scale relative density perturbations evolve nearly proportionally to $\tau^{2/3}$, which is close to the solution (60) for a pressureless fluid. However, since $\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(\tau)=-\tfrac{7}{9}\tau^{-4/3}$, (158b), there is a pressure-gradient-driven fluid flow. Therefore, the solution (109) is invalid for a pressureless fluid. As shown in Sec. X, the equations for scalar perturbations (47) are, for a pressureless fluid, reduced to a single equation, namely the time-independent Poisson equation (56).
A comparison of the solutions (109) and (158a) reveals that the solution proportional to $\tau^{2/3}$ in (109) is also a solution to the homogeneous part of Eq. (157a). The solution proportional to $\tau^{-5/3}$ is the particular solution to the inhomogeneous equation. This solution can only be obtained if the covariant divergence $\vartheta_{{\scriptscriptstyle(1)}}$ is taken into account, as demonstrated in Appendix C.2. Since in the standard equation (156) $\vartheta_{{\scriptscriptstyle(1)}}=0$ (i.e., no fluid flow), it does not produce growing density perturbations.
The presence of the gauge constant $C$, precludes the derivation of solution (109) from the system (157).
XIII.2.1 Cosmological Quantities
Table 1: Planck satellite resultsy
$z(t_{\text{dec}})$ = 1090
$z(t_{\text{eq}})$ = 3387
$cH(t_{\text{p}})$ = $67.66\,\text{km}\,\text{s}^{-1}\text{Mpc}^{-1}$
$T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})$ = $2.725\,\text{K}$
$t_{\text{p}}$ = $13.79\,\text{Gyr}$
$\lvert\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})\rvert \lesssim$ $10^{-5}$
To investigate the evolution of relative density perturbations using Eq. (108), the parameter $\mu_{\text{m}}$ (107) must first be expressed in terms of observable quantities, namely the redshift $z(t_{\text{dec}})$ at the time of decoupling, the present value of the Hubble parameter $cH(t_{\text{p}})$, and the present value of the temperature $T_{{\scriptscriptstyle(0)}}(t_{\text{p}})$ of the cosmological fluid.
The redshift $z(t)$ is defined by the expression
$z(t):=\dfrac{a(t_{\text{p}})}{a(t)}-1,$ (110)
where $a(t_{\text{p}})$ is the present value of the scale factor. The Friedmann equation (11a) shows that, for a flat flrw universe, one may set $a(t_{\text{p}})=1$. Using the background solutions (104), it is possible to express the relevant parameters of the universe in the redshift:
$H(t)$ $=H(t_{\text{p}})\bigl(z(t)+1\bigr)^{3/2},$ (111a)
$t$ $=t_{\text{p}}\bigl(z(t)+1\bigr)^{-3/2},$ (111b)
$T_{{\scriptscriptstyle(0)}\gamma}(t)$ $=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})\bigl(z(t)+1\bigr),$ (111c)
$n_{{\scriptscriptstyle(0)}}(t)$ $=n_{{\scriptscriptstyle(0)}}(t_{\text{p}})\bigl(z(t)+1\bigr)^{3}.$ (111d)
In deriving relation (111c), it is used that, subsequent to decoupling, $T_{{\scriptscriptstyle(0)}\gamma}\propto a^{-1}$. Upon substituting (93) into expression (107) and using the relations (111), we obtain
$\mu_{\text{m}}=\dfrac{2\pi}{\lambda_{\text{dec}}}\dfrac{1}{cH(t_{\text{p}})}\dfrac{1}{z(t_{\text{dec}})+1}\sqrt{\dfrac{5}{3}\dfrac{k_{\text{B}}T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{p}})}{m}},$ (112)
where it is assumed that at decoupling, the matter and radiation temperatures were equal, that is, $T_{\scriptscriptstyle(0)}(t_{\text{dec}})=T_{{\scriptscriptstyle(0)}\gamma}(t_{\text{dec}})$. In this manner, the parameter $\mu_{\text{m}}$ has been expressed in observable quantities. Upon substituting the numerical values from Tables 1 and 2, one obtains
$\mu_{\text{m}}=\dfrac{16.48}{\lambda_{\text{dec}}},\quad\lambda_{\text{dec}}\text{ in pc}.$ (113)
It can thus be concluded that the parameter $\mu_{\text{m}}$ is dependent solely on the initial scale $\lambda_{\text{dec}}$ of a density perturbation.
Finally, the dimensionless time, defined as $\tau:=t/t_{\text{dec}}$, is expressed in terms of the redshift, given by:
$\tau=\left(\dfrac{z(t_{\text{dec}})+1}{z(t)+1}\right)^{3/2},$ (114)
where relation (111b) with $z(t_{\text{p}})=0$ is used.
Table 2: Physical constants
$m$ = $1.6726\times 10^{-27}\,\text{kg}$
pc = $3.0857\times 10^{16}\,\text{m}=3.2616\,\text{ly}$
$c$ = $2.9979\times 10^{8}\,\text{m}\,\text{s}^{-1}$
$k_{\text{B}}$ = $1.3806\times 10^{-23}\,\text{J}\,\text{K}^{-1}$
$G_{\text{N}}$ = $6.6743\times 10^{-11}\,\text{m}^{3}\,\text{kg}^{-1}\,\text{s}^{-2}$
$\text{M}_{\odot}$ = $1.9889\times 10^{30}\,\text{kg}$
$a_{\text{B}}$ = $7.5657\times 10^{-16}\,\text{J}\,\text{m}^{-3}\,\text{K}^{-4}$
XIII.2.2 Initial Values
In order to solve Eq. (108), it is necessary to determine the initial values of the quantities $\delta_{\varepsilon}$ and $\delta^{\prime}_{\varepsilon}$. In addition, values of the local random, nonadiabatic relative pressure perturbation $\delta_{T}$ in expression (103), are required.
Planck Satellite
The Planck observations [50] of the relative perturbations $\delta_{T_{\gamma}}(t_{\text{dec}},\bm{x})$ in the background radiation temperature imply that $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$. In the absence of knowledge regarding the initial growth rate, it is assumed that $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$. Therefore, the initial values for Eq. (108) are as follows:
$\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\lesssim 10^{-5},\quad\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0.$ (115)
As demonstrated in the calculations presented in Sec. XIII.2.3, the outcome of Eq. (108) is almost entirely independent of the initial value for $\delta_{\varepsilon}$, provided that it satisfies the condition $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\rvert\leq 10^{-4}$.
Random Nonadiabatic Pressure Perturbations
The initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ are related to the local random, nonadiabatic relative pressure perturbation $\delta_{T}(t_{\text{dec}},\bm{q})$ by expression (101). Since $w\ll 1$}, very small values of the difference $\delta_{n}(t_{\text{dec}},\bm{q})-\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ result in relatively large positive or negative values of $\delta_{T}(t_{\text{dec}},\bm{q})$.
The fact that $\delta_{T}(t_{\text{dec}},\bm{q})$ takes on random values follows from expression (102). However, using Eq. (101) the randomness of $\delta_{T}(t_{\text{dec}},\bm{q})$ can also be explained as follows. The transition of the universe from the era preceding decoupling to the era following decoupling was rapid and chaotic. A notable decline was observed in the mean particle velocity, and the pressure. As a consequence, the values of both $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})$ and $\delta_{n}(t_{\text{dec}},\bm{q})$ exhibited minor irregularities on the surface of last scattering. Therefore, it is probable that after the transition, the initial values of the difference between the relative density perturbations, represented by $(\delta_{n}-\delta_{\varepsilon})(t_{\text{dec}},\bm{q})$}, were randomly distributed among all density perturbations.
XIII.2.3 Structure Formation
The subsequent analysis will use Eq. (108) to examine the evolution of relative density perturbations. This equation shows that the evolution of a relative perturbation, $\delta_{\varepsilon}(t,\bm{q})$, with an initial scale $\lambda_{\text{dec}}$, was influenced by three factors: the expansion of the universe, pressure perturbations, and gravitation. Relative pressure perturbations $\delta_{p}$, as indicated by (103), consisted of two discrete components: an adiabatic contribution, represented by $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{q})$, and a random (102), nonadiabatic contribution, represented by $\delta_{T}(t_{\text{dec}},\bm{q})$, where the adiabatic component was initially negligible (115).
Figure 1: The graphs show the redshift and time when a relative perturbation in the energy density with initial scale $\lambda_{\text{dec}}$, and initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$ starting to grow at an initial redshift of $z(t_{\text{dec}})=1090$ has become nonlinear, i.e., when $\delta_{\varepsilon}(t,\bm{q})=1$. The graphs are labeled with the initial values of the random, nonadiabatic pressure perturbations $\delta_{T}(t_{\text{dec}},\bm{q})$. For each graph, the Jeans scale is $6.4\,\text{pc}$.
Since $w\ll 1$, (92), the nonadiabatic component could have taken on relatively large positive or negative values for small differences in the relative perturbations of energy and particle number density, as shown in (101). Consequently, there were regions where the total relative pressure perturbation was initially negative after decoupling. This allowed density perturbations to grow rapidly, despite the fact that the expansion of the universe opposed growth. As the adiabatic component of the relative pressure perturbation increased, the total relative pressure perturbation became positive. This significantly reduced the growth rate of density perturbations. The initial growth phase was brief yet long enough for a density perturbation to reach the nonlinear regime several hundred million years after the Big Bang. Conversely, if the nonadiabatic relative pressure perturbation was positive initially, then the total relative pressure perturbation was also positive. This resulted in the formation of voids, which were regions with negative values of $\delta_{\varepsilon}$. In these regions, matter was driven to the edges.
Fig. 1 provides a summary of the evolution of relative density perturbations. It was created using the following methodology. For each value of $\delta_{T}=-0.01$}, 0.05-0.05, 0.10-0.10, 0.15-0.15, and 0.20-0.20, Eq. (108) was solved numerically for a large number of values for the initial perturbation scale $\lambda_{\text{dec}}$ using the initial values (115). The integration started at $z=1090$ and ended when either $z=0$ or when $\delta_{\varepsilon}=1$. Each integration run returned a single point on the graph for a particular choice of the initial scale $\lambda_{\text{dec}}$ when $z>0$ and $\delta_{\varepsilon}=1$. In this case, it is clear that the perturbation has become nonlinear within $13.79\,\text{Gyr}$. On the other hand, if the integration stopped at $z=0$ and $\delta_{\varepsilon}<1$, then the perturbation had not yet reached its nonlinear phase today. Each graph shows the time and scale at which $\delta_{\varepsilon}=1$ for a given value of $\delta_{T}$.
As illustrated in Fig. 1, the optimal scale for growth was approximately $6.4\,\text{pc}$. Perturbations at this scale entered the nonlinear phase as early as $40$ million years ($z\approx 49$) after the Big Bang. Perturbations with scales smaller than $6.4\,\text{pc}$ reached their nonlinear phase at a much later time because their internal gravitation was weaker than for large-scale perturbations. Furthermore, pressure perturbations and expansion of the universe led to oscillatory behavior, as shown in Fig. 2. In contrast, perturbations with a scale greater than $6.4\,\text{pc}$ were less affected by pressure perturbations. However, because of their larger size, the expansion worked against their growth, so that, despite their stronger gravitation, they also reached the nonlinear phase at a later time. Perturbations larger than $70\,\text{pc}$ did not reach their nonlinear regime within $13.79\,\text{Gyr}$.
Perturbations that became nonlinear within $13.79\,\text{Gyr}$ are within the particle horizon at decoupling given by:
$d_{\text{H}}(t_{\text{dec}})=a(t_{\text{dec}})\int_{0}^{ct_{\text{dec}}}\!\!\dfrac{\mathrm{d}\tau}{a(\tau)}=3ct_{\text{dec}}\approx 3.5\times 10^{5}\,\text{pc},$ (116)
where Eq. (111b) and Tables 1 and 2 have been used.
Figure 2: The graphs show the growth rates $\delta^{\prime}_{\varepsilon}$, with initial values $\delta_{\varepsilon}(t_{\text{dec}},\bm{q})\approx 10^{-5}$ and $\delta^{\prime}_{\varepsilon}(t_{\text{dec}},\bm{q})=0$, as function of the redshift $z$, or time in million of years. The initial scales $\lambda_{\text{dec}}$ of the perturbations are measured in parsec. The evolution of relative density perturbations started at $z=1090$.
As Fig. 2 shows, the largest initial growth rates occurred for perturbations that were smaller than $6.4\,\text{pc}$. The smallest perturbations which had become nonlinear within $13.79\,\text{Gyr}$, had a scale of approximately $2\,\text{pc}$, as shown in Fig. 1. Perturbations with scales $2\,{\text{pc}}\lesssim\lambda_{\text{dec}}\lesssim 4.1\,\text{pc}$} oscillated toward the nonlinear phase within $13.79\,\text{Gyr}$. After approximately $14$ million years, the total pressure perturbation (103) had become positive and the growth rate had decreased, resulting in gravitation and, to a lesser extent, expansion becoming the primary drivers of the evolution of density perturbations. Consequently, the most turbulent phase of density perturbations had concluded, and a phase of gradual and consistent gravitational growth toward the nonlinear regime had commenced. As can be seen in Fig. 1, at a redshift of $z\approx 7$, or $600$ million years after the Big Bang, all perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$ had become nonlinear.
Assuming a spherical density perturbation with a diameter equal to $\lambda_{\text{dec}}$, the mass of the perturbation at decoupling is given by
$M(t_{\text{dec}})=\dfrac{4\pi}{3}\bigl(\tfrac{1}{2}\lambda_{\text{dec}}\bigr)^{3}n_{\scriptscriptstyle(0)}(t_{\text{dec}})m.$ (117)
The particle number density at the time of decoupling, $n_{\scriptscriptstyle(0)}(t_{\text{dec}})$, can be derived from the Friedmann equation (75a). For $w\ll 1$ and at time $t_{\text{p}}$ this equation is given by:
$3H^{2}(t_{\text{p}})=\kappa n_{{\scriptscriptstyle(0)}}(t_{\text{p}})mc^{2},\quad\kappa=8\pi G_{\text{N}}/c^{4}.$ (118)
By using the relation (111d), one obtains
$n_{\scriptscriptstyle(0)}(t_{\text{dec}})=\frac{3\bigl[cH(t_{\text{p}})\bigr]^{2}\bigl[z(t_{\text{dec}})+1\bigr]^{3}}{8\pi mG_{\text{N}}}.$ (119)
The relations (117) and (119) and the constants in Tables 1 and 2 reveal that perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$, had masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$.
Because of the steepness of the graphs in Fig. 1 for scales below $6.4\,\text{pc}$, this scale is designated as the relativistic Jeans scale. Its value is
$\lambda_{\text{Jeans}}(t_{\text{dec}}):=\lambda_{\text{Jeans}}a(t_{\text{dec}})\approx 6.4\,\text{pc}.$ (120)
Accordingly, the Jeans mass at decoupling is given by
$M_{\text{Jeans}}(t_{\text{dec}})\approx 2.2\times 10^{4}\,\text{M}_{\odot}.$ (121)
The above analysis shows that, if nonadiabatic pressure perturbations are taken into account, relative density perturbations with initial values $\lvert\delta_{\varepsilon}(t_{\text{dec}},\bm{x})\rvert\lesssim 10^{-5}$ and $\dot{\delta}_{\varepsilon}(t_{\text{dec}},\bm{x})\approx 0$} can give rise to structure in the early universe after the decoupling of matter and radiation.
XIV Summary and Conclusion
This article introduces a relativistic cosmological perturbation theory (62) that describes the evolution of relative energy density and particle number density perturbations in closed ($K=+1$), flat ($K=0$), and open (K=1{K=-1}) Friedmann-Lemaître-Robertson-Walker (flrw) universes, where ${{}^{3}\!}R_{\scriptscriptstyle(0)}=-6K/a^{2}$ is the spatial curvature. It consists of two differential equations: a second-order inhomogeneous evolution equation for the relative energy density perturbation and a first-order entropy equation that determines the evolution of the source term of the first equation. These equations together describe the evolution of nonadiabatic relative perturbations $\delta_{\varepsilon}(t,\bm{x}):=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})/\varepsilon_{{\scriptscriptstyle(0)}}(t)$ and $\delta_{n}(t,\bm{x}):=n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}(t,\bm{x})/n_{{\scriptscriptstyle(0)}}(t)$ in energy and particle number densities, respectively.
The entropy equation implies that before matter and radiation decoupled, both ordinary matter perturbations and Cold Dark Matter (cdm) perturbations are coupled to energy density perturbations by gravitation. Consequently, cdm did not initiate structure formation after decoupling.
After decoupling, pressure gradients became essential for fluid flow, which is crucial for forming structures, since in a pressureless fluid, density perturbations $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ are static.
Newtonian gravity is not a suitable framework for studying the evolution of cosmological density perturbations. This is because the conventional evolution equation, whether derived from general relativity or Newtonian gravity, yields gauge-dependent solutions. Consequently, no distinction exists between sub- and super-horizon density perturbations.
XIV.1 Results for a Flat FLRW Universe
In a radiation-dominated universe, relative density perturbations $\delta_{\varepsilon}$ oscillated with an amplitude that increased proportionally to the square root of time.
After matter and radiation decoupled at $t=t_{\text{dec}}$, the cosmological fluid became a nonrelativistic perfect fluid. According to observations made by the Planck satellite [50], relative density perturbations were of the order of $10^{-5}$ at the time of decoupling. These perturbations are far too small to explain structures in the universe by gravitational growth, if only adiabatic perturbations are considered. This is because the growth of an adiabatic density perturbation slows down due to the increasing pressure perturbation caused by the density perturbation itself. For this reason, this article examines nonadiabatic density perturbations. This was achieved using an equation of state for a nonrelativistic perfect gas that considers not only the particles’ rest energy, but also their kinetic energy. Therefore, in addition to the energy density, the particle number density must also be considered. Then, the total pressure is given by $\delta_{p}(t,\bm{x})\approx\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})+\delta_{T}(t_{\text{dec}},\bm{x})$, where $\tfrac{5}{3}\delta_{\varepsilon}(t,\bm{x})$ is the adiabatic component caused by the density perturbation itself, and a random nonadiabatic component $\delta_{T}(t_{\text{dec}})$. In addition, the solution to the entropy equation is identical to the perturbed equation of state of the energy density, expressed in terms of relative density perturbations and the nonadiabatic component of pressure, that is, $\delta_{n}(t,\bm{x})-\delta_{\varepsilon}(t,\bm{x})\approx-\tfrac{3}{2}w(t)\delta_{T}(t_{\text{dec}},\bm{x})$, where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}\ll 1$. This expression shows that small random fluctuations in both $\delta_{n}(t_{\text{dec}},\bm{x})$ and $\delta_{\varepsilon}(t_{\text{dec}},\bm{x})$ — arising from the rapid and chaotic transition to the post-decoupling era — could lead to significant local random fluctuations in the relative nonadiabatic pressure perturbation, $\delta_{T}(t_{\text{dec}},\bm{x})$, which could take on both positive and negative values. If the total relative pressure perturbation $\delta_{p}(t_{\text{dec}},\bm{x})$ was positive at the moment of decoupling, a void was formed and matter moved toward its edges. Conversely, if $\delta_{p}(t_{\text{dec}},\bm{x})$ was negative, the density perturbation would undergo a brief period of rapid growth. This rapid growth continued until the total relative pressure perturbation $\delta_{p}(t,\bm{x})$ became positive. After that, the density perturbation slowly but steadily grew toward the nonlinear phase, which was reached early on. Perturbations smaller than $70\,\text{pc}$ would reach their nonlinear phase within $13.79$ billion years or less. The relativistic Jeans scale was $6.4\,\text{pc}$, and the corresponding Jeans mass was $2.2\times 10^{4}\,\text{M}_{\odot}$. Perturbations of Jeans scale size reached their nonlinear regime at a redshift of $z\approx 49$ or approximately $40$ million years after the Big Bang. All density perturbations with scales between $2\,\text{pc}$ and $24\,\text{pc}$ reached their nonlinear phase within $z\approx 7$ or $600$ million years. These scales correspond to masses between $6.7\times 10^{2}\,\text{M}_{\odot}$ and $1.2\times 10^{6}\,\text{M}_{\odot}$. This is smaller than what was observed by the James Webb Space Telescope (jwst) [51], which observed mature galaxies. Nevertheless, it is reasonable to assume that the masses continued to increase in the nonlinear phase due to their greater gravitational pull.
XIV.2 Conclusion
Investigating cosmological density perturbations in a flat flrw universe reveals that if a nonrelativistic, monatomic, perfect gas with three degrees of freedom accurately represents the cosmological fluid in our universe after matter and radiation decoupled, then a perturbation theory based on general relativity can explain the origin of the first structures in the universe without relying on cdm. Subsequent observations with the jwst, launched on Christmas Day $2021$, are expected to yield novel insights into the conditions that ensued immediately after the decoupling of radiation and matter, as well as the formation of the earliest structures in the universe. These observations will help us determine if general relativity accurately describes the formation of the first structures in the universe.
Appendix A Derivation of the
Perturbation Theory
Table 3: The coefficients $\alpha_{ij}$ figuring in the equations (124)
$3H(1+p_{\varepsilon})+\dfrac{\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}{2H}$ $3Hp_{n}$ $\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $\dfrac{\varepsilon_{\scriptscriptstyle(0)}(1+w)}{4H}$
$\dfrac{\kappa n_{\scriptscriptstyle(0)}}{2H}$ $3H$ $n_{\scriptscriptstyle(0)}$ $\dfrac{n_{\scriptscriptstyle(0)}}{4H}$
$\dfrac{p_{\varepsilon}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $\dfrac{p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}}{a^{2}}$ $H(2-3\beta^{2})$ 0
$\dfrac{\kappa\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}$ 0 $2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)$ $2H+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{6H}$
$\dfrac{-\,{{}^{3}\!}R_{\scriptscriptstyle(0)}}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ 0 $\dfrac{6\varepsilon_{\scriptscriptstyle(0)}H(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$ $\dfrac{\tfrac{3}{2}\varepsilon_{\scriptscriptstyle(0)}(1+w)}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}$
In this appendix, the perturbation theory (62) is derived. Before undertaking this task, it is necessary to carry out two preliminary steps.
Firstly, as $\theta^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}=0$, (6b), it is unnecessary to consider the gauge-dependent quantity $\theta_{{\scriptscriptstyle(1)}}$. Eliminating the quantity $\theta_{{\scriptscriptstyle(1)}}$ from the differential equations (37b)–(37e) using the algebraic constraint equation (37a) yields a system of four first-order ordinary differential equations suitable for computer algebra software:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}$ $+3H(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)})$
$+\varepsilon_{\scriptscriptstyle(0)}(1+w)\Bigl[\vartheta_{\scriptscriptstyle(1)}+\frac{1}{2H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)\Bigr]=0,$ (122a)
$\dot{n}_{\scriptscriptstyle(1)}$ $+3Hn_{\scriptscriptstyle(1)}+n_{\scriptscriptstyle(0)}\Bigl[\vartheta_{\scriptscriptstyle(1)}+\frac{1}{2H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)\Bigr]=0,$ (122b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}$ $+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\frac{1}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\dfrac{\tilde{\nabla}^{2}p_{\scriptscriptstyle(1)}}{a^{2}}=0,$ (122c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$ $+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$
$-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}+\frac{{{}^{3}\!}R_{\scriptscriptstyle(0)}}{3H}\left(\kappa\varepsilon_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}\right)=0.$ (122d)
Secondly, the energy density and particle number density perturbations (6a) are expressed in the four quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$. Using the background equations (11) to eliminate all time derivatives, and the linearized constraint equation (37a) to eliminate $\theta_{\scriptscriptstyle(1)}$, the following results are obtained:
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\dfrac{\varepsilon_{\scriptscriptstyle(1)}{{}^{3}\!}R_{\scriptscriptstyle(0)}-3\varepsilon_{\scriptscriptstyle(0)}(1+w)(2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)},$ (123a)
$n^{\textnormal{phys}}_{\scriptscriptstyle(1)}=n_{\scriptscriptstyle(1)}-\dfrac{3n_{\scriptscriptstyle(0)}(\kappa\varepsilon_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)})}{{{}^{3}\!}R_{\scriptscriptstyle(0)}+3\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (123b)
From an algebraic standpoint, it is more straightforward to begin by deriving equations for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. This will be accomplished in Sec. A.1. Subsequently, Eqs. (62) for the relative perturbations (61) will be derived in Sec. A.2. The expression (44) will be used to eliminate the time derivative of $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$.
A.1 Evolution Equation for Energy Density Perturbations
The system (122) and the expression (123a) are reformulated using (14a) in a form suitable for implementation in a computer algebra program:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+\alpha_{11}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{12}n_{\scriptscriptstyle(1)}+\alpha_{13}\vartheta_{\scriptscriptstyle(1)}+\alpha_{14}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (124a)
$\dot{n}_{\scriptscriptstyle(1)}+\alpha_{21}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{22}n_{\scriptscriptstyle(1)}+\alpha_{23}\vartheta_{\scriptscriptstyle(1)}+\alpha_{24}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (124b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+\alpha_{31}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{32}n_{\scriptscriptstyle(1)}+\alpha_{33}\vartheta_{\scriptscriptstyle(1)}+\alpha_{34}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (124c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\alpha_{41}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{42}n_{\scriptscriptstyle(1)}+\alpha_{43}\vartheta_{\scriptscriptstyle(1)}+\alpha_{44}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0,\!$ (124d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\alpha_{51}\varepsilon_{\scriptscriptstyle(1)}+\alpha_{52}n_{\scriptscriptstyle(1)}+\alpha_{53}\vartheta_{\scriptscriptstyle(1)}+\alpha_{54}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}$ $=0.\!$ (124e)
The coefficients $\alpha_{ij}$ are given in Table 3.
The quantities in question (123) do not include the gauge function $\psi(\bm{x})$. Consequently, the gauge modes (147) will naturally disappear as a result of deriving the evolution equations for $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$ and $n^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$. This process will be carried out in three steps.
Step 1.
The first step is to remove the explicit occurrence of ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ from Eqs. (124). By differentiating Eq. (124e) with respect to time and eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{{\scriptscriptstyle(1)}}$, $\vartheta_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ with the help of Eqs. (124a)–(124d), one obtains the following equation:
$\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+p_{1}\varepsilon_{\scriptscriptstyle(1)}+p_{2}n_{\scriptscriptstyle(1)}+p_{3}\vartheta_{\scriptscriptstyle(1)}+p_{4}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=0,$ (125)
where the coefficients $p_{1},\ldots,p_{4}$ are given by
$p_{i}=\dot{\alpha}_{5i}-\alpha_{51}\alpha_{1i}-\alpha_{52}\alpha_{2i}-\alpha_{53}\alpha_{3i}-\alpha_{54}\alpha_{4i}.$ (126)
As can be seen from Eq. (125), it follows that
${{}^{3}\!}R_{\scriptscriptstyle(1)}=-\dfrac{1}{p_{4}}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{p_{1}}{p_{4}}\varepsilon_{\scriptscriptstyle(1)}-\dfrac{p_{2}}{p_{4}}n_{\scriptscriptstyle(1)}-\dfrac{p_{3}}{p_{4}}\vartheta_{\scriptscriptstyle(1)}.$ (127)
In this manner, ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ is expressed as a linear combination of the quantities $\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\vartheta_{\scriptscriptstyle(1)}$. Upon replacing ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ in Eqs. (124) by the right-hand side of (127), the following system of equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}+q_{1}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{11}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{12}n_{\scriptscriptstyle(1)}+\gamma_{13}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (128a)
$\dot{n}_{\scriptscriptstyle(1)}+q_{2}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{21}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{22}n_{\scriptscriptstyle(1)}+\gamma_{23}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (128b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+q_{3}\dot{\varepsilon}_{\scriptscriptstyle(1)}^{\textnormal{phys}}+\gamma_{31}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{32}n_{\scriptscriptstyle(1)}+\gamma_{33}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (128c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+q_{4}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\gamma_{41}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{42}n_{\scriptscriptstyle(1)}+\gamma_{43}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (128d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}+q_{5}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\gamma_{51}\varepsilon_{\scriptscriptstyle(1)}+\gamma_{52}n_{\scriptscriptstyle(1)}+\gamma_{53}\vartheta_{\scriptscriptstyle(1)}$ $=0,$ (128e)
where the coefficients $q_{i}$ and $\gamma_{ij}$ are given by
$q_{i}=-\dfrac{\alpha_{i4}}{p_{4}},\quad\gamma_{ij}=\alpha_{ij}+q_{i}p_{j}.$ (129)
It has been achieved that ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occurs explicitly only in Eq. (128d), whereas ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occurs implicitly in the remaining equations. Therefore, Eq. (128d) is no longer required. Accordingly, the remaining four ordinary differential equations are as follows: (128a)–(128c) and (128e) for the four unknown quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
Step 2.
In a manner analogous to the approach undertaken in Step 1, the explicit occurrence of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (128). Differentiating Eq. (128e) with respect to time and subsequently eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{\scriptscriptstyle(1)}$ and $\vartheta_{\scriptscriptstyle(1)}$ with the help of Eqs. (128a)–(128c), results in the following equation:
$q_{5}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+r\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+s_{1}\varepsilon_{\scriptscriptstyle(1)}+s_{2}n_{\scriptscriptstyle(1)}+s_{3}\vartheta_{\scriptscriptstyle(1)}=0,$ (130)
where the coefficients $r$ and $s_{i}$ are given by
$s_{i}$ $=\dot{\gamma}_{5i}-\gamma_{51}\gamma_{1i}-\gamma_{52}\gamma_{2i}-\gamma_{53}\gamma_{3i},$ (131a)
$r$ $=1+\dot{q}_{5}-\gamma_{51}q_{1}-\gamma_{52}q_{2}-\gamma_{53}q_{3}.$ (131b)
As can be seen from Eq. (130), it follows that
$\vartheta{\scriptscriptstyle(1)}=-\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{r}{s_{3}}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{s_{1}}{s_{3}}\varepsilon_{\scriptscriptstyle(1)}-\dfrac{s_{2}}{s_{3}}n_{\scriptscriptstyle(1)}.$ (132)
In this manner, $\vartheta_{\scriptscriptstyle(1)}$ is expressed as a linear combination of $\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$. Upon replacing $\vartheta_{\scriptscriptstyle(1)}$ in Eqs. (128) by the right-hand side of (132), the follwing system of equations is obtained:
$\dot{\varepsilon}_{\scriptscriptstyle(1)}\,$- $\;\gamma_{13}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{1}-\gamma_{13}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{11}-\gamma_{13}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{12}-\gamma_{13}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (133a)
$\dot{n}_{\scriptscriptstyle(1)}\,$- $\;\gamma_{23}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{2}-\gamma_{23}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{21}-\gamma_{23}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{22}-\gamma_{23}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (133b)
$\dot{\vartheta}_{\scriptscriptstyle(1)}\,$- $\;\gamma_{33}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{3}-\gamma_{33}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{31}-\gamma_{33}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{32}-\gamma_{33}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (133c)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}\,$- $\;\gamma_{43}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{4}-\gamma_{43}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{41}-\gamma_{43}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{42}-\gamma_{43}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0,$ (133d)
$\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}\,$- $\;\gamma_{53}\dfrac{q_{5}}{s_{3}}\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+\left(q_{5}-\gamma_{53}\dfrac{r}{s_{3}}\right)\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}$
$+$ $\left(\gamma_{51}-\gamma_{53}\dfrac{s_{1}}{s_{3}}\right)\varepsilon_{\scriptscriptstyle(1)}+\left(\gamma_{52}-\gamma_{53}\dfrac{s_{2}}{s_{3}}\right)n_{\scriptscriptstyle(1)}=0.$ (133e)
It has been achieved that the quantities $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ occur explicitly only in Eqs. (133c) and (133d), whereas they occur implicitly in the remaining equations. Consequently, Eqs. (133c) and (133d) are no longer required. The remaining equations (133a), (133b) and (133e) are three ordinary differential equations for the three unknown quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
Step 3.
The subsequent steps would be to eliminate, in sequence, $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$ from Eq. (133e) with the aid of Eqs. (133a) and (133b). This would result in a fourth-order differential equation for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. It must be noted, however, that this is not a possibility, given that the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$ do not occur explicitly in Eq. (133e). To demonstrate this, Eq. (133e) will be rewritten in the subsequent form:
$\ddot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{1}\dot{\varepsilon}^{\textnormal{phys}}_{\scriptscriptstyle(1)}+a_{2}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=a_{3}\left(n_{\scriptscriptstyle(1)}+\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}\varepsilon_{\scriptscriptstyle(1)}\right),$ (134)
As a consequence of the gauge-invariance of the quantity $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$, the left-hand side of Eq. (134) is also gauge-invariant. It thus follows that the right-hand side of this equation is also gauge-invariant, as will be demonstrated subsequently. Indeed, the result is as follows:
$\dfrac{\gamma_{51}s_{3}-\gamma_{53}s_{1}}{\gamma_{52}s_{3}-\gamma_{53}s_{2}}=-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}.$ (135)
The proof of this equality is straightforward but requires significant computational effort. The algebraic task was performed using the computer algebra system Maxima [52]. The conservation laws (11c) and (11d) and the expressions (6a) yield:
$n_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}=n^{\textnormal{phys}}_{\scriptscriptstyle(1)}-\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}.$ (136)
Substituting (135) into (134) and using (136), we arrive at the evolution equation for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$:
[…]
where the coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by
$a_{1}=-\dfrac{s_{3}}{\gamma_{53}}+\dfrac{r}{q_{5}},\quad a_{2}=-\dfrac{s_{3}}{\gamma_{53}q_{5}},\quad a_{3}=\dfrac{\gamma_{52}s_{3}}{\gamma_{53}q_{5}}-\dfrac{s_{2}}{q_{5}}.$ (138)
The derivation of Eq. (137) shows that both the perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ (19) of the spatial curvature and the divergence $\vartheta_{{\scriptscriptstyle(1)}}$ (20) of the fluid three-vector are included in this equation.
The final step in this process will be deriving an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. The linearized equations (122), from which the evolution equations are derived, are of fourth order. From this system, a second-order equation (137) for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ has been derived. It thus follows that the remaining system (133a)–(133b), and (133e) from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order. As the gauge-invariant quantities, namely $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been used, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated. Consequently, only a first-order evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived. In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the right-hand side of expression (136). By differentiating the left-hand side of this expression with respect to time and using the background equations (11c) and (11d), the first-order equations (122a) and (122b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
$\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right)$
$=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left(n_{\scriptscriptstyle(1)}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}\right).$ (139)
The algebraic task was performed using the Maxima [52] computer algebra software. Due to the equality in Eq. (136), it is possible to replace $n_{\scriptscriptstyle(1)}$ and $\varepsilon_{\scriptscriptstyle(1)}$ in Eq. (A.1) by $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$.
A.2 Perturbation Theory
First Eq. (62a) will be derived. Substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ into Eq. (137) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$ yields the following result:
$b_{1}=2\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1},\quad b_{2}=\dfrac{\ddot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{1}\dfrac{\dot{\varepsilon}_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}+a_{2},\quad b_{3}=a_{3}\dfrac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}}.$ (140)
The coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by (138). With the aid of the computational software Maxima [52], the coefficients (63) of the main text have been calculated.
Finally, Eq. (62b) will be derived. From the definitions (61), it can be deduced that
$n_{\scriptscriptstyle(1)}^{\textnormal{phys}}-\frac{n_{\scriptscriptstyle(0)}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\varepsilon_{\scriptscriptstyle(1)}^{\textnormal{phys}}=n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right).$ (141)
By using the equality (136) and substituting the expression (141) into Eq. (A.1), the following is obtained:
$\dfrac{1}{c}$ $\dfrac{\mathrm{d}}{\mathrm{d}t}\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right]$
$=-3H\left(1-\frac{n_{\scriptscriptstyle(0)}p_{n}}{\varepsilon_{\scriptscriptstyle(0)}(1+w)}\right)\left[n_{\scriptscriptstyle(0)}\left(\delta_{n}-\dfrac{\delta_{\varepsilon}}{1+w}\right)\right].$ (142)
Using Eq. (11d) yields Eq. (62b) of the main text.
Appendix B Verifying the Evolution Equations for Scalar Perturbations
This article does not present a derivation of the linearized equations (13), from which the system of equations (37) for scalar perturbations is derived. Since the latter system is key to the cosmological perturbation theory developed in this article, we will demonstrate its validity. This will be accomplished using three different methods.
The first step in this process is to apply a well-known property of Einstein’s equations and conservation laws. This property stems from the fact that the constraint equations contain only first-order time derivatives of the metric tensor, while the dynamical equations contain second-order time derivatives. It implies that if the constraint equations are satisfied at an initial time, then the solutions to the dynamical equations and conservation laws will also satisfy the constraint equations at all subsequent times. A detailed discussion can be found in Sec. 7.5 of Weinberg’s textbook [42] on the Cauchy problem. The reverse is also true: if solutions to the constraint equations and conservation laws can be found for all times, then they necessarily satisfy the dynamical equations. We will now use this latter property to validate the system of equations (37) for scalar perturbations.
The system (37) encompasses the conservation laws and constraint equations. It consists of four first-order ordinary differential equations and one algebraic equation. Therefore, it can be solved, and its solutions automatically satisfy the system of dynamical equations. This fact will be used to prove the correctness of Eqs. (37).
For $i\neq j$, the dynamical equations (13c) are not coupled to the scalar perturbations. For the case of $i=j$, it is sufficient to consider the contraction of (13c):
$\ddot{h}^{k}{}_{k}+6H\dot{h}^{k}{}_{k}+2\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-3\kappa(\varepsilon_{{\scriptscriptstyle(1)}}-p_{{\scriptscriptstyle(1)}}).$ (143)
Using (20), yields
$(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})+6H(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})-{{}^{3}\!}R_{\scriptscriptstyle(1)}=\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}-p_{\scriptscriptstyle(1)}).$ (144)
Eliminating the second term of equation (144) using the energy density constraint equation (37a) results in
$\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)}+\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(1)}=-\tfrac{3}{2}\kappa(\varepsilon_{\scriptscriptstyle(1)}+p_{\scriptscriptstyle(1)}).$ (145)
As will now be shown, this equation is identical to the time derivative of the energy density constraint equation (37a). Differentiating equation (37a) with respect to time yields
$2\dot{H}(\theta_{\scriptscriptstyle(1)}-\vartheta_{\scriptscriptstyle(1)})+2H(\dot{\theta}_{\scriptscriptstyle(1)}-\dot{\vartheta}_{\scriptscriptstyle(1)})=\tfrac{1}{2}\,{{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+\kappa\dot{\varepsilon}_{\scriptscriptstyle(1)}.$ (146)
Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (11a)–(11c), the momentum constraint equation (37b) and the energy density conservation law (37c) yields the dynamical equation (145). Consequently, Eqs. (37a)–(37c) are correct.
We will now continue with the second method of verifying Eqs. (37). Equations (37d) and (37e) have not yet been verified by the above procedure. This will now be addressed by using the gauge modes. Equations (37) are invariant under the gauge transformations (3) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (9). Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are gauge-dependent. It is therefore evident that the gauge modes
$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,T,p,\theta,$ (147a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}$ $=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H\left(\frac{\tilde{\nabla}^{2}\psi}{a^{2}}-\tfrac{1}{2}\,{{}^{3}\!}R_{\scriptscriptstyle(0)}\psi\right),$ (147c)
must be solutions to Eqs. (37). The gauge modes (147a) follow from (5), with the understanding that $\xi^{0}=\psi(\bm{x})$. The gauge mode (147b) is the covariant divergence of (29). As shown above, the linearized Friedmann equation (37a) is correct. Therefore, it can be used to derive the gauge mode (147c). Substituting the gauge modes (147a)–(147b) into the linearized Friedmann equation results in
${{}^{3}\!}\hat{R}_{\scriptscriptstyle(1)}$ $=4H(\hat{\theta}_{{\scriptscriptstyle(1)}}-\hat{\vartheta}_{{\scriptscriptstyle(1)}})-2\kappa\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$
$=4H\left(\psi\dot{\theta}_{{\scriptscriptstyle(0)}}+\frac{\tilde{\nabla}^{2}\psi}{a^{2}}\right)-2\kappa\psi\dot{\varepsilon}_{{\scriptscriptstyle(0)}}.$ (148)
Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (11a), Eq. (11b), and the energy density conservation law (11c) yields the gauge mode (147c). Expressions (147) imply that the quantities $S_{{\scriptscriptstyle(1)}}$, $\vartheta_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ are invariant under spatial gauge transformations, i.e., transformations for which the condition $\psi(\bm{x})=0$ holds. However, they are not invariant under space-time gauge transformations.
It will now be demonstrated that the gauge modes given by (147) are solutions to Eqs. (37b)–(37e). The first step to doing so is deriving the time derivatives of the gauge modes:
$\dot{\hat{\vartheta}}_{{\scriptscriptstyle(1)}}$ $=-2H\hat{\vartheta}_{{\scriptscriptstyle(1)}},$ (149a)
${{}^{3}\!}\dot{\hat{R}}_{{\scriptscriptstyle(1)}}$ $=-4(\dot{H}-2H^{2})\bigl[\hat{\vartheta}_{{\scriptscriptstyle(1)}}+\tfrac{1}{2}\,{{}^{3}\!}R_{{\scriptscriptstyle(0)}}\psi\bigr],$ (149b)
$\dot{\hat{\varepsilon}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{\varepsilon}_{{\scriptscriptstyle(0)}}=-3\varepsilon_{{\scriptscriptstyle(0)}}(1+w)\bigl[\dot{H}-3H^{2}(1+\beta^{2})\bigr]\psi,$ (149c)
$\dot{\hat{n}}_{{\scriptscriptstyle(1)}}$ $=\psi\ddot{n}_{{\scriptscriptstyle(0)}}=-3n_{{\scriptscriptstyle(0)}}\bigl[\dot{H}-3H^{2}\bigr]\psi,$ (149d)
where Eqs. (11b)–(11d) and Eq. (44) have been used. The validation of the Eqs.  (37b)–(37e) can now be conducted.
First, substitute the gauge modes $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq. (37b), to obtain an equation from which the gauge function $\psi$ disappears. Next, dividing this equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying the result by $\tfrac{3}{2}H$ yields the time derivative of Eq. (11a). Second, substitute the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37c) eliminates the term with $\dot{H}$. Next, dividing the resulting equation by $\psi$ and using the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by the use of Eq. (11c) yields an identity. Third, substitute the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37d), and using the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq. (11c) one arrives at an identity. Finally, by substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq. (37e) the term with $\dot{H}$ vanishes. Using Eq. (11d), one arrives at an identity. It has been demonstrated that the gauge modes (147) satisfy Eqs. (37b)–(37e). Consequently, it can be concluded that these equations are correct. Given the established validity of Eq. (37a), it can be deduced that the system (37) is accurate as well.
As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (151) for the gauge-dependent relative perturbation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs. (37). In fact, the well-known evolution equations for $\delta$ are obtained in the cases of a radiation-dominated universe (154) and a universe after matter and radiation decouple (157).
Consequently, it can be concluded that the system of equations (37) has been verified in three distinct ways.
Appendix C Conventional Evolution Equation for Relative Density Perturbations
In this appendix the usual perturbation equations for relative perturbations in a flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, flrw universe is compared with the perturbation theory developed in this article. To that end, the evolution equation for the relative perturbation $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$ is derived using the systems of background equations (75) and evolution equations (122) for scalar perturbations. In this appendix, we assume $c=1$ and $\Lambda=0$.
In the existing literature, the particle number density is usually not taken into account. Therefore, we take the equation of state for the pressure $p=p(\varepsilon)$, so that $p_{n}=0$}. From (14) and (40) we find that $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$. Using the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs. (122) can be rewritten as follows:
$\dot{\delta}+3H\delta\bigl(\beta^{2}+\tfrac{1}{2}(1-w)\bigr)+(1+w)\left(\vartheta_{\scriptscriptstyle(1)}+\dfrac{{{}^{3}\!}R_{\scriptscriptstyle(1)}}{4H}\right)=0,$ (150a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\frac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0,$ (150b)
${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}+2H\,{{}^{3}\!}R_{\scriptscriptstyle(1)}-2\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)\vartheta_{\scriptscriptstyle(1)}=0,$ (150c)
where Eq. (75b) has been used. First, differentiate equation (150a) with respect to time. Then, eliminate the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ using equations (44), (75a) and (75b) and the perturbation equations (150b) and (150c). Finally, eliminate ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ using Eq. (150a). This results in the system of equations:
$\ddot{\delta}+2H\dot{\delta}\bigl(1+3\beta^{2}-3w\bigr)$
[…]
+ $\qquad+\,4w-6w^{2}+12\beta^{2}w-4\beta^{2}-6\beta^{4}\Bigr)-6\beta\dot{\beta}H\biggr]\delta$
$\quad=-3H\beta^{2}(1+w)\vartheta_{\scriptscriptstyle(1)},$ (151a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H(2-3\beta^{2})\vartheta_{\scriptscriptstyle(1)}+\dfrac{\beta^{2}}{1+w}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (151b)
In this procedure the computer algebra program Maxima [52] has been used. The gauge modes (147) are for the system (151) given by
$\hat{\delta}(t,\bm{x}):=\dfrac{\psi(\bm{x})\dot{\varepsilon}_{\scriptscriptstyle(0)}(t)}{\varepsilon_{\scriptscriptstyle(0)}(t)}=-3H(t)\psi(\bm{x})\bigl(1+w(t)\bigr),$ (152a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}(t)}.$ (152b)
From the Friedmann equation, $3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}$, it follows that $H\neq 0$ for $\varepsilon_{{\scriptscriptstyle(0)}}\neq 0$. This implies that the gauge mode $\hat{\delta}(t,\bm{x})$ is nonzero, regardless of the scale of the density perturbation. Consequently, the general solution of the system (151) has no physical meaning. Therefore, the system (151) cannot be used to study the evolution of density perturbations.
C.1 Radiation-Dominated Era
In this era, the pressure is given by a linear barotropic relativistic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and $w=\tfrac{1}{3}$}. The standard equation is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=0.$ (153)
In contrast, Eqs. (151) result in the system
$\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-\tfrac{4}{3}H\vartheta_{\scriptscriptstyle(1)},$ (154a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+H\vartheta_{\scriptscriptstyle(1)}+\dfrac{1}{4}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (154b)
The gauge modes (152) are solutions to the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
For large-scale perturbations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution to the system (154) is
$\delta(t,\bm{x})=(c_{1}t-2\psi(\bm{x})t^{-1})+\tfrac{9}{8}t^{1/2},$ (155a)
$\vartheta_{\scriptscriptstyle(1)}(t,\bm{x})=-\dfrac{\nabla^{2}\psi(\bm{x})}{a^{2}}+\tfrac{9}{8}t^{-1/2},$ (155b)
where it is used that $H=\tfrac{1}{2}t^{-1}$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}t^{-2}$. The expression between parentheses in (155a) is the solution to the homogeneous part of Eq. (154a). The particular solution $\tfrac{9}{8}t^{1/2}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\tfrac{9}{8}t^{-1/2}$. The right-hand side of Eq. (154a) is not equal to zero, which means that Eq. (153) is incomplete. As a result, it does not provide all physical solutions. Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale perturbations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$. Because the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into Eq. (82), oscillating relative density perturbations (87) with an increasing amplitude are obtained instead of the constant amplitude that follows from (153).
C.2 Era after Decoupling of Matter and Radiation
The standard equation is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=0,$ (156)
where $\beta$ is given by (93). In this era, the equation of state for the pressure is according to thermodynamics given by (89). Therefore, (92) and (93) imply that $w\approx\tfrac{3}{5}\beta^{2}\ll 1$, so that with $T_{\scriptscriptstyle(0)}\propto a^{-2}$ one obtains $\dot{\beta}/\beta=-H$. Using Eq. (75a) one can derive the following result: $6\beta\dot{\beta}H=-2\kappa\varepsilon_{\scriptscriptstyle(0)}\beta^{2}$. Upon substituting the latter expression into (151a) and neglecting $w$ and $\beta^{2}$ with respect to the constants of order one, the system (151) results in
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=-3H\beta^{2}\vartheta_{\scriptscriptstyle(1)},$ (157a)
$\dot{\vartheta}_{\scriptscriptstyle(1)}+2H\vartheta_{\scriptscriptstyle(1)}+\beta^{2}\dfrac{\nabla^{2}\delta}{a^{2}}=0.$ (157b)
The gauge modes (152) are solutions to the system (157) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution. Consequently, for the system (157) $\psi$ is an arbitrary infinitesimal constant $C$. This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (152b), vanishes identically. Consequently, the relativistic gauge transformation (3) with $\xi^{\mu}(t,\bm{x})$ given by (9) reduces to the Newtonian gauge transformation (52). This is to be expected, since a cosmological fluid for which $w\ll 1$} and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (89).
For large-scale perturbations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution to the system (157) is
$\delta(t)=(c_{1}t^{2/3}-2Ct^{-1})+\tfrac{7}{9}t^{-5/3},$ (158a)
$\vartheta_{\scriptscriptstyle(1)}^{\textnormal{phys}}(t)=-\tfrac{7}{9}t^{-4/3},$ (158b)
where $H=\tfrac{2}{3}t^{-1}$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}t^{-2}$, and $\beta^{2}\propto t^{-4/3}$ has been used. The latter proportionality follows from (92) and (93), $T_{\scriptscriptstyle(0)}\propto a^{-2}$, (94), and $a\propto t^{2/3}$, (104). The expression between parentheses in (158a) is the solution to the homogeneous part of Eq. (157a). The particular solution $\tfrac{7}{9}t^{-5/3}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\tfrac{7}{9}t^{-4/3}$.
The right-hand side of Eq. (157a) is nonzero for all scales. Therefore, Eq. (156) is incomplete and does not yield all physical solutions.